A Complete Bibliography of Publications in IEEE Transactions on Software Engineering (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

09 April 2018 Version 1.01

Title word cross-reference

1 [TS85a]. _{K} [Cho89]. γ [JN84, Jaj87]. ∞ [Cho89]. λ [KS83]. N [Avi85, BKL89, San89c].

*MOD [Coo80].

-Acyclic [Jaj87, JN84]. -Expressions [KS83]. -tree [San89c]. -Version [Avi85, BKL89].

1 [Bry81]. 11 [Ano85a]. 13 [Ano87a]. 18 [Ano82b]. 1978 [Shn81]. 1981 [Ano82-30]. 1985 [Ano85-72, Ano85-73].

2.1 [MMS88].

3 [Tra86]. 3081 [IR85]. 3rd [Ano82i].

4 [BHP84]. 4.3BSD [GZ89].

5th [Ano81c].

6 [Ano80a]. 6th [Ano82c].

7 [Ano81b].

’81 [Mus83]. 8th [Ano81d].

9 [Ano83b]. 9th [Ano87b].

ABE [ELHR88]. Absence [MM87]. Absolute [Row88]. Abstract [BHJ^{+}87, BW80, FM85, HM83, Mat84, MY80b, Moi82, Mus80, Pag83, WY84, Whi83, BU86, Kim86]. Abstraction [CL83, Gut80, LL89a, CH88, MI89, Wat88]. Abstractions [Bas84, Cla82, CCW85]. Academic [Wor87]. Access [ABL83, DTV87, EM82, EHS81, KL87, Rou82, WH84, CGW89, TS86]. Accesses [WYD87]. Accommodation [Smo81]. accuracy [GZ89]. achieved [GZ89]. achieves [JM89]. Achieving [FR82, GMK88]. Acknowledgment [Ano80b]. Across [BSP83, Bok81, Nic89, SF88]. Actions [Mck85, Sin85, Wei87a, JC86, Tay86]. Activities [Beh83]. Activity [BP81]. Acyclic [Jaj87, ST87, JN84, KWN88]. Ada [Kar90, MO90, Obe88, TS85a, Ano86a, BTM89, CS85b, FM89, GR88, Ger84, HW89, JKC89, KSB89, KU87, LH83a, LM84, MSS89, Raj87, Sha88, VMBK89]. AdaPIC [WCW89]. Adaptive [ATR82, ELZ86, SC84, YSLC85]. Adding [ADWR86]. Addressing [AB85, SVR86]. adequacy [Wey86, ZG89]. ADI [Tam80]. Adjustment [MM82]. Advance [Ano80d, Ano80c, Ano81f, Ano81e, Ano82d, Ano82e, Ano82f, Ano82g, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano84b, Ano85b, Ano86b]. Advanced [SC87, TS85a]. Advances [Fag86, LS85]. Aesthetics [Gai85]. Affecting [Nes87]. AFFIRM [STE^{+}83, Mus80, STE^{+}82]. aggregation [CG86]. Agreement [CC85b, Per85, PT86]. AI [Bob85, Mos85]. Aided [CS85a, BMPP89, LL89b]. airplane [MW88]. Al [KMB^{+}85, CKST89a]. Alfu [SG81]. Alfu-N2merik [SG81]. Algebra [Ben84, CG85a, PS85, Adl88, Agr88, Bra88, ML89]. Algebraic [Foo87, HMS^{+}84, LM84, Moi82, LL88]. Algorithm [BR87a, Bok81, BDMS81, Car87, CL85, CC89a, CC85b, Dal87, Dea81, FO87, KM83, Kan85, KKT80b, NXG85, SN85, SK85, Aoe89a, BNT86, CKST89a, CKST89b, Cid89, ESL88, Hac89, JM89, KWN88, MI86, Wan96, YLN88, Yan89, ZTSN88].

Algorithms [AB85, AEH84, AHY83,

BGL83, BH83b, BA83, BD81, Cha82, CH81, KS87, LW84, LA81, MS88, Ozs85, RB83, RS88b, SSS82, Smi81a, SS81, Sta85, WY85, YCC89, vBV87, AA89, Bag89, CKM88, JMG88, KGM88, QI88, YW88a]. Allocating [FB89, Kau90, KW85c, Tow86]. Allocation [BD81, BDMS81, CKS85, HFG89, Jom81, Sun87, YSLC85, MCH88, WT86]. Allow [Sil83b]. ALMA [VDD^{+}88]. Alpha [Agr88]. Alphard [BS80]. Alternative [ACM87, HD82, Par80, DBC88]. Alternatives [KL84, RE89]. Ambiguity [Har80]. Analogy [Dav80]. Analyses [CDL84, CH81]. Analysis [ABM87, AR81, BHN80, BSP83, Ben87b, BG85, BA83, BW80, CS85a, EL85, EM82, Fos80, Fre87a, GMT85, Goe80, HKMY87, HP82, HD82, HP80a, HV87, HB85, IV85, Lee87a, Lee87b, Leu83, LS87, LA81, Mac83, MMA87, Men82, MCF87, MR87, NS83, Ng81, NKT87, Ozs85, Pap81, RJ87, RC85, ST87, Sar84, Sca84, SW78, SCD83, SSL85, Smi81a, SK85, SM86, Sun87, TO80, WB85a, WO85, WHH80, YHYM87, ABM88, AS88, Aky88b, Aky88a, ADWR86, BEDBT89, BBG86, BDZ89, CMNK88, DC89, EFRV86, Har88a, How86, JM86, KHR86, Kun86, Lau88, MBB^{+}89, NTT88, OP89, PA86, PS89, RE89, Sch86, SP88, Sym88, TS86, WCW89, YT88, YSD88]. analytic [LL89c]. analytic/empirical [LL89c]. Analyze [HO83, TR89, Vos80]. Analyzers [AR81]. Analyzing [BGL83, BH83b, LH83b, MI89]. Anatomy [SJ83]. Annotated [SLH81a]. Annotation [DM81]. Annotations [Sch82]. Announcement [Ano80c, Ano81g, Ano81f, Ano81e, Ano82d, Ano82e, Ano82f, Ano82g, Ano83c, Ano83d, Ano83e, Ano83-39, Ano84b, Ano84j]. Announcing [Ano80d, Ano82i, Ano82h]. Annual [Ano87b]. Anomalies [ASP82]. Anomaly [JA84, TO80]. Answer [Bob85, PM89]. Answering [OC82].

Antipalindromic [MS87]. Antisampling [Row85]. Apers [CKM88]. APL [KW85b, MM82]. Applicability [Goe85, DL88a]. applicable [FW88]. Application [AB85, Ano87-66, Boe81, Gho85, GCJ^{+}87, Hen80, LNT85, MMA87, MO87, Pap81, RMB^{+}81, RDU85, Smi81a, WO80, BF86, HM83, KB88, Kun86, LNE89, OM88, San89c, Sch89, SS88, TR86]. Applications [Ano82t, Ano82-39, Dow85, Hua80, Hua85, IS89, PBC85, RB83, YO85, Ano86a, Cla88, Dun86, Fis88, KH88, Neu86, Orm88, Yan89]. Applicative [KS83, Pag87]. Applied [Fea82, KW85b, Bar89]. Applying [Hay85, Smi88]. Apprentice [EFRV86, Wat82, Wat85]. Approach [Avi85, BCM87, Bas85a, CS87a, Cri85, Dow85, Fre87a, GMT85, GS78, Gar85, LLS81, LR80, LM84, Nei84, Ng81, Red83, RW87a, SC87, Sha82, SG84, YC80a, YYS81, Zav82, AH89, Arm89, BKHW89, Dow86, EZ89, Gra86, HM83, How86, Kim89, LL88, Mun88, TTNM89, YM89]. Approaches [BR81, Chi85, Kim82, Ale88, RS88a]. Appropriate [Sub85]. Approximate [AS88, OP89, PA86, Sin83, Aky88b]. APS [RDU85]. APSE [Obe88]. arbitration [BDZ89]. Archetype [Mea88]. ARchitects [EFRV86]. Architecture [MB83, Mil87]. Architectures [BD85, HD82, YHYM87, HL88, Mea88, YTT88]. Area [MINH87, THJ87, BDHH89, CV89, LL89c]. ARES [IH86]. Argos [Hol87, TR86]. arithmetic [SD89c]. Array [NH81, PCMP85, Aoe89a, SD89a]. Arrays [Ell81, GL85, ML86]. Arrivals [MP83]. Art [Ano83-40, Jon84]. artificially [Sim86]. ASLAN [AK86]. Aspect [Gai85]. aspects [HSN86b]. Assembled [HO83]. Assemblers [TN80]. Assembly [Reu82]. Assertion [Hol87, Kin80]. Assertions [Bas80a, LB84]. Assessing [SS85]. Assessment [Fry85, TM85, BF86]. assignment [CY89, May89]. Assignments [Bok81]. assistance [RS88a]. Assistant [SLH81a, KK88, Kor88, SSB88]. Assisted [CJL^{+}85, HOPW87, FM89]. assumption [KL86]. Assumptions [Ben87b, Goe85]. Assurance [BP84b, SM85]. Asymptotic [CKS85]. Asynchronous [RH80, CR86]. Atomic [JC86, LG85, Sin85, Tay86]. Atomicity [Sil83a]. Attribute [KP82, Pap81, Des89, YM86]. Attributed [LS83, LNE89]. Auditable [SC87]. Auditing [CO82]. Augmented [CR83]. Augmenting [CS85b]. Authentication [CV89]. Author [Ano86c, Bab82, SLH81b]. Authority [MM87]. authors [Ano80-40, Ano81-38, Ano82-38, Ano83-36, Ano84-28, Ano85-68, Ano85-69, Ano85-70, Ano87-61, Ano87-62]. Automata [Ano81d]. Automated [And88, Hol87, RJ87, SM85, VGV80, KK88, RDU85]. Automatic [Bal85, Bar85, BS82, Rei87, RK84, SW87, Tak80, WS81, BS86, RS88c, YM89, YM86, Tam80]. Automatic-Controller [Tak80]. Automatically [HO83]. Automating [Fic85, Kan85]. Automation [SLTC82, SKY89]. Autonomous [KMB^{+}85]. Autoregressive [SS85]. Availability [MPS85, GMK88, JM89, SM86]. availability/reliability [SM86]. Available [MM87]. Avoidance [Lom80]. Awards [Ano83-38]. Axiomatic [BN80a, Chi85, Ale88]. Axiomatizing [Wey86, ZG89].

B [KW82]. B-Trees [KW82]. Backup [ABG84]. Balancing [CA82, LK87, NXG85, NH85, MI86, NH86, Zho88]. Base [CGP81]. Based [AB85, BCM87, CR87, Chu87, CTDS85, Gud80, HK81, JW83, JS85, KL87, Kim82, KS83, Mac83, RB82b, Rus89, SN85, SKW85, TN80, Wat82, AH89, CMNK88, CKST89a, CKST89b, CV89, DI89, Gol86, HM89, JP89, KA88, LHL89, ML89, MCS88, MW88, RSBV86, Sam89, SSB88, SZ88, TI88, TL89, TTNM89, TMTI86, US86, Wan96, WW88, YLN88, YD86]. Bases [MO87]. Basic [Har85, Jew85]. Basics [Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g]. Basis [Ber87, EL85, Mad88]. Batch [LC84, LB84]. batched [MK89]. Bayesian [Jew85]. Be [Lit80, Sim86]. Behave [Sch82]. Behavior [AEH84, Bry81, CGH83, GMT85, Haa81, Old83, PBC85, PS87, RJ87, WEHR84, FGM89, Lau88]. Benchmarks [YHYM87, HSN86a]. benefits [HB88]. Berkeley [GZ89]. Best [WY85]. Best-First [WY85]. better [DH88]. Between [Dav80, KBB87]. Beyond [Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g]. Binary [DTV87, Man84, CTT89, JP89, MI86, USV88]. binding [Win88]. Bindings [HB85]. Black [RR87]. Blisses [Bre80]. Block [Har85, LF84, LB85]. Block-Structured [LF84]. Blocking [AV89, BD89, Ger84, Per81, Aky88b, Aky88a, BBG86, PA86]. Blocks [SL84]. bookshop [Gra88]. Boolean [Har80]. Bound [WY85, YW88a]. bounds [Row88]. Box [RR87]. Branch [WY85, YW88a]. Branch-and-Bound [WY85, YW88a]. Branches [Chu87, Kau90, Tow86]. Breaker [Ano84c]. broadband [Jal89b]. Broadcast [JL87, Top85, CL89]. broadcasting [CC89b, Geh84]. Broadcasts [BD85]. BSP [Geh84]. bucket [YD86]. Budgets [GM87]. Buffered [EM82]. buffers [OP89]. Building [WPS86]. Bus [EHS81]. Business [Mat87, SLTC82]. Byzantine [CC85b, Per85]. Byzantium [BD85].

C [CGR89, GR88, SS89, VLC88, WO85]. Caching [Ter87]. Cactis [HK88a]. CAD [BKKG88, BKHW89]. CAIS [Obe88]. calculus [OW89]. Call [Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano81g, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano82t, Ano82p, Ano82q, Ano82u, Ano82r, Ano82s, Ano82-39, Ano83s, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano84d, Ano84e, Ano84f, Ano84g, Ano85-35, Ano85-36, Ano85v, Ano85w, Ano85x, Ano85y, Ano85z, Ano85-27, Ano85-28, Ano85-29, Ano85-30, Ano85-31, Ano85-32, Ano85-33, Ano85-34, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano87t, Ano87u, Ano87v, BCL^{+}87, BMOW87, LG85, Kun86, PS88]. Call-by-Need [BMOW87]. call-return [Kun86]. Calls [FLS87]. Cambridge [Har85]. Can [Lit80]. capabilities [ADWR86]. Capability [KL87, IH86]. Capability-Based [KL87]. Capture [DW81]. Capture-Recapture [DW81]. Cartesian [CDL84, Sun87]. Case [MG83, Sil83a, Oss89, San89b, San89a]. Cases [Fos80, BMD89]. Casting [NHS83]. Causal [BH83b]. Cautious [IKK88]. CDL [ST89]. Cedar [Tei85]. Cell [Sip81, Sil84]. Cellular [HS82]. Cellular-Logic [HS82]. center [TR86]. Central [AV89]. Centralized [Tho85]. certain [HSN86a]. Certification [Kar85, Mac82, MM81, RR87]. Certifications [BCFP87]. Certifying [CDM86, CDM89]. chaining [TT86]. Chameleon [MKNS89]. Changes [WB85a]. Changing [KW85a]. Channel [HKMY87]. Channels [Gir87]. characterization [AA89]. Charlotte [FSAC89]. Charts [Pag87]. Checking [Egg81, KL87, YC80a, Lic86]. Checklist [Ano82-30, Ano82-29]. Checkpointing [KT87, SA89]. Cheung [Hir90]. Chi [Ale88]. Chief [Ano84-29]. Choice [McG84]. Clarification [LB88]. Class [Kun82, Agr88, AI89, BYC88, CG86]. Classes [AV89, NH85, NH86]. Classical [Le 87]. classification [MW88]. Cleanroom [SBB87]. clock [GZ89]. Close [CHR82]. Closed [Gou84, Kam84, Aky88b, CS89a, CG86]. closures [Wan96, YLN88]. cloze [HZ86]. Clustering [BKKG88, CF81, CJS87, HB85]. Cobol [Leh89]. CODASYL [LSK88, RK84]. Code [AG83, CLD82, Gaf84, LG84, Lip82, Mac82, MM81, Ram85, Sam82, Lip86, Ste86]. Codes [RW87b, Fal88]. codeword [BL89a]. Coding [KB81]. Coefficient [SS85]. Cognitive [Cou83]. coherency [DIRY89]. Coincident [EL85, LM89]. Collecting [BW84]. Collection [Cha80]. Collision [EHS81]. Collision-Free [EHS81]. Colloquium [Ano81d]. combination [WDIY89]. Combined [Ver82, YW88b]. Combining [BBG86, MG83, RC85, YT88, Fea89]. Come [Ber85b]. Command [Sno83, Mea86, SS88]. Comments [Ale88, Ber85c, GS81, Hir90, HH82, JV83, Kar90, Kau90, LW84, LLS81, LR87, Lin89, Lip86, MO90, Ram87, Rob81, Shn81, Ste86, ZG89, vBV87, CKM88, Ten88]. Commercial [Jef87]. Common [Obe88]. Communicating [Gou84, GL85, Rus80, CS89a]. Communication [Ano80s, EHS81, Gar85, LA81, MY80a, Nat85, RJ87, STE^{+}82, STE^{+}83, TB85, ZR87, vBGS87, BWWH88, BS86, CHKH88, CL88, EFK89, Kam86, Lin89, Nat86, PT86, SVR86, WW89]. Communications [Gai85, PR85, HW89, Kar90]. Commutable [Sin86]. Compact [DTV87]. Compaction [AM88, Ver82]. Companson [FWA87]. Comparative [WS83]. Comparing [BR81, BS87, HHR89, DBC88]. Comparison [BZ80, BN80a, BD81, Chi85, Jom81, Ale88, BKL89, HL88, Leh89, Nta88, TT86]. Comparisons [IMOL84, Dow86]. CompCon [Ano80r]. Competing [Goe80, SW78, AGCL86]. Compilation [CVGM80, CH81, LF84, PP83, LQ89, SF88].

Compile [Har88b, SHMM89].

Compile-time [Har88b]. Compiler

[BS82, FK80, ML81, PCMP85, Rei87, SJ83, CTYY89, MDSW88, VLC88]. Complementary [BR87b]. Complete [Mag87]. completely [MK89]. Completeness [How82, MS85, Jal89c, VS89]. Complex [Hen80, PCW85, KCB88]. Complexity [BZ80, BH83a, KR87, Pas85, Sta82, Tai80, DL88a, LDF88, Sha88, Wey88]. Component [YT87]. Components [LM87, LM84, Nei84]. composing [Mea86]. composite [Des89, KB88]. Composition [MO87]. comprehensibility [HZ86]. Comprehensive [Jai87]. Compression [Bas85b, Kun82, YL80, CIK88]. COMPSAC [Mus83]. Computation [Che83, Whi83, CCG^{+}88, USV88]. computational [VS89]. computations [HYSY88]. computed [TT86]. Computer [Ano80s, Ano80-30, Ano80x, Ano80-31, Ano80y, Ano80-32, Ano80z, Ano80-33, Ano80-27, Ano80-34, Ano80-28, Ano80-35, Ano80-29, Ano81-31, Ano81x, Ano81-32, Ano81-30, Ano81y, Ano81-33, Ano81z, Ano81-34, Ano81-27, Ano81-35, Ano81-28, Ano81-36, Ano81-29, Ano82-31, Ano82w, Ano82-32, Ano82-30, Ano82x, Ano82-29, Ano82-33, Ano82y, Ano82-34, Ano82z, Ano82-35, Ano82-27, Ano82-36, Ano82-28, Ano82-40, Ano83s, Ano83-29, Ano83w, Ano83-30, Ano83x, Ano83-31, Ano83y, Ano83-32, Ano83z, Ano83-33, Ano83-27, Ano83-34, Ano83-28, Ano84z, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85-52, Ano85-40, Ano85-53, Ano85-41, Ano85-54, Ano85-42, Ano85-55, Ano85-43, Ano85-56, Ano85-44, Ano85-57, Ano85-45, Ano85-58, Ano85-46, Ano85-59, Ano85-47, Ano85-60, Ano85-48, Ano85-61, Ano85-49, Ano85-62, Ano85-50, Ano85-63]. Computer [Ano85-51, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87w, Ano87x, Ano87y, Ano87z, Ano87-41, Ano87-29, Ano87-42, Ano87-30, Ano87-43, Ano87-31, Ano87-44, Ano87-32, Ano87-45, Ano87-33, Ano87-46, Ano87-34, Ano87-47, Ano87-35, Ano87-48, Ano87-36, Ano87-49, Ano87-37, Ano87-50, Ano87-38, Ano87-51, Ano87-39, Ano87-52, Ano87-40, Bab81, BHN80, Bas85d, Beh83, Ber83, BCL^{+}87, BDMS81, EHS81, Fry85, Gai85, HOPW87, HH82, Jai87, KK83, LL89b, NKT87, Nis83, PP83, TB85, TL83, WL85, Was85, Wes82, WC80, YSLC85, BMPP89, EZ89, Gra86, Kam86, Neu86, Sch89, SJ86, Woo86, YH88]. Computer-aided [LL89b]. Computers [Ber85b, Sum87, CHKH88]. Computing [Ano82i, BDMS81, GMGK84, LC84, LS85, Sil84, Svo84, YYS81, BDHH89, CK88a, CK88b, HYSY88, Kim89, Len89, PI86, TMTI86]. Concept [Chu87, MY80a, Sil84, SY86]. Concepts [LB85, NHM^{+}87, Sum87]. Conceptual [Fre87a, Kun89, LM89]. Concerning [RR87]. Concurrency [BGL83, BH83b, Car87, FO87, HD89, Kha84, KBFL85, KW82, LB87, Ozs85, Tay86, ABM88, Bas88, DIRY89, IKK88, ST89, YT88]. concurrency-coherency [DIRY89]. Concurrent [BCFP87, CS84, Dav87, Mag87, Man84, MY80a, MR87, Mur80, PS87, RH80, SPSP87, TO80, Top85, WS83, YC80a, ADWR86, BEDBT89, BDV86, CH88, EFRV86, FM89, DV88, FGM89, GS88, GWW89, Kim88, LB88, MCS88, MI89, YH88, CGR89, GR88, Pas85]. Condensation [AC82]. condition [FN89, WW89]. conditions [CS89a]. Conference [Ano81c, Ano82i, Ano82t, Ano82c, Ano87b, Ano87-69, Dhi86a, Dhi86b]. Conferences [Mar85]. Confidence [Cav85]. Configuration [Ber84, Ber85c, KM85]. Configurations [NS87, PA86]. conformance [BDZ89, MMS88]. congestion [Fon89]. Conic [MKS89]. consensus [AA89, LA86]. Consequence [CV81]. consideration [FGM89].

Considerations [AFS81, BDBPLR84, CF84, Fer86, Pas85, Sam82, YL80, KS89]. Consistency [GM82, JM89, Lic86]. Consistent [SA89, BKL89]. Constants [Kam84]. Constrained [ADWR86, RW87b]. Constraints [BR87a, BDRZ85, Das85]. Construct [Kes81, LB85, XH89]. Constructing [Dal87, Fea89, Fre87a, MKS89, Nei84, SA88, YCC89]. Construction [BBG^{+}85, JS87, BMPP89, CCRDL88, Gol86, SRVO88]. Constructor [FKM85]. Constructs [Das85, RS83, ST89]. containing [Kau90, Tow86]. Content [DW81]. Contention [Smi81b]. Contents [Ano86v, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano84-30, Ano84-31, Ano84-32, Ano84-33, Ano84-34, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s]. Context [Bas84, Kun89, HH86]. Continuously [MM87]. Control [ABL83, BZ80, BGL83, BH83b, Car87, CO82, CP87, CL83, Cot84, Dav87, EHS81, FO87, KMS82, KMB^{+}85, LA83, LB87, LR80, LM82, Lud83, Oku85, Ozs85, RMB^{+}81, SKhL89, WH84, YC80a, Bas88, BM86, Dhi86a, Dhi86b, HD89, IKK88, JM89, LL89a, Sie88a, Sie88b, War86, WCW88, WCW89, YH88]. Controllability [Woo86]. Controlled [BR81, Fry85, Rom87, Woo86]. controlled-share [Woo86]. Controller [Tak80]. Controllers [Ram85]. Controlling [BH81, BP88]. Controls [Est87, DIRY89]. Conversation [Kim82, TH86, WF89]. Conversion [MFKS85, Lam88b, Lam88a]. Cooperating [SL84]. Cooperative [KS87]. Coordination [Sin85, ABM88, Kim88]. Copies [CGP81, YMW^{+}85]. Copy [MW82]. copyright [Ano80-36, Ano81-37, Ano82-37, Ano83-35, Ano84-27, Ano85-64, Ano85-65, Ano85-66, Ano86-28, Ano87-53]. Corporate [LPBD87]. Correct [BS80, Bli81, Cri84]. Correction [BC80, CKST89a, CDM89, KBFL85, Lam88a, Mac82, NH86, PC86a, San89b, STE^{+}83]. corrections [Lip86]. Correctness [Bas85d, ESL88, Kin80]. Corrigendum [Oss82]. Cost [Jef87, Sta82, YMW^{+}85, BMD89, Hag89]. Costs [War83, BP88, Gaf86]. Counting [Ben84]. Coupled [Ber85a, Sco87]. coupling [CDY86, WDIY89]. Course [BDBPLR84, BR87b, Kee87, MBGH89]. Courses [LM87]. cover [Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano84r, Ano84s, Ano84t, Ano84u, Ano84v, Ano84w, Ano84x, Ano86-29, CS89a]. Covers [Gou84, CIK88]. Covert [Gir87, HKMY87]. Crash [Shr81, SS83]. create [Smi88]. Criteria [IMOL84, RW87b, Tai80, CPRZ89, FW88]. Criterion [Nis83]. Critical [KSB89, LLS81, MO90, SCD83, Dun86, Neu86]. Critique [GS81]. Cross [TN80]. Cryptography [Gud80]. Crystal [DFS87]. CSM [ZX87]. CSMA [ZR87]. CSP [JC86, Sha85, Sil83b]. CTDNet [GWW89]. Curve [Par80, WEHR84]. Customized [CG85c, RS88a]. Cycle [BD89, CJS87, KS87, MS87, SJ87, WEHR84, DBC88, EZ89, TSSP86, YNTL88]. Cyclic [BD89, OP89]. cyclomatic [RM88].

D [Kau90, CF81]. DAG [BKKG88]. Data [Ano86w, Ano87-69, BH80, Bab81, BD82, BSP83, BW84, Bas85b, BL84, Bec80, BS80, Ber85a, Ber85b, BW80, BDRZ85, CA85, Chu87, Cla82, CCW85, CGP81, DAT87, FM85, Goo81, GM87, HR81, HH82, HB85, JA84, JW83, JTS^{+}85, KM83, Kun82, LK83, Lee87a, Luc85, ML86, MI89, Moi82, Mus80, PR85, Phi84, RW85, Red83, Rus89, Sar84, Sin83, TO80, TMB80a, TMB80b, WYD87, WB85a, Whi83, WO85, Won83, YWH82, YMW^{+}85, YL80, ZY87, Adl88, AA89, BL89a, BEDBT89, BNT86, BU86, CPRZ89, DH88, DIRY89, FW88, Leh89, Lip86, MKNS89, MIBY88, OM88, PS89, Tra86, War86, Wey86, WDIY89, ZG89]. data-translation [MKNS89]. Database [ACM87, Ano83t, ABG84, BHN80, BCFP87, BDF^{+}82, BDRZ85, CNW83, CC80, CF81, CR87, CCW85, DH84, DAH^{+}87, FKM85, FO87, GM82, GM85, HK84, HD82, HR82, HS82, JW83, JV83, KL84, KD83, Lee87b, LB84, LB85, MINH87, MW82, MIMH85, Ng81, OO85, Ozs85, Per84, PB87, RK84, Rou82, SL87, SSL85, SW85, SK82, SG84, SS81, Vos80, YHYM87, YMW^{+}85, BBG^{+}88, BR89a, Bra88, CGW89, CYDA88, CL89, CMNK88, CY89, DH88, DL88b, DYS89, Gho86, Gho89, HL88, HSN86a, HK88a, IKK88, IH86, JKT88, JC88, KL88, KH88, LSK88, MCS88, OM88, Orm88, QI88, RFS88, TMB86, VS89, Wan96, YLN88, YBL88]. Databases [Ano83-40, Bas85b, CO82, DAT87, Geh82, Gho85, JS87, Lee87a, Lee87b, Mot87, OC82, SA89, Tho85, WL85, BKKG88, GMK88, HD89, HSN86b, KB88, LSK88, LNE89, PV88, RS88c, TAO89]. Dataflow [HYSY88, KBB87, OJ84]. DBMS [BD82]. Deadline [CC89a, WR80]. Deadline-Oriented [WR80]. Deadlock [CJS87, Ger84, GS80, HR82, JV83, Lom80, SN85, WB85b, CKST89a, CKST89b, ESL88, WW89]. Deadlocks [ACM87, Lin89, MSS89, Nat86]. Dealing [ACM87]. debugger [BDV86, BTM89]. Debugging [GMGK84, Tra83, FM89, LL89a, Moh88]. December [Ano82-30]. decentralized [Gra86, LA86]. Decide [BR87a]. Decision [BR87a, Mur80, SP88]. Decision-Free [Mur80]. Decomposition [CC80, CG86, Adl88, AI89, Hać86]. Deduction [Yan89, YCC89]. defect [YSD88]. Defining [CCW85, CS89a]. Definition [DH84, RS83, HW86b]. degrading [MD86]. Delay [Von88, Cho89, Har89]. Delay-independent [Von88]. Delimited [Pro82]. demonstration [FM89]. Denial [Gli84]. Denial-of-Service [Gli84]. Dependability [Lap84]. Dependence [DJ88, ZY87]. dependences [AM86]. Dependencies [Lee87a]. Dependent [AV89, Bas84, MP83, AS88]. Depth [Cha82]. Derivation [Dea81, MI86, Tam80]. Derived [NS83]. Deriving [DB84]. Describing [KP82, BMP86]. Description [DV88, Gie82, SS89, Tak80, Oss89]. Design [AS85, AFS81, ASHP87, AR81, BD87, BG85, BW80, BDRZ85, BR87b, CNW83, CF84, CHP81, Com81, DFS87, Du85, GS78, GCC^{+}87, Gud80, HKMY87, Haw83, Jai87, Kan85, KL84, Kem85, LLS81, LB85, LA81, LM84, ML81, Mac83, MMA87, MY80b, Ng81, NW89, OC82, RH85, RB83, Red83, Rei81, RK84, SvBC87, SSB88, Smo81, SG84, SK85, Sum87, TH86, WL85, Wei87b, YYS81, YC83, YC85, ADWR86, Bag89, BHP84, BZ86, Bra88, CCA86, CMNK88, DYS89, EFRV86, GS88, Gol86, HSN86b, Jal89a, KK88, LHL89, LSK88, Lic86, LL89b, Mea88, ML86, Neu86, PC86a, PC86b, RH86, San89b, San89a, Sch89, SZ88, Tri89, Von88, WF89, YT86]. designer [RS88c]. designing [CL88]. Designs [LG84, BMP86]. detailed [SL89]. Detect [Kem85]. detecting [Lin89, Nat86, Zei89]. Detection [CJS87, Den87, GS80, HR82, JA84, JV83, MSS89, PPR^{+}83, Ram87, SN85, TO80, WB85b, CKST89a, CKST89b, Cid89, ESL88, PS88, WW89]. deterioration [Leu86]. Determination [Ell81]. determining [BMD89]. Developing [Blu87, RM87, WPSK86]. Development [AG83, BDV86, BK89, BR81, Bas85c, Bas80b, Beh83, BDBPLR84, Ber87, Bli81, BP81, Coo84, Cot84, DB85, EZ89, Fic85, FM85, Hay85, How84, JMG88, Kun89, Mat87, NSM85, Ost83, Par80, RMB^{+}81, Rei85, SBB87, WB85a, Won84, AN88a, Boo86, CZ88, DBC88, Dro88, Gaf86, HN86, LL89b, LV89, Orm88, SJ86, WCW89, XH89]. Device [Rav80]. Devices [HS82, Har88a]. DEVS [CZ88]. Diagnosis [FWA87]. Diagnostics [FL80]. diagram [Adl88, War86]. Diagrammatic [Pag83]. Diagrams [JNS83, Was85, BNT86]. Differ [BGL83]. Differences [SW87]. Different [Har80]. difficulties [Sym88]. Difficulty [Tra83]. digital [Aoe89a]. dimensional [RH86]. Direct [BU86, Moi82, BDF^{+}82]. Directed [CHP81, KP82, Rub83, ST87, SK82, DJ88, Eic88, Hay86]. Discarding [SL87]. Discipline [HP80b]. Discrete [BCM87, Mol85, RJ87, HS89, RMM88]. Discrete-Event [BCM87]. Discussion [Goe80]. Disk [Sun87]. Dispatching [KKT80b]. Display [AC82, Bas85a, CRW83]. Dissemination [Top85]. Distribute [SS81]. Distributed [AB85, AH87, ASHP87, Ano82i, Ano83-40, AHY83, ABG84, Ber85a, Bha82, BH83b, BJRE85, Bok81, BCFP87, CG85b, CY83, CL85, Che83, CR87, CA82, CJS87, CGP81, Coo80, CS87b, Dal87, DAT87, FGL82, FR82, GMGK84, GS80, HS87, HR82, Hua85, IS89, JV83, Jai87, KS87, Kes81, KU87, KT87, KM85, KHR86, LW84, LNT85, Len89, Lis82, LS85, MPS85, MINH87, MW82, MIMH85, Nat85, Nes87, NXG85, Ozs85, PPR^{+}83, Per84, PT86, RH85, Ram87, RM87, RS82, Sch89, Sco87, SSL85, SKhL89, Shr81, Sil84, SN85, SS83, Smo81, SA89, SBD^{+}85, Sta85, Svo84, SSP85, Ter87, TB85, TL83, TMB86, Vos80, WL85, Wei87a, Won83, WB85b, YYS81, YC83, YMW^{+}85, YCT^{+}85, ZX87, vBV87, Bag89, Bas88, BYC88, BDHH89, BR89a, CK88a, CK88b, CKM88, CKST89a]. distributed [CKST89b, Cid89, CY89, Dav89, ELZ86, ESL88, FB89, FSAC89, GMK88, GF88, Hać86, Hac89, HD89, ISXC86, JMG88, JKT88, Kes88, Kim89, LTC89, LY86, Lin89, LL89c, MKS89, MD86, MCH88, Nat86, PI86, PV88, RE89, Sin86, TL89, Tri89, TH86, VMBK89, Von88, Wan88, WW89, YM89, YH88, YBL88, YGBC89]. distributing [JKC89, LWL89]. Distribution [BD89, BHJ^{+}87, CNW83, Le 87, MP83, Row88, TTNM89]. Distributions [WR80]. distributive [KH88]. diverse [BEB^{+}86]. Division [EM82]. DMIN [BDMS81]. Do [Mos85]. document [Kim86]. Documentation [ATOK85, HW86a, HW86b, Leh89]. Documenting [BM82]. Does [Mos85]. Domain [AS85, Bar85, CHR82, WC80, WW88, Zei89]. domain-based [WW88]. Domain-Specific [Bar85]. dominance [YW88b]. Double [TL83, Aoe89a]. double-array [Aoe89a]. Double-Loop [TL83]. Draco [Fre87a, Nei84]. Drafting [NXG85]. Drawings [RT81]. Driven [CR83, KKT80b, Zho88]. due [Har89]. Dynamic [BD81, BDMS81, CF81, DB85, Goo81, Jom81, KS87, KM85, Leu86, SKhL89, Sil83b, TS86, WYD87, Won83, YBL88, YD86, Dav89, Hać86, Har89, Zho88]. dynamics [AH89].

Eager [WS85]. Earliest [CC89a]. East [Ano82d]. East82 [Ano81e, Ano82e]. East84 [Ano83e]. Echo [Cha82]. economic [Lev86]. Economics [Boe84]. Economies [BK89]. Eden [ABLN85, PNP88]. Edge [Ano87-59, Ano87-60]. Editing [Wat82]. Editor [Rub83, Kim86, Ano80t, Ano80u, Ano81u, Ano81t, Ano81v, Ano84-29, Cur83, Fai87]. Editor-in-Chief [Ano84-29]. Editorial [Ano80v, Ano80w, Ano81s, Ano81w, Ano82v, Ano83u, Ano83v, Ano84j, Ano84h, Ano84i, Ano84y, Ano85-37, Ano85-38, Ano85-39, Ano86x, Ano86y, Ano86z, Ano87-27, BS81, Bha82, Cha80, Mus83]. editors [TTTI88, Ano87-28, BZ86]. Edmas [AH87]. Education [Fre87b, LPBD87]. Educational [RW87a]. Effect [IR85, SSL85, MBB^{+}89]. Effectiveness [BS87]. Effects [BH80, Bak88, CK88a, TS86]. Efficiency [KB81]. Efficient [BS80, BMOW87, CGW89, CC85b, KBD85, LA86, YW88a, Aoe89a, Aoe89b, Cid89, Kim88, LTC89, RFS88]. Effort [AG83, Par80, LV89]. elaborations [Fea89]. electric [San89b, San89a]. Electrical [Dav80]. Element [BH80, Nta84]. Elements [Ber84, Ber85c, Fre87b]. elimination [HM89]. Embedded [Zav82]. Emerald [BHJ^{+}87]. Empirical [BH83a, LNT85, LC87, PBC85, SBB87, SCD83, SjYTP85, SE84, CCA86, Leh89, LL89c, NS88, TK89]. Empiricist [VW84]. End [Rei87]. Enforce [ZY87]. Engineering [Ano81c, Ano82c, Ano82-40, Ano82-39, Ano83t, Ano84y, Ano87b, Ano87-69, Ard87, BW84, BP83, BDBPLR84, Boe81, Boe84, BP84a, CGP87, DN85, Fre87b, Kee87, LG84, LPBD87, LWSS80, LM87, McG84, Mos85, Red83, Sca84, SLH81a, Sno84, Stu84, Sub85, TPW81, WPSK86, WB85a, Ano86w, BSH86, DL88b, ELHR88, KB88, MBGH89, RB89, Sim86, Ano80a, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano81b, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82b, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano83b, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano84a, Ano85a, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano87a, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l]. Engineering [Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s]. Enhanced [NSM85]. Enhancing [OC82, SY86]. Ensure [Sil83a]. Entity [BL84, JNS83, Ng81, PS85, SZ88, MMT^{+}89]. Entity-Relationship [JNS83, Ng81, PS85, MMT^{+}89]. Environment [BS84, Fer84, Jef87, Lei80, MMF81, MS88, Obe88, Ost83, Pas85, RY85, Rus89, Stu84, Wor87, ABM88, AN88a, BDHH89, CCG^{+}88, CL88, CY89, Ede88, ELHR88, Fis88, HW86a, HW86b, LY86, Moh88, SP86, TMTI86, VDD^{+}88, Win88, WT86, YTT88, ZS86, TS85a]. Environmental [Kar85]. Environments [BS81, Gib87, Goo81, SKW85, Ano86a, BR88, GS88, HN86, HK88a, Mad88]. EQL [JG89, Jay88]. Equational [MS85]. Equations [Foo87, Sch82]. Equivalence [CG85a, JNS83, SS84, LNE89]. equivalent [OW89]. ergodicity [FN89]. Error [CR86, DW81, Mil85, SjYTP85, SL84, Kor88, TK89, Fos80]. error-locating [Kor88]. Error-Prone [SjYTP85]. Errors [Ben87b, BC80, EL85, Gla81, IV85, Kem85, GF88, Zei89]. ESPRESO [Lud83]. Essay [Sta84]. Essential [Chu87, Fre87b]. ESTCA [Fos80]. Estelle [VLC88]. Estelle-C [VLC88]. Estimating [CK85, DW81, Gaf84, Lip86, Sch85]. Estimation [Chu87, Ros85b, Row85, TTNM89]. Euclid [KS86]. Evaluating [CGP87, KH88, SSS82, WB85a, Wey88]. Evaluation [AGCL86, ABM87, ABHM85, Chi85, CP87, DN84, JTS^{+}85, KMS82, KS83, Lap84, Mag87, MS88, MIMH85, Pee81, RH80, SBB87, SL84, Tho85, VW84, YMW^{+}85, Ale88, Bag89, CPRZ89, Fer86, Kim89, KL86, LSK88, MCS88, RB89, SP88]. Event [BCM87, HS89, LL89a, RMM88]. Evolution [Ano83s, Ard87]. Evolving [NS87]. exact [Aky88b]. Example [Bal81, CF80, FM85, JW83, SZ88, TAO89]. examples [BR89b, SP88]. Excellence [Ano83-37]. Exception [BBG^{+}85]. exclusion [JP89]. Executable [SC88, UUD85, EZ89, ZS86]. Execution [AE87, DE82, LC84, Old83, Pla84, Sta82, SK85, GWW89, MBB^{+}89, TMTI86, VMBK89, YT88]. Exercises [BD87]. expansion [KWN88]. Expansive [HM84]. Experience [AS85, CS85a, CGR89, Com81, CA84, DFS87, FSAC89, HKMY87, Kee87, Sum87, WHH80]. Experiences [Bas85c, Mat84, CCRDL88]. Experiment [BR81, Boe81, BGS84, LWSS80, Rom87]. Experimental [JV85, Ost83, Kim89, KL86, LV89]. Experimentation [BSH86]. Experiments [Dun86, Nis83, RP88]. Expert [Doy85, Kus89, NSM85, Sub85, CMNK88, DYS89, RB89]. Explanation [NSM85]. Explicit [CL83, NSM85]. Exponential [LA83, Mil86, Per81, SS84]. express [Agr88]. Expressing [Das85]. Expressions [KS83, MK89, OJ84, ADWR86]. Extended [CF81, GM82, KW85a, MW82, Ozs85, AM86, VS89, WW88]. Extendible [Men82]. Extending [HY87, SP86, Sil83b, Was85, WN89]. extensible [BBG^{+}88]. Extension [BI82, Jaj87, Agr88, War86]. Extensions [Dow86, Jew85]. external [WT89]. extraction [KA88, PM89].

Facilitating [HS87]. facilities

[GR88, YTT88]. Facility [BCL^{+}87, VPH83]. Factors [Nes87, Rei81]. Failure [CR87, Sha85, TR86]. Failure-Handling [CR87]. failures [LM89, SM86]. fake [PC86a, PC86b]. Fall [Ano80r]. False [WB85b]. Family [BH83a, SK82, FW88]. Fast [BD85, Gol81, Reu80, Har88a, KWN88, YGBC89]. Fault [AK83, ABHM85, Avi85, BJRE85, Cri85, GS78, KU87, LLS81, NKT87, Sch85, SGM87, TMB80a, BYC88, JC86, Kim89, LC86, PI86]. fault-tolerance [JC86]. Fault-Tolerant [Avi85, BJRE85, Cri85, KU87, NKT87, Sch85, SGM87, Kim89, LC86, PI86]. Faults [Gaf84, Lip82, Lip86, PT86, Ste86, TTNM89]. FDDI [SJ87]. Feasibility [BR87a].

Features [ABL83, Sta82, ZS86]. Feedback [Bas85c, Per81, HK88b]. FFT [BA83]. Field [JTS^{+}85, Bak88]. File [BH80, BN80b, CF84, Du85, Gud80, LS83, MIMH85, Smi81a, YL80, YSLC85, DC89, Hać86, Hac89, Leu86, WT86, YD86]. Files [CDL84, FWA87, KL84, Sun87, WYD87, JM89, MK89, RS88b]. Filing [MS84]. Finding [Ano85-67, TL89, LTC89]. Finite [Gou84, GL85, CS89a, OP89]. First [CKS85, Rub83, WY85, MBGH89]. First-Fit [CKS85]. Fit [CKS85]. Flex [CG85c]. Flexible [MO87, SJ86, IH86]. floating [Bar89]. floating-point [Bar89]. Flow [BZ80, Che83, HK81, HR81, JA84, LK83, Mac82, MM81, RW85, Sar84, TO80, WO85, YC80a, Adl88, BNT86, CPRZ89, FW88, PS89, War86, YH88]. flowchart [SD89c]. Flowcharts [CV81, Dav80]. FOnetiks [SG81]. Forever [Ber85b]. Foreword [Ano84k, Ano84l, Ano86-27, Ano88, Ber85b, BL86, Bun87, CS87b, GN85, GB85, GB86, How85, JT87, LS85, Mar85, Mos85, Won85]. Form [BI85, Le 87, Ano80-36, Ano81-37, Ano82-37, Ano83-35, Ano84-27, Ano85-64, Ano85-65, Ano85-66, Ano86w, Ano86-28, Ano87-53]. Formal [Ale88, Bar89, BMPP89, Ber87, BBG^{+}85, BP81, CY83, Chi85, DE82, Hay85, Kem85, LR80, cLH89, MG83, Rei81, SL89, SS83, WEHR84, BF86, CPRZ89, Gra88, Hen86, Mea86]. Formalism [Ozs85, CZ88]. Formalizing [BG85]. formally [GS88]. Formatter [Fea82]. Forms [SLTC82, CMNK88]. formula [SD89a]. Fortran [BSP83]. Forward [Mil85, Tay86]. founded [GS88]. Four [Chi85, Har80, Ale88]. FP [KM87, RP88]. FP-Shell [KM87]. Fragmentation [Leu83, Leu86]. Fragmented [YCT^{+}85, CL89]. Fragtypes [Mad88]. Framework [AK83, Gou83, HP80b, CZ88, Ott89]. Free [EHS81, JS87, Mur80]. Freezing [FR82]. Front [Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano84r, Ano84s, Ano84t, Ano84u, Ano84v, Ano84w, Ano84x, Rei87, Ano86-29]. Function [AG83, Bas80b, Beh83, FLS87, GMT85, Sym88, CDY86, FSAC89]. Functional [BN80a, BG85, Hen86, How80, Jal89a, Orm88, Sha82, SVR86, YWH82, HHR89, How86]. Functions [DB84]. Further [Ng81]. Future [Mar85].

G [Bry81, CKM88]. gains [Kam86]. Gambit [BDRZ85]. Gandalf [HN86]. Gannon [LLS81]. General [Cha82, CS84, FO87, MP83, PS85, AS88, AM88, CK88b, Joe89, US86]. general-order-statistics [Joe89]. general-purpose [CK88b]. Generalization [DH84]. Generalized [BBG86, Bec80, HV87, JW83, AI89, DYS89]. Generate [Wes82]. generated [ISXC86]. generating [TTTI88]. Generation [Bre80, CA85, CLP84, Lic86, Pol86, SP88]. Generator [BS82, Gib87, WS81]. Generic [VDD^{+}88]. Genesis [RGP88, BBG^{+}88]. geometric [TTNM89, USV88]. Global [FGL82, Rob81, ZTSN88]. Globally [SA89]. Go [Ber85b]. Good [Lit80, Sta82]. gracefully [MD86]. Gradient [LK87]. Grammar [KP82, Rei81]. Grammars [KP82, LS83, Pap81, Rus89, YM86]. Grant [LM82]. Graph [Che83, Eic88, NS83, ZTSN88]. Graphical [GKS83, Ede88, Fis88, Leh89, OW89]. Graphics [Rei81, FM89]. graphics-assisted [FM89]. Graphs [Cha82, KBB87, Old83, ST87, SK82, YG81, GWW89, YCC89, Yan89]. Gray [CLD82, Fal88]. grid [Oss89]. group [Bra88]. group-select [Bra88]. groups [San89c]. Growth [SS85, YO85, Mil86]. Guaranteed [Ano87-67, Ano87-68, Lei80, LY86]. Guest [Ano80v, Ano80w, Ano81w, Ano84y, BS81, Bha82, Cha80, Mus83, Ano87-28, BZ86, Cur83, Fai87]. guided [BMPP89]. Gutenberg [SVR86].

H [Ale88, Cho89]. Halstead [FZ89, Tra86]. Handle [KW85a]. Handling [BBG^{+}85, CR87, GF88, SD89a]. Hard [Lei80, ZRS87, ZR87, DM89, LY86]. Hard-Real-Time [Lei80, DM89, LY86]. Hardware [GM87, IV85, SM86]. Hardware-Related [IV85]. Hash [DHJS83]. Hashing [CLD82, Men82]. Having [MS87, Sin85]. HDLC [HMS^{+}84]. Heap [GKS83]. Heapsort [WT89]. Heterogeneous [BCL^{+}87, Gib87, LA83, Rus89, BDHH89]. Heuristic [DB84]. Heuristics [Omi89]. Hierarchical [BHN80, KD83, YG81, CZ88, MW88, Neu86, USV88]. hierarchy [San89c]. Higgens [HK88b]. High [Cav85, CLP84, CG85c, Dea81, GKS83, Lei84, ML81, MC84, SPSP87, BWWH88, BF86, GMK88, JM89, JC88, TMTI86]. High-Level [CG85c, Dea81, GKS83, Lei84, ML81, MC84, BWWH88]. Higher [Mat84, Sha89]. higher-level [Sha89]. Highly [CH81]. Hints [Ter87]. historical [TAO89]. HOL [Gaf86]. Homogeneity [Bry81]. homogeneous [ELZ86]. Horizon [MDSW88]. horizontal [AN88a, MDSW88]. Horizontally [Gie82]. Human [Gai85, LWSS80, Rei81, SLH81a, Was85]. Humanized [ATOK85]. Hybrid [EM82]. hyper [TTNM89]. hyper-geometric [TTNM89]. hyperbolic [Fon89]. Hyperexponential [Fer83].

IAI [LPBD87]. IASTED [Dhi86a, Dhi86b]. IBM [IR85, MA87]. IC* [CCG^{+}88]. ICALP [Ano81d]. Icon [CJL^{+}85, CGH83]. Icon-Assisted [CJL^{+}85]. Identification [SW87]. identifiers [SVR86]. Identifying [SjYTP85]. idle [TL89]. IEEE [Ano80-36, Ano80-30, Ano80x, Ano80-31, Ano80y, Ano80-32, Ano80z, Ano80-33, Ano80-27, Ano80-34, Ano80-28, Ano80-35, Ano80-29, Ano81-37, Ano81-31, Ano81x, Ano81-32, Ano81-30, Ano81y, Ano81-33, Ano81z, Ano81-34, Ano81-27, Ano81-35, Ano81-28, Ano81-36, Ano81-29, Ano82-37, Ano82-31, Ano82w, Ano82-32, Ano82-30, Ano82x, Ano82-29, Ano82-33, Ano82y, Ano82-34, Ano82z, Ano82-35, Ano82-27, Ano82-36, Ano82-28, Ano83-35, Ano83-29, Ano83w, Ano83-30, Ano83x, Ano83-31, Ano83y, Ano83-32, Ano83z, Ano83-33, Ano83-27, Ano83-34, Ano83-28, Ano83-38, Ano84-27, Ano84z, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85-64, Ano85-65, Ano85-66, Ano85-52, Ano85-40, Ano85-53, Ano85-41, Ano85-54, Ano85-42, Ano85-55, Ano85-43, Ano85-56, Ano85-44, Ano85-57, Ano85-45, Ano85-58, Ano85-46, Ano85-59, Ano85-47, Ano85-60]. IEEE [Ano85-48, Ano85-61, Ano85-49, Ano85-62, Ano85-50, Ano85-63, Ano85-51, Ano85-67, Ano86-28, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87-53, Ano87-41, Ano87-29, Ano87-42, Ano87-30, Ano87-43, Ano87-31, Ano87-44, Ano87-32, Ano87-45, Ano87-33, Ano87-46, Ano87-34, Ano87-47, Ano87-35, Ano87-48, Ano87-36, Ano87-49, Ano87-37, Ano87-50, Ano87-38, Ano87-51, Ano87-39, Ano87-52, Ano87-40, BMP86, Ano80a, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano81b, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82h, Ano82b, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano83b, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano84a, Ano85a, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano87a]. IEEE [Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Shr85]. If [Bob85]. II [Lee87b, Sie88b]. Image [CJL^{+}85, CYDA88, CIK88, OM88]. images [KA88, USV88]. Imaging [Ano82h]. immediacy [San89c]. Impact [BD82, Rom87, Gaf86, LDF88]. Implement [CL83]. Implementation [Bal81, Bec80, BDF^{+}82, BR87b, DFS87, FL80, Foo87, FM85, GCC^{+}87, KM87, KU87, Moi82, RDU85, Reu82, Whi83, vBGS87, And88, Aoe89b, BU86, BS86, JG89, KCB88, Kim88, Mea88, PNP88, VLC88]. Implementations [KMS82]. Implementing [BJRE85, CG85b, MC84, KL88]. Implication [hSKN89]. Implications [WEHR84, Bra88]. Important [Ano82h]. Improve [BL89b]. Improved [Lit80, Wan96, YLN88]. Improvement [Bas84, BR88, Hac89]. improvement-oriented [BR88]. improvements [Sym88]. Improving [AYS84, BL89a, Car87, TMB80a]. Ina [Ber87, WN89]. INCAS [NHM^{+}87]. Incidence [DAS^{+}87]. Including [ML89]. inclusion [AM86]. Incomplete [KW85a]. Inconsistency [PPR^{+}83, Ram87]. Incorporating [KKT80a]. Incremental [MMF81, PS89]. independence [KL86]. Independent [AE87, KW85c, ML81, SS84, Von88]. Index [Ano80a, Ano81b, Ano82b, Ano83b, Ano84a, Ano85a, Ano87a, Ano86c, Des89]. Indexes [ISR83, Omi89]. Induced [Com81]. Inductive [Bas80a, Kin80]. Industrial [BDBPLR84]. Industry [Hay85]. Inference [CO82, CP87, DM81, Joe89, SD89c, SD89a]. influence [SM86]. Information [Ano80-40, Ano80-37, Ano80-38, Ano80-39, Ano81-38, Ano82-38, Ano83-36, Ano84-28, Ano85-68, Ano85-69, Ano85-70, Ano87-61, Ano87-62, Ano87-59, Ano87-60, Blu87, CR87, CTDS85, HK81, KA88, KW85a, Lee87a, Lee87b, Mac82, MM81, RW85, SL87, Top85, UUD85, YT87, Gra88, Lau88, MK88, NS88, Row88, TI88, WPS86, WPSK86].

Information-Based [CR87].

Information-Theoretic [Lee87a, Lee87b]. inherently [BYC88]. Initialization [ABG84]. Input [VPH83]. Insertion [Pro82]. Inspection [BL89b]. Inspections [WF84, Fag86]. Instances [JS87, Jaj87, JN84]. Institute [Ard87, SKW85]. insularity [Fer86]. Integer [BR87a, SD89c]. integers [SD89c, SD89a, SD89b]. Integrated [DIRY89, GS88, TB85, YNTL88]. Integration [BL84, DH84, DAT87, Mot87, YWH82, LNE89]. Integrity [BDRZ85, LB84, LB85, BF86]. intelligence [HYSY88]. intelligent [CYDA88, ELHR88, Sim86]. Interaction [KB81, Was85]. interactions [EFK89]. Interactive [Bas85a, BDRZ85, BH81, GKS83, LWSS80, MR87, Rei81, VPH83, AN88b, CL88, WPS86, WPSK86, ZTSN88]. Interconnecting [BCL^{+}87]. Interconnection [Ber83, YT87]. Interface [CS84, Obe88, TI88, WCW89, YTT88, SS89]. Interfaces [Bas85a, Chi85, DN85, MRS85, Ale88, KCL88, SC88]. Interfacing [PR85]. Interinput [Le 87]. Intermediate [Mac82, MM81]. Internal [RK84, Smi81b]. International [Ano81d, Ano81c, Ano82i, Ano82c, Ano87b, Ano87-69]. Internet [SP86]. Interorganization [Est87]. Interoutput [Le 87]. interpolating [LHL89]. Interpretation [BD82, FLS87, KP82, IH86, MZ88]. Interpreter [KP82, MM82, Pap81]. Interpreters [KP82]. Interprocess [SVR86, EFK89]. Interrogator [MCF87]. intersection [Row88]. Interval [FyY83]. Intervals [BCFP87, HD89]. Intervention [Abb84]. Introduction [Cur83, Fai87, Sch87a, Shr85]. Intrusion [Den87]. Intrusion-Detection [Den87]. intuition [BMPP89]. intuition-guided [BMPP89]. Invariants [Ell81, MSS89, Tam80]. inversion [YM86].

Investigation [Gou83, LV89, Wan96]. Investment [Ano87-67, Ano87-68]. Invited [Ano83-38]. IOTA [YN85]. irreducible [CIK88]. ISO [Egg81]. Isolation [LP80]. Isomorphisms [KBB87]. Issue [CS87b, HP82, Ano87-54, Ano87-55, Ano87-56, Ano87-57, Ano87-58, BZ86]. Issues [Ano83-40, Bha82, LB85, LA81, SW85, TPW81, THJ87]. Items [Bab81, HH82, Luc85]. Iterations [Tur84]. Iterative [Bas80b, DB82, PS89, SD89b].

J [ZG89]. Jan [CDM89]. JASMIN [LWL89]. Jo [Ber87, WN89]. Job [AV89, NH85, NH86]. Join [Ano83-37, JS87, MS88, Omi89, Des89, PV88, QI88]. Join-Like [JS87]. joins [CL89]. Journal [Ano82h]. JSD [Cam86]. July [Shn81]. justification [Lev86].

K-d [CF81]. KBEmacs [Wat85]. Kernel [GCJ^{+}87, OC82]. Kestrel [SKW85]. Key [NHM^{+}87, WH84, LHL89]. Keys [MS87]. killing [PS88]. Kit [Bev89]. Knee [Nis83]. knot [Cid89]. Knowledge [DK84, Gol86, JS85, KB81, SKW85, SE84, Wat82, YT87, MK88, MW88, Sam89, SSB88]. Knowledge-Based [JS85, SKW85, Gol86, Sam89, SSB88].

Laboratory [WB85a]. lambda [GWW89]. LAN [Gir87, Kes88, SZ88, Sum87]. Language [ABL83, Ber87, CJL^{+}85, CG85c, Coo80, Dea81, HS87, JW83, JTS^{+}85, ML81, MY80a, MC84, PCMP85, Reu82, RS83, Sco87, ST89, Sno83, Tak80, UUD85, VPH83, WO85, ZX87, AK86, BDV86, BHP84, CTYY89, GR88, HB88, JG89, JKC89, JC88, KCL88, KS86, LBY88, MMT^{+}89, Mea86, ML89, SRVO88, Sha89, SY86, TAO89, TTTI88, TMTI86, XH89, ZS86, SS89]. language-based [TMTI86]. language-oriented [TTTI88]. Languages [Ano81d, CLP84, CL83, LF84, Lom85, MC84, OO85, Pag87, HYSY88, OW89, Rus88, SS88, VS89]. Large [FWA87, KL84, LC84, YT87, ZY87, RGP86]. Large-Scale [YT87]. last [Mat86]. layered [XH89]. layout [BNT86]. Learned [BDBPLR84]. Learning [SP88, YW88b, DJ88]. Lee [GS81]. Length [Bab81, HH82, Luc85]. Lessons [BDBPLR84]. Letter [Shn81]. Level [AB85, CLP84, CG85c, Dea81, GKS83, Lei84, ML81, MC84, SPSP87, BWWH88, JC88, Sha89, TMTI86, YW88a]. leveling [PI86]. Levels [Mat84]. Leveraged [Ano87-67, Ano87-68]. Library [HP80a]. Life [Dun86, WEHR84, DBC88, EZ89, TSSP86, YNTL88]. Life-critical [Dun86]. life-cycle [YNTL88]. lifecycle [RS88a, VDD^{+}88]. Like [JS87]. Limitations [Goe85]. Line [Bry81, Lip82, Ste86]. Linear [Ben84, BR87a, Kun82, USV88]. Linearization [TYZ89]. Lines [AG83]. Linguistic [Lis82]. linking [LQ89]. Lisp [FLS87]. List [Ano81a, Ano82a, Ano83a, Ano85-71, Ano87-63, Ano87-64, Ano87-65, Sam82]. Liveness [GL85]. Lloyd [GS81]. Load [CA82, LK87, NXG85, NH85, PI86, SC84, SS84, AS88, ELZ86, NH86, Zho88]. Load-Independent [SS84]. Load-leveling [PI86]. Loads [HFG89]. Loads/Stores [HFG89]. Local [Ber83, CJS87, HFG89, MINH87, RB82b, THJ87, WL85, BDHH89, CL89, CV89, LL89c, WT86, YGBC89]. Locality [KD83]. Locally [AH87]. locating [Kor88]. Location [AE87]. Lock [MFKS85, WH84, CC89b, LHL89]. Locking [MFKS85, PB87, SK82, Sil83a, Tho85, TR89, Eic88, Hać86, Har89, LB88]. locks [HD89]. Locus [SP86]. Logging [Reu80]. Logic [HS82, JW83, Dam88, KL88, MZ88, PM89, SC88, WN89]. Logical [CNW83, Rou82]. Long [Smi81a]. Look [CHR82]. Loop [DB84, Ell81, JS87, TL83]. Loop-Free [JS87]. loops [Kau90, Tow86]. Loosely [Ber85a, Sco87]. Low [AB85]. LR [WS81]. Lucid [SG89].

M [CKM88, Bry81, Cho89]. M/G/1 [Bry81]. M/H [Cho89]. Machine [BDF^{+}82, HD82, ML81, Bra88, QI88]. Machine-Independent [ML81]. Machines [AYS84, Gou84, GL85, SSS82, Aoe89b, CS89a, HYSY88, TL89]. Macro [Reu82, TN80]. Macro-Based [TN80]. macroscopic [LQ89]. Mail [AH87]. Mailbox [Fer84]. mainframe [HL88]. Maintainability [Pee81, Rom87]. Maintainable [BS84]. Maintaining [NS87]. Maintenance [Ano83g, Ben87a, KR87, Man84, NSM85, Sch87a, Sch87b, WEHR84, YC80b, YC85, BEDBT89, DJ88, YNTL88]. Major [TPW81]. Make [Sch82]. Making [BS84, Lom85]. Management [ACM87, Ano84y, BHN80, Ber84, Ber85c, CRW83, CLP84, Coo84, DK84, cLH89, Sil83b, TPW81, Wei87a, BBG^{+}88, BR89b, Gho86, JC88, RB89, TMB86]. manager [KS89, LWL89, SHP88]. Managers [OJ84]. Managing [How84, Sca84, War83, Har88a]. manipulating [MSNT88]. Manipulation [JW83]. Manufacturing [Gho85]. Many [NH85, NH86]. map [TI88, MMS88]. Mapping [AFS81, SS89]. Maps [Kun82, KA88]. Markov [Sie88a, Sie88b]. mass [RE89]. Master [Ard87]. Match [CLD82, Du85, Fal88, YD86]. Matching [AYS84, KO83, MM87, Aoe89b, Rus88]. Mathematical [Gou83, Jom81, MBGH89, RR87]. Mathemtical [KB88]. Matrix [Gol81]. Matrx [DAS^{+}87]. maximal [CIK88]. Maximum [Che83, How84]. Mean [Aky88a, Cho89]. Measure [Geh82, Kha84, KBFL85]. Measurement [IV85, DI89, HZ86, YM89]. measurement-based [DI89]. Measurements [BC80, HW89, Kar90].

Measures [BZ80, SSS82, YC80b, YC85, DL88a, RM88, Wey88]. Measuring [Beh83, Ben87b, CGH83]. Mechanical [BS86]. Mechanism [BH81, HS82, LF84, SC87, GWW89, LHL89, PS88, SF88, Sin86]. mechanisms [CV89, LL89a]. Mechanization [Kim82]. Medical [Ano82h]. Meeting [SKY89]. Membership [Ano87-66]. Memory [AEH84, Bab81, BH81, BD81, BDMS81, HH82, Jom81, LF83, Smi81b, Har88a, Har88b, NW89, QI88, RMM88, YW88a]. Mermaid [YCT^{+}85]. Message [BCM87, CF84, Kes88, Man86]. Message-Based [BCM87]. Messages [Sip81]. Meta [KP82]. metaprogramming [Lev86]. Method [AYS83, AYS84, Bab81, Bec80, Ber87, BL89b, CF81, Cla82, CHP81, GCJ^{+}87, Gol81, Haw83, HH82, KW82, LP80, LK87, Luc85, Oku85, Per84, RC85, Sun87, TR89, YL80, CIK88, Che86, Hir90, Lev86, Mea86, Mea88]. Methodology [BW84, Bas85c, BL84, CC80, Mat87, Pee81, Raj87, RMB^{+}81, RM87, SvBC87, Wei87b, HS88, KH88, SZ88, TR86, WPSK86]. Methods [BP83, CP87, Das85, DTV87, ADWR86, Bar89, BZ86, BF86, Hag89, Kin80, SL89]. Metric [BSP83, Lee87a]. Metrics [Cur83, HK81, JV85, KR87, LC87, LV89, Sha88]. Microarchitecture [Gie82]. microarchitectures [MDSW88]. Microcode [ML81, AN88a]. Microprocessor [Mil87, CV89]. microprocessor-based [CV89]. Microprocessors [CG85c]. Microprogrammed [Gie82]. microprogramming [Dam88]. Migration [Mil87, Smi81a, Hac89, Win88]. MIKE [TL83]. MIMD [JLN83]. Minimal [Dal87, YCC89]. Minimization [HFG89]. MIS [Jef87, Sno84]. Mixed [HS87]. MLS [HY87]. mobile [Fon89]. Model [AFS81, BD82, BL84, BH83b, Che80, CR87, Den87, DK84, Har85, HV87, IMOL84, Jai87, Jew85, Jom81, KW85a, KKT80b, LK87, cLH89, MPS85, PS85, Par80, SS83, Vos80, YWH82, YG81, BMD89, CCG^{+}88, CTT89, CZ88, EZ89, Fon89, Hać86, JKT88, KB88, MZ88, PM89, RS88c, SS89, SHMM89, TK89, TSSP86, TMB86, Wan88, WW88, WCW88, YNTL88]. Modeled [SK82]. Modeling [Ber83, CA85, Dow85, EFRV86, GM85, Hua85, JL87, Kun89, Lau88, MMA87, MB83, NH81, Ozs85, SGM87, WY85, WEHR84, YO85, YWH82, Dow86, HSN86b, LM89, MCS88, OM88, Sch86]. Models [AV89, BN80a, Goe80, Goe85, Jef87, KS83, KKT80a, NSM85, NS83, Pas85, SW78, STE^{+}82, STE^{+}83, TB85, TM85, YO85, AI89, BBG86, DBC88, DH88, HYSY88, HS88, Joe89, Mil86, MIBY88, YSD88]. Modifications [YT87, TS86]. Modified [YC83, CKST89a, CKST89b]. Modula [All80]. Modular [Man86, PCW85, YN85, WE86]. Modularization [Sil84]. Modularized [CH81]. Module [Sha82, Hay86]. Modules [GHM87, LR80, FB89]. Modulo [Sun87]. Molecule [XH89]. Monitoring [DB85, Ger84, Pla84]. Monitors [Kim82, Rav80, WS83]. MSG.84 [BG85]. Multiaccess [Fer84, WL85]. Multiattribute [CF81, CLD82]. Multicomputer [DFS87, NHM^{+}87]. Multidatabase [DH84]. Multidimensional [LB87, RS88b]. Multilanguage [Gib87, Fis88]. Multilevel [DAH^{+}87, LF83, MIBY88, MI89]. Multilinear [YL80]. multilingual [DH88]. Multilist [HK84]. multimicro [HL88]. Multiparty [EFK89]. Multiple [AV89, Mot87, NH85, Rei85, WYD87, Whi83, YMW^{+}85, CGR89, Kau90, NH86, SJ86, Tow86, US86]. Multiprocessor [BA83, DM89, DB82, MB83, ZY87, NW89, QI88]. Multiprogrammed [BH81, Nis83, Har88b]. Multiprograms [CHP81]. Multiproject [BGS84]. Multisystem [WDIY89, CDY86, DIRY89]. Multiversion [EL85, KL86, LM89]. multiversioning [LWL89]. Muse [HOPW87]. Mutation [How82]. Mutual [PPR^{+}83, Ram87, JP89]. MVS [MA87]. MVS/XA [MA87]. Myth [Doy85].

N [CKST89a, Lin89]. N2merik [SG81]. NAG [HP80a]. Name [AB85]. Naming [Win88]. Natarajan [Lin89]. Natural [JTS^{+}85, KCL88]. Navigation [CJL^{+}85]. Nearest [KM83]. Necessary [FN89]. Need [BMOW87, Goo81]. needs [Sim86]. negative [LTC89]. NEST [AE87]. Nested [RS82, Jal89a]. Net [BBG86, HV87, KS83, Ozs85, AI89, BWWH88, DC89, HS89, HSN86b, MCS88, MZ88, MSS89, PM89]. Nets [CR83, DAS^{+}87, Gar85, KBB87, LS87, Mag87, Mol85, NHS83, RH80, Vos80, YC83, BM86, MBB^{+}89, WF89]. Network [BD85, BD89, BD82, JS87, KKT80a, LP80, Per81, RB82b, SSL85, TL83, WL85, YSLC85, BEDBT89, BBG86, BDHH89, CL89, Hać86, JKT88, LL89c, Mea86, RS88c, SA88, WT86]. Networks [BI82, BI85, Ber83, Dav80, Est87, EHS81, JL87, Kam84, Le 87, MY80b, MC81, Oss82, PR85, SS84, TB85, THJ87, AS88, Aky88b, Aky88a, CHKH88, CV89, CG86, EZ89, FN89, Jal89b, LTC89, OP89, PA86, Sch89, YGBC89]. Newton [LHL89]. Next [CA85]. Nexus [Tri89]. Noisy [KO83]. Nominations [Ano83-38]. Non [MFKS85, Sil83a, HSN86a]. Non-Two-Phase [MFKS85, Sil83a]. NON-VON [HSN86a]. Nonclustered [Omi89]. Nondeterminism [Ben84]. nonhomogeneous [Joe89]. nonhomogeneous-Poisson-process [Joe89]. Noninterference [HY87]. Nonlinear [RW87b, TYZ89]. Nonprocedural [PP83]. Nonqueueing [Fer84]. Nonsensitive [Sin83].

Nontraversible [Che87]. Nonunique [Sin85]. Normalization [Kam84]. Norton [BI82]. Notation [Pag83]. Note [Ano81t, Ano84-29, Ano87-28, Bas80a, Dav87, Gli84]. Notes [Gut80]. Notice [Ano80t, Ano80u, Ano81u, Ano81v]. notion [JKC89]. NP [Mag87]. NP-Complete [Mag87]. null [KCL88]. Number [CGP81, Gaf84, Lip82, Ste86, Bar89, BMD89, Lip86, RM88, TTNM89]. Numerical [ASP82, Gho89].

Object [AH87, Boo86, Sno83, HB89, Jal89a, MK88]. Object-Oriented [AH87, Sno83, Boo86, HB89, Jal89a, MK88]. Objects [BJRE85, Pag83, Jal89a, Jal89b, KB88, KCB88, Kim86, PNP88]. Observations [PBC85]. Obsolete [SL87]. obtained [Woo86]. Occurring [BR87a]. Off [Bry81]. Off-Line [Bry81]. Office [CC85a, SLTC82, SKY89]. Omega [WO85]. On-Line [Bry81]. One [Ano87-67, Ano87-68]. online [DM89]. Only [CG85b, Wei87a]. Open [OJ84, Per81, FN89, PA86]. Operating [ASHP87, BC80, Gli84, IR85, MA87, Pas85, RB82b, SG84, TL83, Bev89, FSAC89, KS89, Sam89, Tri89]. Operation [JS87, Lap84, Bra88]. Operational [Ben87b, BS84, MD86, Smo81, Zav82]. Operations [BGL83, Cha82, KBD85, KCB88, Gho89]. Operator [CS84]. operators [OW89]. Optimal [Bok81, BDMS81, CL85, GS78, HK84, ISXC86, ISR83, KK83, LLS81, NH85, NH86, Nic89, Pro82, SKhL89, WT86, BMD89, CY89]. Optimality [CKS85]. Optimistic [Car87, Bas88, LWL89]. Optimization [AHY83, Ber83, BMOW87, CKM88, CG85a, CS89b, CGP81, LW84, RH85, Sam82, SC84, Arm89, Kus89]. Optimizing [CL89, FK80, FLS87, KGM88, PV88, RP88].

Optimum [Sip81]. Order

[Tra83, Joe89, Mil86]. Ordering [Mck85]. Organization [CF81]. organizing [KB88]. Oriented [AH87, Che80, LK83, Reu80, Sno83, WR80, vBV87, BR88, Boo86, HB89, Jal89a, MK88, TTTI88]. orphan [HM89, PS88]. OSI [Kar90, HW89]. OSI-style [Kar90, HW89]. Other [CKM88]. Output [VPH83]. overflow [YD86]. Overhead [KKT80a, YL80, HW89, Kar90]. Overview [Row85, Cam86, Fis88, Tri89].

P [CKM88]. page [Ano84c, LWL89]. Paged [Jom81]. Paging [CH81]. Palindromic [MS87]. Panel [Ano85-35, Ano85-36]. Paper [LLS81, KA88]. paper-based [KA88]. Papers [Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano81g, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano82t, Ano82p, Ano82q, Ano82r, Ano82s, Ano82-39, Ano83s, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano84d, Ano84e, Ano84f, Ano85-35, Ano85-36, Ano85y, Ano85z, Ano85-27, Ano85-30, Ano85-32, Ano85-33, Ano85-34, Ano87t, Ano84g, Ano85v, Ano85w, Ano85x, Ano85-28, Ano85-29, Ano85-31, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano87u, Ano87v]. Paradigm [Blu87, KB81, RB83, Sch89]. Paradigms [Bob85]. Parallel [Cha82, CS85b, CK85, JLN83, KW85c, LA81, RB83, RMM88, SS81, TS85b, CCG^{+}88, Fea89, May89, MI86, MZ88, NTT88, SRVO88, XH89, YM89]. Parameterized [Gog84]. Parameters [SSL85]. PARLOG [Cla88]. Parser [WS81]. Parsers [AYS83]. Parsing [CK85, Dea81, Rus88]. Part [Lee87a, Lee87b]. Partial [CVGM80, CLD82, Du85, Fal88, Pol86, YD86]. Participation [Ano82u]. Partition [Hua85, RC85, YGBC89]. partitioned [HB89]. Partitioning [IS89, ISXC86, JKC89, Nic89, RS88b, WDIY89]. partly [MK89]. Pascal [CVGM80, FL80, GM82, Egg81, FK80, FZ89, GKS83, Pas85, SJ83]. PASM [TS85b]. passing [Man86]. Past [Den80]. Path [Chu87, OJ84, PM87, Rou82, WHH80, CPRZ89]. Paths [Ben84, Che87, LTC89]. Pattern [AYS84, Aoe89b, Rus88]. Patterns [Smi81a]. PECAN [Rei85]. PELAS [Kor88]. PELAS-program [Kor88]. percent [Mat86]. Perfect [CLD82, DHJS83]. Performance [ACM87, AA89, BEDBT89, BHN80, Bas84, BA83, BW80, Car87, CDL84, CH81, CGH83, Des89, DB82, GMT85, Gar85, GM85, HD82, HV87, HB89, KMS82, KS89, LNT85, MM82, Mag87, Mur80, NTT88, NH81, Pas85, PB87, RE89, RH80, RW87b, ST87, Sch85, SSL85, SSS82, Sun87, TB85, Tho85, TR89, YM89, BL89a, Bag89, CHKH88, Dow86, Fer86, Hac89, HL88, HSN86a, Kun86, MIBY88, Sam89, Woo86]. performing [IH86]. periodic [BEDBT89]. Persistent [Gla81]. Person [BL89b]. Personal [Sum87]. Perspective [Bal85]. Perspectives [RB82a]. Perturbation [Zei89]. Perturbations [Zei83]. pessimistic [JM89]. Petri [MCS88, AI89, BBG86, BWWH88, BM86, CR83, DAS^{+}87, DC89, Gar85, HS89, HV87, HSN86b, KBB87, KS83, LS87, Mag87, MBB^{+}89, Mol85, MSS89, NHS83, Ozs85, PM89, RH80, WF89, YC83]. Petri-net-based [MCS88]. Phase [MFKS85, PB87, Sil83a, Har89, LB88]. phenomena [BBG86]. PHILAN [Kes88]. Philosophies [Har80]. philosophy [Kim88]. Physical [WYD87, LSK88, RS88c]. PICQUERY [JC88]. Pictorial [CF80, JC88, RFS88]. Pipelined [IS89, MS88, MCS88]. PL [Sar84]. PL/I [Sar84]. Placement [WYD87]. Planning [Coo84]. PODS [BEB^{+}86]. Point [KM83, Bar89, CTT89, Sym88]. Pointers [Lom85, ML86]. Points [Beh83]. Poisson [Joe89]. policies [MBB^{+}89]. Policy [FyY83]. polynomial [LHL89]. Port [MY80a, RSBV86]. port-based [RSBV86]. Post [Sha85]. Post-Failure [Sha85]. POSTGRES [SHP88]. Practical [AYS83, CS85a]. Practice [BP83, BMP86, Hag89]. Practices [Sno84, CCA86]. Precedence [AYS83]. Predicate [Vos80, MZ88]. predicate-transition [MZ88]. Predicate/Transition [Vos80]. Predicate/Transition-Nets [Vos80]. Predictability [DI89]. Predicting [War83]. Prediction [AG83, TK89]. predictions [AGCL86]. predictors [TT86]. Prefix [PM87]. Preliminary [Ano83-39]. Prescribed [Mur80]. presence [PT86]. Present [Den80]. Presentation [BC80, Mat84]. Preserving [CH88]. Press [Ano81-30, Ano82-30, Ano82-29]. Pretty [Gai85, Rub83]. Prevention [KS87]. PRIM [Gol81]. Primitives [BDHH89, Nat85, WS83]. Principles [Com81, BR89b, MBGH89, Smi88]. Printing [Rub83]. Priority [KKT80b, SN85, CKST89a, CKST89b]. Priority-Driven [KKT80b]. Privilege [RSBV86]. Probabilistic [EM82, PS87]. Probabilities [MP83]. probe [CKST89a, CKST89b, OM88, Sar84]. Probes [Pro82]. Problem [Che83, Du85, JNS83, KM83, KO83, Ros85a, Ver82, BKL89, KCL88, LC86, MKNS89, May89]. Problems [Ber83, HR81, Mag87, Sin85, ABM88, Har88a, YW88b]. Procedural [Cla82]. Procedure [BCL^{+}87, LH83a, LG85, HZ86, PS88, PM89]. Procedures [CC85a, HMS^{+}84, SLTC82, MCH88, Ten88]. Process [Bag89, BM86, Lud83, MY80b, RMB^{+}81, SS85, Adl88, CHKH88, DI89, Hac89, Joe89, LL89b, Moh88, NW89, PC86a, PC86b, SVR86, TR86, WCW89]. Process-translatable [BM86]. Processes [DB82, Geh84, Lom80, MC81, Oss82, RK83, Rus80, SL84, WS83, Wes82, Kim88]. Processing [CJL^{+}85, CL85, CS85b, GM87, Har80, JLN83, Kes81, MS88, Omi89, Per84, PCMP85, Sam82, SC87, SLTC82, hSKN89, TB85, TS85b, Won83, YCT^{+}85, Kes88, LB88, MD86, MCS88, NTT88, OM88, YGBC89]. Processor [Bok81, Fer83, NH85, CGR89, Kau90, NH86, PT86, Tow86, US86]. Processors [KW85c, NH81, FB89, Nic89]. Product [BI85, CDL84, Le 87, Sun87, Mun88]. Product-Form [BI85, Le 87]. Production [Rei87, TMTI86]. Productivity [Beh83, BL89b, How84]. Products [BR87b]. Professionals [Ano80s, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, CLP84]. profile [LQ89]. Prog [BL89b]. Program [AGPT85, Ano81s, Ano83f, Ano83g, Ano85b, AC82, AR81, BBG^{+}85, BM84, BP81, Cha80, Che84, Che87, CLP84, CS87a, CH81, Com81, Cot84, DE82, DM81, Fea82, GMT85, How80, JS85, KB81, Kin80, LK83, LF83, LPBD87, LR87, LS83, Mac82, Mac83, MM81, ML86, NS83, Old83, Pag87, PBC85, Rei85, Tai80, Ten88, Wat82, Wat88, Wei84, Wes82, WO80, WC80, Zei83, Ano86b, Arm89, BMPP89, CCRDL88, Dro88, Gol86, Har88b, How86, JKC89, Kor88, KHR86, Kun86, LL89a, Lic86, LQ89, RH86, TK89, Ard87]. Programmer [Kim88, Wat82, Wat85]. Programmer-transparent [Kim88]. Programming [Abb84, Ano81d, Bal85, BS81, Bar85, BN80a, Bob85, Coo80, Cri85, Dav87, Gog84, Goo81, HS87, HP80b, Jon84, Kes81, Lom85, MY80a, MMF81, Obe88, RGP86, RW87b, RS83, SPSP87, SE84, TS85a, Tur84, VW84, WS83, WO85, YN85, ZX87, BMPP89, CCG^{+}88, CL88, CH88, Ede88, Gol86, HHR89, Hen86, KL86, Mad88, MMT^{+}89, ML89, ML86, SY86]. Programs [ASP82, BM82, Bas85d, Bas80b, Bli81, Com81, CGH83, Cri84, Ell81, Haa81, HO83, KS83, LC84, Lis82, Nat85, NHS83, Pro82, PP83, RM87, Sar84, Sch85, Sco87, Sha85, WHH80, YG81, BYC88, FM89, FZ89, DV88, FGM89, GF88, ISXC86, JKC89, KSB89, Kau90, KL88, Leh89, MSNT88, MO90, MZ88, Nic89, PM89, SD89c, SD89a, SD89b, SRVO88, Tow86, TMTI86, VMBK89, XH89, YM89]. Progress [Gou84, CS89a]. Project [Ano84y, BP84a, BS84, Kee87, cLH89, Ost83, TPW81, AH89, BMP86, BR88, BEB^{+}86, BR89b, HK88a, NHM^{+}87]. projection [Che86, Hir90]. Projections [LS84]. Projects [BSP83, How84, Sca84, Wor87]. Prolog [Bob85, CGW89, Sub85]. prologue [BZ86]. Promoting [Mat84]. Prone [SjYTP85]. Proof [MS85, PM89, ESL88]. Proofs [MC81, Oss82]. Properties [NS83, SJ87, AM86, HB89, JM86, USV88]. Property [Kam84]. Proposals [Ano85-35, Ano85-36]. PROSPEC [CL88]. PROTEAN [BWWH88]. Protection [HS82, San89c]. Protocol [Har85, Hol87, LS84, Lam88b, MCF87, RDU85, SvBC87, SJ87, And88, Che86, Hir90, KWN88, LTC89, LL88, Mea88, SZ88, SC88, SL89, ZTSN88, Lam88a]. Protocols [ABM87, Gar85, GM85, HR82, JV83, JL87, LB87, MFKS85, SK82, Sil83a, STE^{+}82, STE^{+}83, ZR87, vBV87, vBGS87, Bas88, BWWH88, BS86, CL88, LA86, SA88, VLC88]. Prototype [SBD^{+}85]. Prototyping [BGS84, BM86, CCRDL88, DV88, Hen86, HB88, JMG88, LBY88, SS88, Wan88]. PROUST [JS85]. prover [SD89b]. PROVIDE [Moh88]. Providing [KCL88, RS88a, Kes88]. Proving [GM82, GL85, Hol87]. PSQL [RFS88]. Psychology [Cou83]. Public [Com81]. Publication [Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g]. Publications [Ano80-30, Ano80-31, Ano80-32, Ano80-33, Ano80-34, Ano80-35, Ano81-31, Ano81-32, Ano81-33, Ano81-34, Ano81-35, Ano81-36, Ano82-31, Ano82-32, Ano82-33, Ano82-34, Ano82-35, Ano82-36, Ano83-29, Ano83-30, Ano83-31, Ano83-32, Ano83-33, Ano83-34, Ano85-52, Ano85-53, Ano85-54, Ano85-55, Ano85-56, Ano85-57, Ano85-58, Ano85-59, Ano85-60, Ano85-61, Ano85-62, Ano85-63, Ano87-41, Ano87-42, Ano87-43, Ano87-44, Ano87-45, Ano87-46, Ano87-47, Ano87-48, Ano87-49, Ano87-50, Ano87-51, Ano87-52]. published [Lip86]. Purpose [CS84, CK88b].

quadtrees [USV88]. Qualified [GMT85, HR81]. Quality [BP84b, Chu87, Oku85, SM85, Dhi86a, Dhi86b, KCL88]. quantified [SD89a]. Quantitatively [BR81]. Queries [AHY83, CG85a, CLD82, CL85, CS89b, Har80, LW84, Per84, hSKN89, Won83, Agr88, CKM88, Fal88, IH86, MCS88, PV88]. Query [BHN80, CF80, JW83, MS88, OO85, YCT^{+}85, JC88, OM88, TAO89, VS89, YGBC89]. Query-by-Example [JW83]. Query-by-Pictorial-Example [CF80]. Quest [Ano83-37]. Question [All80, Bob85, OC82]. Question-Answering [OC82]. Queue [MP83, Per81, Woo86]. Queueing [BI82, BI85, Bry81, HP82, Kam84, LA83, NKT87, SS84, AS88, Aky88a, BBG86, CG86, FN89, Hać86, JKT88, OP89]. Queues [LP80, PA86]. Queuing [KKT80a, KKT80b, Aky88b]. Quicksort [Dav88, BL89a]. quorum [AA89]. quorum-consensus [AA89].

R [Ano83s]. races [KSB89, MO90]. Raid [BR89a]. Rajdoot [PS88]. Ramifications [SS85]. ramming [BL89b]. Random [CDL84, DN84, SS85, WYD87, Nic89]. Randomized [CC85b, Per85]. randomly [ISXC86]. range [Fal88]. rapid [BM86, Hen86, SS88]. rational [PC86a, PC86b]. Rayleigh [Par80]. Read [CG85b, Wei87a, Len89]. Read-Only [CG85b, Wei87a]. read-single [Len89]. readability [Ten88]. Real [Abb84, AK83, Das85, Haa81, HKMY87, Hua85, JL87, JV85, KS86, Lei80, Pla84, SG89, TSSP86, VPH83, War83, ZRS87, ZR87, AK86, BL86, DM89, FGM89, JM86, Kim89, LY86, LBY88, San89b, San89a]. Real-Time [Abb84, AK83, Das85, Haa81, Hua85, JL87, JV85, KS86, Pla84, VPH83, War83, ZRS87, ZR87, SG89, TSSP86, AK86, BL86, FGM89, JM86, Kim89, LBY88, San89b, San89a]. realizable [HB88]. Reasoning [AN88b, Doy85, PS87, Sha89, DJ88]. Recapture [DW81]. Reclamation [WS85]. Recognizing [San89c]. recommended [BMP86]. Reconfiguration [Sha85]. Recoverability [Shr81]. recovering [Kim88]. Recovery [ABG84, CTT89, GM85, KT87, Mil85, Reu80, SL84, SS83, CR86, KGM88, Tay86]. rectangles [CIK88]. Recursion [CS87a, Sch82]. Recursive [FLS87, TR89, Agr88, TYZ89]. redefined [May89]. Reducing [AYS83]. Reduction [AGPT85, Hag89, Ott89]. Redundancy [MIMH85, TMB80a, TMB80b, Man86]. Redundant [Won83]. Referees [Ano81a, Ano82a, Ano83a, Ano80b, Ano85-71, Ano87-63, Ano87-64, Ano87-65]. Reference [KD83, Smi81a]. Refinement [KB81, Raj87, Jal89a]. Reflexive [DAS^{+}87]. Regeneration [PNP88]. Register [HFG89]. registration [Ano86w]. reimplementation [Wat88]. Rejection [AV89]. Related [IV85, THJ87]. Relational [AFS81, Bec80, CG85a, CS89b, Jaj87, KW85a, Lee87a, Lee87b, MS88, WY84, YCT^{+}85, Agr88, Bra88, CCRDL88, Gho86, Gho89, IH86, JN84, LL88, ML89, OW89, Wan96, YLN88]. relations [CY89, YW88b]. Relationship [BL84, JNS83, Ng81, PS85, LV89, MMT^{+}89]. Release [KK83]. Reliability [Bha82, Che80, Goe80, GB85, Goe85, IMOL84, IR85, Jew85, Lit80, MA87, RH85, RB82a, Ros85b, Ros85a, ST87, SW78, SGM87, Sie88a, Sie88b, SS85, TM85, YO85, AGCL86, Bak88, BL86, CDM86, CDM89, Dhi86a, Dhi86b, Dun86, GB86, HH86, Joe89, KHR86, LDF88, Mil86, Sch86, SY86, SM86, WW88]. Reliable [BD85, RMB^{+}81, Kes88, KS86, TMTI86, TH86, WPS86]. Reliably [Ber85a]. Rematerialization [Won83]. Remote [AE87, BCL^{+}87, LG85, PS88]. Rendezvous [GR88]. Reordering [HP82]. Repetitive [FLS87]. Replicated [FWA87, SL87, AA89, DC89, JM89, PNP88]. replication [Hac89]. reply [Bab82, Hir90, SLH81b, ZG89]. represent [War86]. Representation [DAS^{+}87, DK84, Mac82, MY80b, MM81, WY84, YC83, YT87, HS89, MK88]. Representative [JN84, Jaj87]. Representing [Bab81, HH82, Luc85, MRS85, YG81, MSNT88]. request [CDY86]. Required [Nta84]. Requirements [CR83, Hen80, MMA87, Smo81, Zav82, ZRS87, HM83, NS88]. Research [Ost83, SKW85, VW84]. resequencing [Cho89]. residual [TTNM89]. Resilient [Jal89b, MW82, Svo84]. Resistance [Shr81]. resolution [CKST89a, CKST89b, ESL88]. Resource [Lei84, NH81, OJ84, Ram85, Sil83b, Sin85, Sum87, ZRS87, DI89, SP88]. Resources [Mur80, SKhL89]. Response [Lei80, CK88a, LY86]. responses [KCL88]. responsive [Smi88]. Restoration [Rus80]. Restructuring [LF83, PBC85, Har88b]. Results [ASP82, BP83, CC89a, Fer83, LM87, SS84, TMB80b, NS88]. retargetable [MDSW88]. Retraining [BDBPLR84]. retries [US86]. Retrieval [CTDS85, Du85, JTS^{+}85, TI88, YD86]. Return [Ano87-67, Ano87-68, Kun86]. Reusability [BM84, Che84, Jon84, Ker84, RGP88]. Reusable [HM84, LG84, LM84, Mat84, Nei84]. Reuse [Pol86, Sta84]. Revealing [WO80]. reverse [CTT89]. Review [ABLN85]. Reviews [WF84]. Revisited [Fre87b, SCD83, Dav88]. revocation [RSBV86]. revolutionary [BKHW89]. Rights [LM82]. Rigorous [Cot84, Cri85]. RIM [DAS^{+}87]. Ring [SJ87, Har85]. Risk [Fry85]. Robust [Cri84]. Role [AS85]. Roles [MRS85, SK85]. Rollback [KT87, US86]. Rollback-Recovery [KT87]. Round [MS87]. Routing [ATR82, AV89, YBL88]. RPC [Gib87]. RT [AK86]. RT-ASLAN [AK86]. RTAG [And88]. RUBRIC [CTDS85]. Rule [CTDS85, Hua80, Ros85a, SHP88, TS86]. Rule-Based [CTDS85]. Rules [DM81, Kim88, SD89c, SD89a, TYZ89]. Run [BD82, FL80]. Run-Time [BD82, FL80]. Runtime [Egg81].

SADAT [VGV80]. Safe [Lom85, Phi84]. Safety [JM86, LH83b, LS87, BL86, OW89]. Saguaro [ASHP87]. Salient [ZS86]. Sampled [FyY83]. Sampling [DW81]. Sanf [SG81]. SARA [EFRV86]. SAT [HY87]. scalars [ML89]. Scale [BK89, Boe81, YT87, AI89]. Scatter [Hol87]. Schedule [MINH87]. Scheduler [SKY89]. Schedulers [Lei84, IKK88, Woo86]. Schedules [Abb84]. Scheduling [CC89a, Smi81b, Sta85, TS85b, ZRS87, CK88a, CK88b, DM89, LC86, NW89]. Schema [BD82, BL84, RK84, Rou82, LNE89, War86]. Schemas [AFS81, CNW83, Lee87b]. Schemata [Wes82]. Scheme [CLD82, DHJS83, EM82, JLN83, Kim82, LB84, MRS85, MW82, Reu80, ZY87, Dav89, Lin89, Nat86, TH86, Wan96, WF89, YLN88]. Schemes [Jaj87, YMW^{+}85, JN84, Kim89, KH88]. Science [AG83, Ano82-40, Cou83, KW85b, SCD83, RM88, SHMM89]. Scientific [Bas85b, SW85]. Scope [CL83, FLS87].

Screen [RS83]. Scribe [Oss89]. SE [Ano80a, Ano81b, Ano82b, Ano83b, Ano85a, Ano87a]. SE-11 [Ano85a]. SE-13 [Ano87a]. SE-18 [Ano82b]. SE-6 [Ano80a]. SE-7 [Ano81b]. SE-9 [Ano83b]. Search [Man84, WY85, Aoe89a, MK89, YW88b]. Searching [Hol87, SS81]. Second [Ano82-39]. Secondary [Leu83, Har88a]. Secrets [MM87]. Section [Mus83, Sch87a]. Secuity [MCF87]. Secure [CC89b, GCC^{+}87, Gud80, Mac83, SG84, GCJ^{+}87]. Security [BH80, DAH^{+}87, GCJ^{+}87, Nes87, OC82, SF88, TS86]. SEE [Stu84]. SEES [RY85]. Segmentation [BH80]. select [Bra88]. Selecting [RW85]. Selection [Chu87, HK84, ISR83, RW87a, CTT89, CPRZ89]. Selective [AR81]. Selectors [Lei84]. Semantic [FKM85, Haw83, HK88b, TI88]. Semantics [AGPT85, CG85a, Jay88, MBB^{+}89]. semaphores [JP89]. Semiautomatic [VLC88, vBGS87]. Semijoin [MINH87]. Sensitive [Fos80, Jef87]. Separate [CVGM80, LF84]. Sequences [BR87b, MS87]. Sequential [Geh84, IS89, MK89]. Series [Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g]. Server [AV89, BN80b, CF84, LA83]. Servers [Fer83, BEDBT89, Fon89, WT86]. Service [Gli84, MP83, Bak88, Kes88, RE89, SC88]. Servicing [WR80]. Session [Ano85-35, Ano85-36, Wat85]. Set [ASP82, FyY83, ISR83, Obe88, OW89, Row88, SD89a, WCW89]. Sets [Den80, How82]. several [YSD88]. shadow [KGM88]. Shapiro [LLS81]. share [Woo86]. Shared [NH81, NW89, QI88, RMM88]. shared-memory [QI88]. Shared-Resource [NH81]. Sharing [Fer83, Sum87, DIRY89, ELZ86, WDIY89, YD86]. Sharing-Results [Fer83]. Shell [KM87]. shipping [CDY86]. Shop [Sno84]. short [HD89]. short-term [HD89]. Shortage [McG84]. Shortest [Bok81, MINH87, LTC89]. Shrivastava [GS81]. significance [TI88]. SIMD [CS85b, SSS82]. Simple [CC85b, LF84, SF88]. simplicity [FSAC89]. simplification [WF89]. simplified [Ott89]. Simplifld [SG81]. Simulation [BCM87, LH83a, Mil87, AH89, EFRV86, GS88, MCH88, RMM88, Zho88]. simulations [HS89]. Single [Bas88, PB87, WH84, LHL89, Len89]. Single-Key-Lock [WH84, LHL89]. Single-Site [PB87, Bas88]. Site [ABG84, PB87, Bas88, CY89]. Size [Sip81]. sizes [Row88]. SLAN [BHP84]. SLAN-4 [BHP84]. Sliced [KKT80b]. Slicing [LR87, Wei84]. Small [Ano83s, Boe81, BS84, Com81, SJ83, Sno84, Gra88]. Small-Scale [Boe81]. Social [Sca84]. Society [Ano81-30, Ano82-30, Ano82-29, Ano87w, Ano87x, Ano87y, Ano87z, Ano80-37, Ano80-38, Ano80-39, Ano80-30, Ano80x, Ano80-31, Ano80y, Ano80-32, Ano80z, Ano80-33, Ano80-27, Ano80-34, Ano80-28, Ano80-35, Ano80-29, Ano81-31, Ano81x, Ano81-32, Ano81y, Ano81-33, Ano81z, Ano81-34, Ano81-27, Ano81-35, Ano81-28, Ano81-36, Ano81-29, Ano82-31, Ano82w, Ano82-32, Ano82x, Ano82-33, Ano82y, Ano82-34, Ano82z, Ano82-35, Ano82-27, Ano82-36, Ano82-28, Ano83-29, Ano83w, Ano83-30, Ano83x, Ano83-31, Ano83y, Ano83-32, Ano83z, Ano83-33, Ano83-27, Ano83-34, Ano83-28, Ano84z, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85-52, Ano85-40, Ano85-53, Ano85-41, Ano85-54, Ano85-42, Ano85-55, Ano85-43, Ano85-56, Ano85-44, Ano85-57, Ano85-45, Ano85-58, Ano85-46, Ano85-59, Ano85-47, Ano85-60, Ano85-48, Ano85-61]. Society [Ano85-49, Ano85-62, Ano85-50, Ano85-63, Ano85-51, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87-41, Ano87-29, Ano87-42, Ano87-30, Ano87-43, Ano87-31, Ano87-44, Ano87-32, Ano87-45, Ano87-33, Ano87-46, Ano87-34, Ano87-47, Ano87-35, Ano87-48, Ano87-36, Ano87-49, Ano87-37, Ano87-50, Ano87-38, Ano87-51, Ano87-39, Ano87-52, Ano87-40]. SODOS [HW86a, HW86b]. Sofiware [IS89]. Software [AS85, AG83, AK83, ABHM85, Ano80a, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano81c, Ano81b, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82c, Ano82b, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82-39, Ano83g, Ano83b, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano84y, Ano84a, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85a, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87a, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ard87, Avi85, BK89, BR81, BW84, BS87, BP83, BDBPLR84, Ben87a, BD87]. Software [Ber84, Ber85c, Boe81, Boe84, BW80, BP84a, BS84, BP84b, BKHW89, BR87b, CGP87, CRW83, Cav85, CCRDL88, Che80, Coo84, Cou83, Cur83, DB85, Dow85, DN85, DW81, EL85, Fic85, Fon89, Fre87a, Fre87b, GS78, Gla81, Goe80, GB85, Goe85, GM87, HN86, Hag89, Hay85, HH86, Hen80, HK81, HM84, How84, Hua85, IMOL84, IV85, JV85, Jew85, KR87, Kar85, Kee87, Ker84, KK83, KW85b, LG84, Lap84, LPBD87, LWSS80, LLS81, LM87, LH83b, LDF88, LC87, Lit80, cLH89, Lud83, Mat84, MMA87, McG84, MY80b, MO87, Mos85, NS87, Nei84, Oku85, Ost83, Par80, Pee81, PM87, Pro82, RMB^{+}81, RB82a, RW85, Red83, Rom87, Ros85b, Ros85a, Sca84, SW78, Sch87a, Sch87b, Sch86, SGM87, SBB87, SCD83, SjYTP85, SS85, SKW85, SM85, Sno84, Sta84]. Software [SW87, Stu84, Sub85, TO80, TMB80a, TPW81, THJ87, TM85, War83, WPSK86, WB85a, Wei87b, WEHR84, Won84, Wor87, YO85, YC80b, YYS81, YC83, YC85, YT87, YTT88, AGCL86, AH89, ABM88, ADWR86, Bak88, BBG86, BMP86, BSH86, BR88, BHP84, BZ86, BEB^{+}86, BF86, BP88, BR89b, BKL89, BMD89, CCA86, CDM86, CDM89, DBC88, Dow86, Dun86, Fag86, Gaf86, GB86, HZ86, HW86a, HW86b, HK88a, Joe89, KK88, LL89b, LV89, LM89, LBY88, Mil86, MBGH89, Mun88, Pol86, RM88, RS88a, RB89, San89a, Sch89, SJ86, SP88, Sha89, SHMM89, Sim86, Smi88, SY86, SM86, TTNM89, TR86, TSSP86, TH86, WW88, Wey86, Wey88, WCW89, YT86, YNTL88, YSD88, ZG89, RY85, Stu84, San89b, Ano85i, Shr85]. Solution [TR89, Hać86]. Solutions [Foo87, Sin85, JP89]. solving [MKNS89]. Soma [Kes81]. Some [ASP82, AEH84, CC89a, Dow85, FGM89, HSN86b, LLS81, Mat84, NS83, PBC85, RR87, SD89c, SS84, TMB80b, WB85a, YC80b, vBV87, BBG86, Dow86, Har88a, NS88, Nta88]. Sorting [AEH84, Ver82, LL89c]. Source [AG83]. Space [Bok81, BMOW87, MR87]. Space-Efficient [BMOW87]. Spanner [ABM87]. Spanning [Dal87]. spatial [MK88, OM88]. SPD [ATOK85]. Special [Cha80, CS87b, Mus83, Sch87a, BZ86]. Specific [Bar85]. Specification [ABM87, Bas85a, BS80, Ber87, BBG^{+}85, CVGM80, CY83, Chi85, Cla82, Fea82, Gar85, Hay85, Hay86, HMS^{+}84, LR80, LM84, Lud83, Moi82, MS84, Mus80, RK83, SLTC82, STE^{+}82, STE^{+}83, UUD85, Was85, WE86, Zav82, vBV87, ABM88, Ale88, AK86, BHP84, BWWH88, Gra88, HM83, Hen86, Oss89, SG89, SS88, Wan88, ZS86]. Specifications [Bas80b, BG85, Foo87, Kem85, MS85, MG83, Phi84, PP83, BU86, Fea89, HS88, Jal89c, SA88, SC88]. Specified [Bli81]. Specifying [BGS84, Cla82, CG85c, Hen80, DV88, ST89]. Speedup [CK85]. SQL [CG85a]. Stability [Sta85, YC80b, YC85, CK88a]. staffing [AH89]. Stage [BD89]. Standard [Egg81]. Standards [Ano82-39, BP84a, BMP86]. Star [CL85, YSLC85, CA84]. State [AV89, Ano83-40, Gou84, GL85, Jon84, MP83, MR87, Rus80, Sch87b, STE^{+}82, STE^{+}83, Was85, CS89a, ZTSN88]. State-Dependent [MP83]. State-of-the-Art [Ano83-40]. State-Space [MR87]. Statements [Zei83]. States [FGL82, SA89]. Static [BD81, FLS87, HP80a, Jom81, Sar84, TO80, Tho85, TR89, Aoe89b, MSS89, RS88b, YT88]. statistic [Mil86]. Statistical [Bas85b, CO82, Gho85, Gho86, Joe89, KBD85, Oku85, OC82, OO85, Ros85b, SW85, TR86]. statistics [Joe89]. Status [RB82a]. Steady [MP83]. Steady-State [MP83]. Step [Rub83, SLH81a]. Steps [TS85a]. still [VMBK89]. Stochastic [BBG86, DC89, HS89, Mol85, WY85, AI89, MBB^{+}89]. Stopping [Ros85a]. Storage [CKS85, Leu83, Sip81, YL80, Har88a, Leu86, RE89]. Stores [HFG89]. Storing [Ber85a]. Strategies [ACM87, BS87, CS89b, BEDBT89, Nta88]. Strategy [LK83, PM87, WC80, DBC88, YGBC89]. Streets [BD85]. strength [Ott89]. String [AYS84, Aoe89b]. Structural [MMA87, NS83, TTNM89, Nta88]. Structure [CRW83, HK81, HB85, KM83, MS87, Pag87, PCW85, Rom87, Sta82, Aoe89a, BL89a, BEDBT89, Kim86, Leh89, Oss89, SZ88, YD86]. Structured [BN80a, CS85a, CC80, LF84, LM84, Pag83, Reu82, VW84, vBV87, Arm89, MK89, ML86, YM89]. Structures [BS80, BDRZ85, Har80, HK84, Lee87b, LS83, TMB80a, TMB80b, Lic86, MSNT88, MIBY88]. Structuring [Shr81]. Stub [Gib87]. Students [Ano82-40]. Studies [SE84]. Study [BH83a, BH81, DHJS83, Har80, IR85, JV85, LNT85, LC87, MG83, Mil87, PB87, SjYTP85, WS83, Bev89, CCA86, DL88a, DI89, LL89c, MCH88, Oss89, SL89, TK89, TR86, Zho88]. style [HW89, Kar90, KL88]. Subdomains [WO80]. Subject [EL85]. Sublanguages [Bec80]. submodule [SA88]. Subroutines [RR87]. subset [Wan96, YLN88]. Substring [KO83]. Subsystems [Lom80]. Subtasks [KW85c]. Successful [Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Won84]. Sufficient [WW89, FN89]. Summary [Ano83g, Goe80]. Superviews [Mot87]. Support [Bob85, KL84, KBD85, Lis82, MC84, Obe88, RGP88, Rei85, RY85, Sco87, SCD83, SBD^{+}85, DJ88, DL88b, EFRV86, HW86a, HW86b, HK88a, TMTI86, VDD^{+}88]. Supporting [BHN80, PS88, WCW89]. Survey [BP83, Jon84, LM87, Gol86, YT86]. survivability [MD86]. SYGRAF [KL88]. Symbol [CL83]. Symbolic [AGPT85, CLD82, DE82, Doy85, YT88]. Symmetrical [Per81]. Synchronization [Nat85, Bag89, EFK89, GZ89, Sin86]. Synchronized [DB82, SSP85, FN89]. Synchronizing [RK83]. synchronous [AM88]. Syntactic [BH83a]. Syntax [CHP81, KP82, Pag83, Rub83]. Syntax-Directed [KP82, Rub83]. Synthesis [BI85, Bas80a, JP89, KB81, Mur80, OJ84, RM88, Smi88, ZTSN88]. Synthesizer [RDU85]. Synthesizing [Ram85, Wan96, YLN88]. System [AH87, ASHP87, ABG84, BGL83, Ber85a, Bok81, BC80, BH81, CS85a, Cot84, CTDS85, DH84, EFRV86, FGL82, GMGK84, Gud80, HKMY87, HOPW87, HO83, HB85, IR85, Jom81, Ker84, KKT80a, LA83, LB84, Lom85, Lud83, Mac83, MINH87, Mat87, MW82, MS84, MA87, Mus80, Nes87, NH85, Nis83, OC82, Pas85, Rei81, RK84, RB82b, SL87, Sar84, SC84, SKhL89, SS83, Smo81, SG84, Sta82, SW87, Stu84, Sum87, TL83, TS85b, UUD85, WY84, YHYM87, YYS81, YC83, YCT^{+}85, YN85, AH89, AN88b, Bar89, BBG^{+}88, Bev89, BR89a, CYDA88, CMNK88, DH88, DYS89, DC89, FB89, FSAC89, Gra88, Hać86, HH86, JKT88, Kau90, Kor88, KS89, LSK88, MKNS89, Man86, MW88, NH86, NS88, RSBV86, RFS88, Sam89, Sch89, SRVO88, SS88, TTTI88, TL89, Tow86]. system [Tri89, TSSP86, YW88a, ABLN85, RY85, Tra86]. Systematic [BBG^{+}85, Dro88]. Systems [ACM87, AK83, Ano82i, Beh83, BCL^{+}87, Bha82, BA83, Blu87, BCFP87, Bry81, CRW83, CY83, CR87, CA82, CR83, CS87b, Das85, DK84, Doy85, DB82, FO87, Fre87a, Fry85, GS78, GS80, Gli84, HS87, Hen80, HR82, JV83, Jai87, JLN83, KL87, KD83, KU87, KT87, KM85, Lap84, LLS81, LC84, LM84, LS85, MS85, Mag87, MPS85, MO87, MR87, Mur80, MIMH85, NS87, NSM85, NKT87, PPR^{+}83, PCW85, PB87, PS87, RH85, RH80, Ram87, Rei85, RS82, Rus80, Sam82, SC87, SSL85, Shr81, SK82, Sub85, Ter87, TR89, VPH83, Vos80, WL85, WB85b, YMW^{+}85, Zav82, ZRS87, ZY87, AK86, ADWR86, BL86, BM86, CR86, CK88a, CK88b, CCRDL88, CL89, Cho89, Dav89, DJ88, DL88b, ELZ86, ELHR88, EFRV86, GS88, Gra86, HHR89, Har88a]. systems [Har88b, HW89, HB89, JM86, Kar90, KH88, KS86, Kus89, Lau88, MKS89, MD86, MI89, NTT88, Neu86, NW89, PI86, Pol86, RE89, RB89, SSB88, SJ86, Sie88a, Sie88b, Sin86, Smi88, TH86, Von88, WPS86, WPSK86, WE86, YH88, YBL88]. Systolic [GL85].

T [Hir90]. T.-Y [Hir90]. Table [Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano84-30, Ano84-31, Ano84-32, Ano84-33, Ano84-34, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, CL83]. Tables [BR87a, Gho86]. TABS [SBD^{+}85]. Take [LM82]. Takes [Bob85]. TAME [BR88]. Tandem [PA86]. Task [KKT80b, TS85b, CTT89, MCH88]. Tasking [Ger84, BTM89, Sha88]. Tasks [LH83a, WS83, ZRS87, DM89]. taxonomy [CK88b]. TDMS [Har80]. Teaching [Wei87b]. Technical [ABLN85, Ano83-37, Mar85]. Technique [ATR82, Sch85, PNP88, US86]. Techniques [Che83, Gie82, HKMY87, Hen80, Kar85, KMB^{+}85, CCRDL88, Gol86, MI89, YT86, Zei89]. Technologies [CGP87]. Technology [ATOK85, RW87a]. TEMPO [GZ89]. temporal [WN89]. Term [Smi81a, HD89]. Termination [FR82, GL85, SSP85]. Terms [WEHR84]. Test [Chu87, Fos80, How82, RW85, Red83, SvBC87, Tai80, BMD89, Wey86, ZG89]. Testable [SC87]. testbed [JKT88]. Testing [BS87, BS82, CS85a, Cha80, CS87a, Chu87, CHR82, Dow85, DN84, Gho85, GCJ^{+}87, Gou83, How80, How82, Jal89c, Kar85, Kem85, LK83, MG83, Nta84, PM87, RC85, RR87, SvBC87, Tai80, VGV80, WO80, WC80, WHH80, Zei83, BDZ89, Dow86, FW88, Hay86, How86, MMS88, Mun88, Nta88, RY85, SL89]. Text [AC82, Fea82]. textual [Leh89]. Their [Hen80, NSM85, OJ84, Sta82, LV89]. Them [Das85]. Theorem [BI82, SD89b]. Theoretic [CS87a, Lee87a, Lee87b]. Theoretical [EL85, TMB80b]. Theories [Lit80, WO80]. Theory [GHM87, Mil85, RR87, SCD83, FZ89, LNE89, Tra86, BR89b]. Theory-W [BR89b]. Third [Ano87-69, McG84]. Thorough [Wan96]. Threaded [USV88]. throughout [WCW89]. throughput [Aky88b, OP89]. Tidier [RT81]. Time [Abb84, ATR82, All80, AI89, AK83, BD89, BD82, Bok81, CR83, Das85, EM82, FL80, Haa81, Hua85, JL87, Jef87, JV85, KS86, KK83, KKT80b, Le 87, Lei80, Mol85, Pla84, SJ87, Sta82, TAO89, VPH83, WR80, War83, YMW^{+}85, ZRS87, ZR87, AK86, BL86, DM89, FGM89, HD89, Har88b, JM86, Kim89, LY86, LBY88, MI89, San89b, San89a, Sha89, SHMM89, SG89, TSSP86]. Time-by-example [TAO89]. Time-Driven [CR83]. Time-Sensitive [Jef87]. Time-Sliced [KKT80b]. time pad [Sin86]. Timed [Gar85, HV87]. Times [Lei80, LY86]. Timestamp [HM89, LB87]. Timestamp-based [HM89]. Timestamps [BCFP87, Car87]. Timing [CR83, Das85, AM88, GF88, JM86, War86]. Tinkertoy [Ede88]. Titles [Ano82-30, Ano82-29]. Token [MW82, SJ87]. Tolerance [AK83, ABHM85, TMB80a, JC86]. Tolerant [Avi85, BJRE85, Cri85, GS78, KU87, LLS81, NKT87, Sch85, SGM87, BYC88, Kim89, LC86, PI86]. Tool [ABM87, BW80, BDRZ85, VGV80, WO85, BWWH88, FM89, DV88, JMG88, MSNT88, WCW89]. Toolpack [Ost83]. Tools [BP83, CA85, VPH83, MMS88]. Tour [Tei85]. Towsley [Kau90]. Trace [BDZ89, GKS83, HS88, Zho88]. trace-driven [Zho88]. tradeoffs [NW89, Woo86]. Training [LPBD87]. Traits [CA84]. Transaction [MPS85, Reu80, SC87, IKK88, KS89, LB88, YBL88]. Transaction-Oriented [Reu80]. Transactions [Ano80a, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano81b, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano82h, Ano82b, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano83b, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano84a, Ano85a, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano87a, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, BDRZ85, CG85b, GM82, RS82, Sin83, SBD^{+}85, Sin86]. transfer [RSBV86, Sie88a, Sie88b]. Transformation [BHN80, BS80, BM84, CC85a, War86]. Transformational [Bal81, Dea81, Fic85]. Transformations [Ano81s, Bas84, Che84, Wil81, BMPP89]. transformations-computer-aided [BMPP89]. Transition [STE^{+}82, STE^{+}83, Was85, vBV87, MZ88, ZTSN88]. Transition-Nets [Vos80]. Transition-Oriented [vBV87]. translatable [BM86]. Translating [CG85a]. Translation [BHN80, VMBK89, MKNS89, Wat88]. Translations [Rus89]. transparency [SP86]. transparent [Kim88]. Transport [LM82]. Transpose [Gol81]. Traversal [Che83]. Tree [Bok81, CF81, Har80, CTT89, Kun86, MK89, San89c]. tree-structured [MK89]. Trees [Dal87, DTV87, KW82, Man84, RT81, MI86, SP88]. Trends [Ano82t]. True [MW82]. True-Copy [MW82]. Trusted [MM87]. TUNEX [Sam89]. tuning [Sam89, SRVO88]. Tutorial [Ano80d, Ano80c, Ano81f, Ano81e, Ano82d, Ano82e, Ano82f, Ano82g, Ano83c, Ano83d, Ano83e, Ano85-72, Ano85-73]. Two [BD89, BL89b, BR87b, CP87, DTV87, HKMY87, LA83, MFKS85, PB87, RH86, Sil83a, HHR89, Har89, LB88, Lip86, MCH88, Nic89, YW88a]. Two-dimensional [RH86]. two-level [YW88a]. Two-Person [BL89b]. two-phase [Har89]. Two-Server [LA83]. Two-Stage [BD89]. Type [FM85, Gut80, Mus80, Phi84, WY84, Wil81, SF88, SS89]. Types [BHJ^{+}87, BW80, Goo81, Moi82, Whi83, BU86]. Typestate [SY86]. Typical [LM87].

U [Ale88]. UIMS [HK88b]. Uncertainty [Bas85d]. Undergraduate [LM87]. Understanding [BM82, Ben87a, BP88, JS85, Kan85, DH88]. Undo [Reu80]. Unidirectional [LM82]. unified [Mea88]. Unifying [SG84]. union [Row88]. Units [Geh82, SF88]. Universal [BN80b, JS87, MRS85]. UNIX [GZ89, DI89, MS84, PR85, RB82b, Sam89, Ker84]. Unlimited [Bab81, HH82, Luc85]. Unreliable [LC84]. unsuccessful [MK89]. update [KH88]. upon [TS86]. usage [DI89]. Use [CLP84, KR87, KW85a, KU87, Sno84, HW86a]. Used [SL84]. User [Bas85a, BGL83, CHKH88, Che80, Chi85, DN85, NS88, WPSK86, Ale88, TI88, YTT88]. User-Oriented [Che80]. User-process [CHKH88]. Users [EM82]. Users-Probabilistic [EM82]. Using [ATR82, CR83, DE82, DTV87, Gar85, GL85, HKMY87, JS87, KMB^{+}85, KP82, LH83a, LS87, LM84, Mac82, Mag87, MY80b, MM81, Omi89, Ozs85, RH80, RW85, ST87, Sch82, SS85, STE^{+}82, STE^{+}83, Vos80, Won83, YWH82, YHYM87, YG81, YC83, Aoe89a, CIK88, CC89b, Gaf86, JC86, MSS89, RMM88, SG89, TH86, VLC88, Woo86, WF89, ZTSN88]. Utilizing [UUD85].

Valid [BW84]. Validating [Das85, FM89, FZ89, Tra86]. Validation [AG83, BSP83, Hol87, RMB^{+}81, KWN88, NS88, TMB86]. value [Aky88a, KCL88]. Variable [FyY83]. Variable-Interval [FyY83]. Variables [DB85, LH83a, Len89]. Variations [Luc85]. Vector [PCMP85]. Veda [JMG88]. Vehicle [KMB^{+}85]. Verification [CC85a, CY83, CV81, DE82, HOPW87, HMS^{+}84, Hua80, JL87, LS84, LB84, RC85, Sha82, STE^{+}82, STE^{+}83, YH88, Bev89, BWWH88, BS86, Che86, Hir90, LL88, SD89c, SD89a]. Verify [Gou84, CS89a]. verifying [CL88, SD89b]. Version [Avi85, Bas85c, Cot84, DL88b, Gut80, HY87, Wei87a, BKL89, PS89]. Versions [Car87, KL84, CGR89]. Versus [BGS84, vBV87, Ten88]. Vertical [Mil87, Win88]. Vertical-Migration [Mil87]. vertically [HB89]. Very [CLP84, GKS83, SPSP87]. via [LS84, Wat88]. View [DH84, HM84, YWH82]. Views [CCW85, DAH^{+}87, Rei85]. Virtual [ATR82, AEH84, BD81, BDMS81, LF83, Mot87, ZR87, Har88a, Har88b]. Visibility [Mck85, WCW88]. Visible [BS84]. visual [CTYY89, TI88]. visualization [Moh88]. Vol [Ano80a, Ano81b, Ano82b, Ano83b, Ano85a, Ano87a]. voting [Dav89]. VSWS [FyY83].

W [BR89b]. Waiting [ATR82, WR80]. Waiting-Time [WR80]. Walkthroughs [WF84]. Wang [Ard87]. Washington [Ano85-72, Ano85-73]. watchdog [US86]. Waveform [Pap81]. Way [Ano85-67]. Weak [AYS83, How82]. Weapon [Stu84]. Web [MSNT88]. Week [Ano81f, Ano81e, Ano82d, Ano82e, Ano82f, Ano82g, Ano83c, Ano83d, Ano83e, Ano85-72, Ano85-73]. Week80 [Ano80d, Ano80c]. Weighted [Fer83]. weights [LTC89]. Well [Pro82, GS88]. Well-Delimited [Pro82]. well-founded [GS88]. West81 [Ano81f]. West82 [Ano82f, Ano82g]. West83 [Ano83c, Ano83d]. Weyuker [ZG89]. Whether [Sim86]. which [JM89]. Window [Mar85]. Windows [Gai85]. Within [Whi83, Kau90, Tow86]. without [FR82, ML86, SVR86]. Work [Ben87a]. Working [ASP82, Den80, FyY83]. Workload [IR85]. Workshop [Ano82-39, Ano83g]. Workstation [CA84, TL89]. workstation-based [TL89]. write [Len89]. Written [GM82, Shn81].

X.21 [RDU85]. XA [MA87]. Xenix [GCC^{+}87, GCJ^{+}87]. Xerox [CA84].

Year [Bal85]. Yielding [LM84].

References

Ahamad:1989:PCQ

[AA89] M. Ahamad and M. H. Ammar.

Performance characterization of quorum-consensus algorithms for replicated data. IEEE Transactions on Software Engineering, 15(4):492–496, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=16608.

Ahamad:1985:ANB

[AB85] M. Ahamad and A. J. Bernstein. An application of name based addressing to low level distributed algorithms. IEEE Transactions on Software Engineering, SE-11(1):59–67, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701898.

Abbott:1984:ISR

[Abb84]  C. Abbott.  Intervention

schedules for real-time programming. IEEE Transactions on Software Engineering, SE-10(3):268–274, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010235.

Attar:1984:SIR

[ABG84] R. Attar, P. A. Bernstein, and N. Goodman. Site initialization, recovery, and backup in a distributed database system. IEEE Transactions on Software Engineering, SE-10(6): 645–650, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010293.

Anderson:1985:SFT

[ABHM85] T. Anderson, P. A. Barrett, D. N. Halliwell, and M. R. Moulding. Software fault tolerance: An evaluation. IEEE Transactions on Software Engineering, SE-11(12):1502–1510, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701973.

Ancilotti:1983:LFA

[ABL83] P. Ancilotti, M. Boari, and N. Lijtmaer. Language features for access control. IEEE Transactions on Software Engineering, SE-9(1):16– 25, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703008.

Almes:1985:EST

[ABLN85] G. T. Almes, A. P. Black, E. D. Lazowska, and J. D. Noe. The Eden System: A technical review. IEEE Transactions on Software Engineering, SE-11(1):43–59, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701897.

Aggarwal:1987:STS

[ABM87] S. Aggarwal, D. Barbara, and K. Z. Meth. Spanner: A tool for the specification, analysis, and evaluation of protocols. IEEE Transactions on Software Engineering, SE-13(12):1218–1237, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702175.

Aggarwal:1988:SES

[ABM88] S. Aggarwal, D. Barbara, and K. Z. Meth. A software environment for the specification and analysis of problems of coordination and concurrency. IEEE Transactions on Software Engineering, 14(3):280–290, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4649.

Archer:1982:DCP

[AC82] J. Archer and R. Conway. Display condensation of program text. IEEE Transactions on Software Engineering, SE-8(5): 526–529, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702980.

Agrawal:1987:PAS

[ACM87] R. Agrawal, M. J. Carey, and L. W. Mcvoy. The performance of alternative strategies for dealing with deadlocks in database management systems. IEEE Transactions on Software Engineering, SE-13(12):1348–1363, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702184.

Adler:1988:ADF

[Adl88] M. Adler. An algebra for data flow diagram process decomposition. IEEE Transactions on Software Engineering, 14(2):169–183, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4636.

Avrunin:1986:CEA

[ADWR86] G. S. Avrunin, L. K. Dillon, J. C. Wileden, and W. E. Riddle. Constrained expressions: Adding analysis capabilities to design methods for concurrent software systems. IEEE Transactions on Software Engineering, SE-12(2):278–292, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312944.

Agrawal:1987:LIR

[AE87] R. Agrawal and A. K. Ezzat.

Location independent remote execution in NEST. IEEE Transactions on Software Engineering, SE-13(8):905–912, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702309.

Alanko:1984:VMB

[AEH84] T. O. Alanko, H. H. A. Erkio, and I. J. Haikala. Virtual memory behavior of some sorting algorithms. IEEE Transactions on Software Engineering, SE-10(4):422–431, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010255.

Al-Fedaghi:1981:MCD

[AFS81] S. Al-Fedaghi and P. Scheuermann. Mapping considerations in the design of schemas for the relational model. IEEE Transactions on Software Engineering, SE-7(1):99–111, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702806.

Albrecht:1983:SFS

[AG83] A. J. Albrecht and J. E. Gaffney. Software function, source lines of code, and development effort prediction: A software science validation. IEEE Transactions on Software Engineering, SE-9(6): 639–648, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703110.

Abdel-Ghaly:1986:ECS

[AGCL86] A. A. Abdel-Ghaly, P. Y. Chan, and B. Littlewood. Evaluation of competing software reliability predictions. IEEE Transactions on Software Engineering, SE-12(9):950–967, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313050.

Ambriola:1985:SSP

[AGPT85] V. Ambriola, F. Giannotti, D. Pedreschi, and F. Turini. Symbolic semantics and program reduction. IEEE Transactions on Software Engineering, SE-11(8):784–794, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702087.

Agrawal:1988:AER

[Agr88] R. Agrawal. Alpha: an extension of relational algebra to express a class of recursive queries. IEEE Transactions on Software Engineering, 14(7):879–885, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=42731.

Almes:1987:EOO

[AH87] G. T. Almes and C. L. Holman. Edmas: An objectoriented, locally distributed mail system. IEEE Transactions on Software Engineering, SE-13(9):1001–1009, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702322.

Abdel-Hamid:1989:DSP

[AH89]  T. K. Abdel-Hamid.  The

dynamics of software project staffing: a system dynamics based simulation approach. IEEE Transactions on Software Engineering, 15(2):109– 119, February 1989. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
21738.

Apers:1983:OAD

[AHY83] P. M. G. Apers, A. R. Hevner, and S. B. Yao. Optimization algorithms for distributed queries. IEEE Transactions on Software Engineering, SE-9(1):57–68, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703012.

See comments [LW84, CKM88].

Ammar:1989:TSD

[AI89] H. H. Ammar and S. M. R. Islam. Time scale decomposition of a class of generalized stochastic Petri net models. IEEE Transactions on Software Engineering, 15(6):809–820, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24734.

Anderson:1983:FSF

[AK83] T. Anderson and J. C. Knight. A framework for software fault tolerance in real-time systems. IEEE Transactions on Software Engineering, SE-9 (3):355–364, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703064.

Auernheimer:1986:RAS

[AK86] B. Auernheimer and R. A. Kemmerer. RT-ASLAN: A specification language for realtime systems. IEEE Transactions on Software Engineering, SE-12(9):879–889, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313044.

Akyildiz:1988:MVA

[Aky88a] I. F. Akyildiz. Mean value analysis for blocking queueing networks. IEEE Transactions on Software Engineering, 14(4):418–428, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4663.

Akyildiz:1988:EAT

[Aky88b] I. F. Akyildiz. On the exact and approximate throughput analysis of closed queuing networks with blocking. IEEE Transactions on Software Engineering, 14(1):62–70, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4623.

Alexander:1988:C

[Ale88] H. Alexander. Comments on “Formal specification of user interfaces: a comparison and evaluation of four axiomatic approaches” by U. H. Chi. IEEE Transactions on Software Engineering, 14(4):438–439, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4665. See

[Chi85].

Allchin:1980:MQT

[All80] J. E. Allchin. Modula and a question of time. IEEE Transactions on Software Engineering, SE-6(4):390–391, July/ August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702747.

Arisawa:1986:PEI

[AM86] H. Arisawa and T. Miura. On the properties of extended inclusion dependences. IEEE Transactions on Software Engineering, SE-12(11): 1098–1101, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313001.

Allan:1988:CGS

[AM88] V. H. Allan and R. A. Mueller. Compaction with general synchronous timing. IEEE Transactions on Software Engineering, 14(5):595–599, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6137.

Aiken:1988:DEH

[AN88a] A. Aiken and A. Nicolau. A development environment for horizontal microcode. IEEE Transactions on Software Engineering, 14(5):584–594, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6136.

Ambriola:1988:RAI

[AN88b] V. Ambriola and D. Notkin.

Reasoning about interactive system. IEEE Transactions on Software Engineering, 14 (2):272–276, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4645.

Anderson:1988:API

[And88]  D. P. Anderson.  Auto-

mated protocol implementation with RTAG. IEEE Transactions on Software Engineering, 14(3):291–300, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4650.

Anonymous:1980:IIT

[Ano80a]  Anonymous.  1980 index

IEEE Transactions on Software Engineering vol. SE-6. IEEE Transactions on Software Engineering, SE-6 (6):604, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702788.

Anonymous:1980:AR

[Ano80b] Anonymous. Acknowledgment of referees. IEEE Transactions on Software Engineering, SE-6 (1):110–111, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702699.

Anonymous:1980:AATb

[Ano80c]  Anonymous.  Advance an-

nouncement … tutorial week80. IEEE Transactions on Software Engineering, SE-6(6):604, November/December 1980. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702793.

Anonymous:1980:AATa

[Ano80d]  Anonymous.  Advance an-

nouncing … tutorial week80. IEEE Transactions on Software Engineering, SE-6(5): 502, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702771.

Anonymous:1980:ITSa

[Ano80e]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-6(1): c1, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702686.

Anonymous:1980:ITSb

[Ano80f] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-6(2):c1, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702706.

Anonymous:1980:ITSc

[Ano80g] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-6(3):c1, May/June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702722.

Anonymous:1980:ITSd

[Ano80h] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-6(4):c1, July/ August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702736.

Anonymous:1980:ITSe

[Ano80i]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-6(5): c1, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702753.

Anonymous:1980:ITSf

[Ano80j]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-6(6): c1, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702772.

Anonymous:1980:CPa

[Ano80k]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-6(1): 112, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702705.

Anonymous:1980:CPb

[Ano80l]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-6(2):232, March/April 1980.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702721.

Anonymous:1980:CPc

[Ano80m]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-6(3):312, May/June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702735.

Anonymous:1980:CPd

[Ano80n]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-6(5):502, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702769.

Anonymous:1980:CPe

[Ano80o]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-6(5):502, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702768.

Anonymous:1980:CPf

[Ano80p]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-6 (6):604, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702789.

Anonymous:1980:CPg

[Ano80q]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-6 (6):604, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702790.

Anonymous:1980:CF

[Ano80r]  Anonymous.  CompCon 80

Fall. IEEE Transactions on Software Engineering, SE-6 (4):398, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702752.

Anonymous:1980:CCP

[Ano80s] Anonymous. Computer communication professionals . . .. IEEE Transactions on Software Engineering, SE-6(1): 112, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702703.

Anonymous:1980:ENa

[Ano80t]  Anonymous.  Editor’s no-

tice. IEEE Transactions on Software Engineering, SE-6 (1):1, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702688.

Anonymous:1980:ENb

[Ano80u]  Anonymous.  Editor’s no-

tice. IEEE Transactions on Software Engineering, SE-6 (5):401, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702755.

Anonymous:1980:GEa

[Ano80v]  Anonymous.  Guest edito-

rial. IEEE Transactions on Software Engineering, SE-6 (2):117, March/April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702708.

Anonymous:1980:GEb

[Ano80w]  Anonymous.  Guest edito-

rial. IEEE Transactions on Software Engineering, SE-6 (6):505, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702774.

Anonymous:1980:ICSb

[Ano80x]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-6(1):c2, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702687.

Anonymous:1980:ICSd

[Ano80y]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-6(2):c2, March/April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702707.

Anonymous:1980:ICSf

[Ano80z]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-6(3):c2, May/June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702723.

Anonymous:1980:ICSh

[Ano80-27] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-6(4):c2, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702737.

Anonymous:1980:ICSj

[Ano80-28] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-6(5):c2, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702754.

Anonymous:1980:ICSl

[Ano80-29] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-6(6):c2, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702773.

Anonymous:1980:ICSa

[Ano80-30] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-6(1):112–e–112– e, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702704.

Anonymous:1980:ICSc

[Ano80-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-6(2):232– a–232–a, March/April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702720.

Anonymous:1980:ICSe

[Ano80-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-6(3): 312–a–312–a, May/June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702734.

Anonymous:1980:ICSg

[Ano80-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-6(4):398– c–398–c, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702751.

Anonymous:1980:ICSi

[Ano80-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-6(5):502–c–502– c, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702770.

Anonymous:1980:ICSk

[Ano80-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-6(6):604–i–604– i, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702792.

Anonymous:1980:ICF

[Ano80-36] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-6(1):112, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702701.

Anonymous:1980:IASa

[Ano80-37] Anonymous. Information about the Society. IEEE Transactions on Software Engineering, SE-6(1):112–c–112– c, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702702.

Anonymous:1980:IASb

[Ano80-38] Anonymous. Information about the Society. IEEE Transactions on Software Engineering, SE-6 (4):398–a–398–b, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702750.

Anonymous:1980:IASc

[Ano80-39] Anonymous. Information about the Society. IEEE Transactions on Software Engineering, SE-6(6):604–h–604– h, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702791.

Anonymous:1980:IA

[Ano80-40] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-6(1):112, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702700.

Anonymous:1981:RL

[Ano81a]  Anonymous.  1980 referees

list. IEEE Transactions on Software Engineering, SE-7 (1):148–149, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702813.

Anonymous:1981:IIT

[Ano81b]  Anonymous.  1981 index

IEEE Transactions on Software Engineering vol. SE-7. IEEE Transactions on Software Engineering, SE-7 (6):611, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702895.

Anonymous:1981:ICSb

[Ano81c]  Anonymous.  5th interna-

tional conference on software engineering. IEEE Transactions on Software Engineering, SE-7(1):151, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702816.

Anonymous:1981:ICA

[Ano81d]  Anonymous.  8th Interna-

tional Colloquium on Automata, Languages and Programming (ICALP 81). IEEE Transactions on Software Engineering, SE-7(3):332, May/ June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702846.

Anonymous:1981:AATb

[Ano81e]  Anonymous.  Advance an-

nouncement … tutorial week east82. IEEE Transactions on Software Engineering, SE-7(6):611, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702897.

Anonymous:1981:AATa

[Ano81f]  Anonymous.  Advance an-

nouncement … tutorial week west81. IEEE Transactions on Software Engineering, SE-7(5):527, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702881.

Anonymous:1981:AAC

[Ano81g] Anonymous. Announcement & call for papers. IEEE Transactions on Software Engineering, SE-7(2):257, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702834.

Anonymous:1981:ITSa

[Ano81h]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-7(1): c1, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702794.

Anonymous:1981:ITSb

[Ano81i] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-7(2):c1, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702819.

Anonymous:1981:ITSc

[Ano81j] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-7(3):c1, May/June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702837.

Anonymous:1981:ITSd

[Ano81k] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-7(4):c1, July/ August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702849.

Anonymous:1981:ITSe

[Ano81l]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-7(5): c1, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702868.

Anonymous:1981:ITSf

[Ano81m]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-7(6): c1, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702882.

Anonymous:1981:CPa

[Ano81n]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-7(1): 151, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702818.

Anonymous:1981:CPb

[Ano81o]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-7(2):257, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702836.

Anonymous:1981:CPc

[Ano81p]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-7(3):332, May/June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702848.

Anonymous:1981:CPd

[Ano81q]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-7(4):448, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702867.

Anonymous:1981:CPe

[Ano81r]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-7(5):527, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702879.

Anonymous:1981:EPT

[Ano81s] Anonymous. Editorial: Program transformations. IEEE Transactions on Software Engineering, SE-7(1):1–2, January/February 1981. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702796.

Anonymous:1981:ENb

[Ano81t] Anonymous. Editor’s note.

IEEE Transactions on Software Engineering, SE-7(4):352, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702853.

Anonymous:1981:ENa

[Ano81u]  Anonymous.  Editor’s no-

tice. IEEE Transactions on Software Engineering, SE-7 (2):161, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702821.

Anonymous:1981:ENc

[Ano81v]  Anonymous.  Editor’s no-

tice. IEEE Transactions on Software Engineering, SE-7 (6):529, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702884.

Anonymous:1981:GE

[Ano81w]  Anonymous.  Guest edito-

rial. IEEE Transactions on Software Engineering, SE-7 (6):530, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702885.

Anonymous:1981:ICSc

[Ano81x]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-7(1):c2, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702795.

Anonymous:1981:ICSf

[Ano81y]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-7(2):c2, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702820.

Anonymous:1981:ICSh

[Ano81z]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-7(3):c2, May/June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702838.

Anonymous:1981:ICSj

[Ano81-27] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-7(4):c2, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702850.

Anonymous:1981:ICSl

[Ano81-28] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-7(5):c2, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702869.

Anonymous:1981:ICSn

[Ano81-29] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-7(6):c2, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702883.

Anonymous:1981:ICSe

[Ano81-30] Anonymous. IEEE Computer Society Press. IEEE Transactions on Software Engineering, SE-7(2):257–a–257–b, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702833.

Anonymous:1981:ICSa

[Ano81-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-7(1): 151–j–151–j, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702817.

Anonymous:1981:ICSd

[Ano81-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-7(2):257– d–257–d, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702835.

Anonymous:1981:ICSg

[Ano81-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-7(3): 332–a–332–a, May/June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702847.

Anonymous:1981:ICSi

[Ano81-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-7(4):448– a–448–a, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702866.

Anonymous:1981:ICSk

[Ano81-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-7(5):527–b–527– b, September/October 1981.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702880.

Anonymous:1981:ICSm

[Ano81-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-7(6):611–f–611– f, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702896.

Anonymous:1981:ICF

[Ano81-37] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-7(1):151, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702815.

Anonymous:1981:IA

[Ano81-38] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-7(1):150, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702814.

Anonymous:1982:RL

[Ano82a]  Anonymous.  1981 referees

list. IEEE Transactions on Software Engineering, SE-8 (3):293–294, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702945.

Anonymous:1982:IIT

[Ano82b]  Anonymous.  1982 index

IEEE Transactions on Software Engineering vol. SE-18. IEEE Transactions on Software Engineering, SE-8 (6):621, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702998.

Anonymous:1982:ICSj

[Ano82c]  Anonymous.  6th interna-

tional conference on software engineering. IEEE Transactions on Software Engineering, SE-8(4):446, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702970.

Anonymous:1982:AATa

[Ano82d]  Anonymous.  Advance an-

nouncement … tutorial week east 82. IEEE Transactions on Software Engineering, SE-8(1):83, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702913.

Anonymous:1982:AATb

[Ano82e]  Anonymous.  Advance an-

nouncement … tutorial week east82. IEEE Transactions on Software Engineering, SE-8(2):160, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702925.

Anonymous:1982:AATc

[Ano82f]  Anonymous.  Advance an-

nouncement … tutorial week west82. IEEE Transactions on Software Engineering, SE-8(5):532, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702983.

Anonymous:1982:AATd

[Ano82g]  Anonymous.  Advance an-

nouncement … tutorial week west82. IEEE Transactions on Software Engineering, SE-8(6):621, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702999.

Anonymous:1982:AIN

[Ano82h] Anonymous. Announcing an important new journal IEEE Transactions on Medical Imaging. IEEE Transactions on Software Engineering, SE-8 (2):160–a–160–a, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702924.

Anonymous:1982:AIC

[Ano82i]  Anonymous.  Announcing

the 3rd international conference on distributed computing systems. IEEE Transactions on Software Engineering, SE-8(3):294, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702946.

Anonymous:1982:ITSa

[Ano82j]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-8(1): c1, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702898.

Anonymous:1982:ITSb

[Ano82k] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-8(2):c1, March/ April 1982. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702914.

Anonymous:1982:ITSc

[Ano82l] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-8(3):c1, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702930.

Anonymous:1982:ITSd

[Ano82m] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-8(4):c1, July/ August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702950.

Anonymous:1982:ITSe

[Ano82n]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-8(5): c1, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702973.

Anonymous:1982:ITSf

[Ano82o]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-8(6): c1, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702984.

Anonymous:1982:CPa

[Ano82p]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-8(1): 83, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702911.

Anonymous:1982:CPb

[Ano82q]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-8(2):160, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702929.

Anonymous:1982:CPd

[Ano82r]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-8(3):294, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702949.

Anonymous:1982:CPe

[Ano82s]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-8 (6):621, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703002.

Anonymous:1982:CPT

[Ano82t] Anonymous. Call for papers 1983 trends and applications conference. IEEE Transactions on Software Engineering, SE-8(6):621, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703000.

Anonymous:1982:CPc

[Ano82u] Anonymous. Call for participation. IEEE Transactions on Software Engineering, SE-8(2):160, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702926.

Anonymous:1982:E

[Ano82v] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-8(4):297, July/ August 1982. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702952.

Anonymous:1982:ICSb

[Ano82w]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-8(1):c2, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702899.

Anonymous:1982:ICSe

[Ano82x]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-8(2):c2, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702915.

Anonymous:1982:ICSh

[Ano82y]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-8(3):c2, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702931.

Anonymous:1982:ICSk

[Ano82z]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-8(4):c2, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702951.

Anonymous:1982:ICSm

[Ano82-27] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-8(5):c2, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702974.

Anonymous:1982:ICSo

[Ano82-28] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-8(6):c2, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702985.

Anonymous:1982:ICSf

[Ano82-29] Anonymous. IEEE Computer Society Press checklist of new titles. IEEE Transactions on Software Engineering, SE-8(3):294–b–294–b, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702947.

Anonymous:1982:ICSd

[Ano82-30] Anonymous. IEEE Computer Society Press checklist of new titles December 1981. IEEE Transactions on Software Engineering, SE-8 (2):160–d–160–d, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702927.

Anonymous:1982:ICSa

[Ano82-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-8(1): 83–f–83–f, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702912.

Anonymous:1982:ICSc

[Ano82-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-8(2):160– e–160–e, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702928.

Anonymous:1982:ICSg

[Ano82-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-8(3): 294–c–294–c, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702948.

Anonymous:1982:ICSi

[Ano82-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-8(4):446– c–446–c, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702971.

Anonymous:1982:ICSl

[Ano82-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-8(5):532–a–532– a, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702982.

Anonymous:1982:ICSn

[Ano82-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-8(6):621–h–621– h, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703001.

Anonymous:1982:ICF

[Ano82-37] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-8(1):83, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702909.

Anonymous:1982:IA

[Ano82-38] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-8(1):82, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702908.

Anonymous:1982:SSE

[Ano82-39] Anonymous. The second software engineering standards applications workshop call for papers. IEEE Transactions on Software Engineering, SE-8(4):446, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702972.

Anonymous:1982:SCS

[Ano82-40] Anonymous. Students in computer science and engineering! IEEE Transactions on Software Engineering, SE-8(1):83, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702910.

Anonymous:1983:RL

[Ano83a]  Anonymous.  1982 refer-

ees list. IEEE Transactions on Software Engineering, SE-9(3):372, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703067.

Anonymous:1983:IIT

[Ano83b]  Anonymous.  1983 index

IEEE Transactions on Software Engineering vol. SE-9. IEEE Transactions on Software Engineering, SE-9 (6):747, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703120.

Anonymous:1983:AATa

[Ano83c]  Anonymous.  Advance an-

nouncement … tutorial week west83. IEEE Transactions on Software Engineering, SE-9(4):538, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703092.

Anonymous:1983:AATb

[Ano83d]  Anonymous.  Advance an-

nouncement … tutorial week west83. IEEE Transactions on Software Engineering, SE-9(5):634, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703106.

Anonymous:1983:AATc

[Ano83e]  Anonymous.  Advance an-

nouncement tutorial week east84. IEEE Transactions on Software Engineering, SE-9(6):747, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703125.

Anonymous:1983:AP

[Ano83f] Anonymous. Advance program. IEEE Transactions on Software Engineering, SE-9(4):538, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703090.

Anonymous:1983:APS

[Ano83g] Anonymous. Advance program summary software maintenance workshop. IEEE Transactions on Software Engineering, SE-9(5):634, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703104.

Anonymous:1983:ITSa

[Ano83h]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-9(1): c1, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703003.

Anonymous:1983:ITSb

[Ano83i] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-9(2):c1, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703024.

Anonymous:1983:ITSc

[Ano83j] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-9(3):c1, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703044.

Anonymous:1983:ITSd

[Ano83k] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-9(4):c1, July/ August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703070.

Anonymous:1983:ITSe

[Ano83l]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-9(5): c1, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703093.

Anonymous:1983:ITSf

[Ano83m]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-9(6): c1, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703107.

Anonymous:1983:CPa

[Ano83n]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-9(1): 115, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703023.

Anonymous:1983:CPb

[Ano83o]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-9(2):213, March/April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703043.

Anonymous:1983:CPc

[Ano83p]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-9(3):372, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703069.

Anonymous:1983:CPd

[Ano83q]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-9(4):538, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703089.

Anonymous:1983:CPe

[Ano83r]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-9(5):634, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703103.

Anonymous:1983:CPS

[Ano83s] Anonymous. Call for papers small computer (r)evolution. IEEE Transactions on Software Engineering, SE-9(6):747, November/December 1983. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703122.

Anonymous:1983:DE

[Ano83t]  Anonymous.  Database en-

gineering. IEEE Transactions on Software Engineering, SE-9(2):213, March/April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703040.

Anonymous:1983:Ea

[Ano83u] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-9(2):117, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703026.

Anonymous:1983:Eb

[Ano83v] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-9(3):217, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703046.

Anonymous:1983:ICSb

[Ano83w]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-9(1):c2, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703004.

Anonymous:1983:ICSd

[Ano83x]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-9(2):c2, March/April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703025.

Anonymous:1983:ICSf

[Ano83y]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-9(3):c2, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703045.

Anonymous:1983:ICSh

[Ano83z]  Anonymous.  IEEE Com-

puter Society. IEEE Transactions on Software Engineering, SE-9(4):c2, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703071.

Anonymous:1983:ICSj

[Ano83-27] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-9(5):c2, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703094.

Anonymous:1983:ICSl

[Ano83-28] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-9(6):c2, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703108.

Anonymous:1983:ICSa

[Ano83-29] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-9(1):115–b–115– b, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703022.

Anonymous:1983:ICSc

[Ano83-30] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-9(2):213– d–213–d, March/April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703042.

Anonymous:1983:ICSe

[Ano83-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-9(3): 372–a–372–a, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703068.

Anonymous:1983:ICSg

[Ano83-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-9(4):538– c–538–c, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703091.

Anonymous:1983:ICSi

[Ano83-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-9(5):634–c–634– c, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703105.

Anonymous:1983:ICSk

[Ano83-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-9(6):747–n–747– n, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703124.

Anonymous:1983:ICF

[Ano83-35] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-9(1):115, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703021.

Anonymous:1983:IA

[Ano83-36] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-9(1):114, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703020.

Anonymous:1983:JUQ

[Ano83-37] Anonymous. Join us in the quest for technical excellence. IEEE Transactions on Software Engineering, SE-9(6):747, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703123.

Anonymous:1983:NII

[Ano83-38] Anonymous. Nominations invited for IEEE awards. IEEE Transactions on Software Engineering, SE-9(2):213– b–213–c, March/April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703041.

Anonymous:1983:PA

[Ano83-39] Anonymous. Preliminary announcement. IEEE Transactions on Software Engineering, SE-9(6):747, November/ December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1703121.

Anonymous:1983:SAI

[Ano83-40] Anonymous. State-of-the-art issues in distributed databases. IEEE Transactions on Software Engineering, SE-9(3):218, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1703047.

Anonymous:1984:IIT

[Ano84a] Anonymous. 1984 index IEEE Transactions on Software Engineering. IEEE Transactions on Software Engineering, SE-10(6):858–862, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010315.

Anonymous:1984:AA

[Ano84b]  Anonymous.  Advance an-

nouncement. IEEE Transactions on Software Engineering, SE-10(1):124, January/ February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010208.

Anonymous:1984:BP

[Ano84c] Anonymous. [breaker page].

IEEE Transactions on Software Engineering, SE-10(1): nil1, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010190.

Anonymous:1984:CPa

[Ano84d]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-10 (2):220, March/April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010225.

Anonymous:1984:CPb

[Ano84e]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-10 (4):472, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010263.

Anonymous:1984:CPc

[Ano84f] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-10(5): 610–612, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010284.

Anonymous:1984:CPd

[Ano84g] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-10(6): 863–864, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010316.

Anonymous:1984:Ea

[Ano84h] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-10(1):1, January/February 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010191.

Anonymous:1984:Eb

[Ano84i] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-10(2):125, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010213.

Anonymous:1984:EAa

[Ano84j]  Anonymous.  Editorial an-

nouncement. IEEE Transactions on Software Engineering, SE-10(5):473, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010268.

Anonymous:1984:Fa

[Ano84k] Anonymous. Foreword. IEEE Transactions on Software Engineering, SE-10(2):126–127, March/April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010214.

Anonymous:1984:Fb

[Ano84l] Anonymous. Foreword. IEEE Transactions on Software Engineering, SE-10(5):474–477, September/October 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010269.

Anonymous:1984:FCa

[Ano84m]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(1): c1, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010188.

Anonymous:1984:FCb

[Ano84n]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(1): c2, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010209.

Anonymous:1984:FCc

[Ano84o]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(2): c2, March/April 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5010226.

Anonymous:1984:FCd

[Ano84p]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(2): c1, March/April 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5010211.

Anonymous:1984:FCe

[Ano84q]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(3):c2, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010242.

Anonymous:1984:FCf

[Ano84r]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(3):c1, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010228.

Anonymous:1984:FCg

[Ano84s]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(4):c2, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010264.

Anonymous:1984:FCh

[Ano84t]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(4):c1, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010244.

Anonymous:1984:FCi

[Ano84u]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(5): c1, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010266.

Anonymous:1984:FCj

[Ano84v]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(5): c2, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010285.

Anonymous:1984:FCk

[Ano84w]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(6): c2, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010317.

Anonymous:1984:FCl

[Ano84x]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, SE-10(6): c1, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010287.

Anonymous:1984:GES

[Ano84y]  Anonymous.  Guest edito-

rial: Software engineering project management. IEEE Transactions on Software Engineering, SE-10(1):2–3, January/February 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010192.

Anonymous:1984:ICS

[Ano84z] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-10(1):nil1, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010189.

Anonymous:1984:ICF

[Ano84-27] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-10 (1):121–122, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010206.

Anonymous:1984:IA

[Ano84-28] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-10(1):123, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010207.

Anonymous:1984:NEC

[Ano84-29] Anonymous. A note from the Editor-in-Chief. IEEE Transactions on Software Engineering, SE-10(6):613, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5010289.

Anonymous:1984:TCa

[Ano84-30] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-10(2):nil1, March/April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010212.

Anonymous:1984:TCb

[Ano84-31] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-10(3):nil1, May/June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010229.

Anonymous:1984:TCc

[Ano84-32] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-10(4):nil1, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010245.

Anonymous:1984:TCd

[Ano84-33] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-10(5):nil1, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010267.

Anonymous:1984:TCe

[Ano84-34] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-10 (6):nil1, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010288.

Anonymous:1985:IIT

[Ano85a] Anonymous. 1985 index IEEE Transactions on Software Engineering vol. SE-11. IEEE Transactions on Software Engineering, SE-11(12):1517, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701975.

Anonymous:1985:AP

[Ano85b] Anonymous. Advance program.

IEEE Transactions on Software Engineering, SE-11(3):311–313, March 1985. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702007.

Anonymous:1985:BBSa

[Ano85c] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(3):315–a–315–a, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702010.

Anonymous:1985:BBSb

[Ano85d] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(4):443–c–443–c, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702030.

Anonymous:1985:BBSc

[Ano85e] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(5):496–b–496– b, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702043.

Anonymous:1985:BBSd

[Ano85f] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(7):655–c–655–c, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702073.

Anonymous:1985:BBSe

[Ano85g] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(8):823–c–823–c, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702096.

Anonymous:1985:BBSf

[Ano85h] Anonymous. Beyond the basics:

A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(9):992–b–992–b, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702122.

Anonymous:1985:IS

[Ano85i]  Anonymous.  IEEE Soft-

ware. IEEE Transactions on Software Engineering, SE-11(9):829–a–829–d, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702099.

Anonymous:1985:ITSa

[Ano85j] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(1):c1, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701890.

Anonymous:1985:ITSb

[Ano85k] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(2):c1, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701980.

Anonymous:1985:ITSc

[Ano85l] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(3):c1, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701997.

Anonymous:1985:ITSd

[Ano85m] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(4):c1, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702013.

Anonymous:1985:ITSe

[Ano85n] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(5):c1, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702031.

Anonymous:1985:ITSf

[Ano85o] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(6):c1, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702044.

Anonymous:1985:ITSg

[Ano85p] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(7):c1, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702058.

Anonymous:1985:ITSh

[Ano85q] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(8):c1, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702074.

Anonymous:1985:ITSi

[Ano85r] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(9):c1, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702097.

Anonymous:1985:ITSj

[Ano85s]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11 (10):c1–1252, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701911.

Anonymous:1985:ITSk

[Ano85t] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(11):c1, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701942.

Anonymous:1985:ITSl

[Ano85u] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-11(12):c1, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701959.

Anonymous:1985:CPa

[Ano85v]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-11 (1):142, January 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1701907.

Anonymous:1985:CPb

[Ano85w]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-11(2):240, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701994.

Anonymous:1985:CPc

[Ano85x]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-11 (3):314, March 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702008.

Anonymous:1985:CPd

[Ano85y]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-11 (4):442, April 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702027.

Anonymous:1985:CPe

[Ano85z]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-11 (5):496, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702041.

Anonymous:1985:CPf

[Ano85-27] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-11 (6):572, June 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702055.

Anonymous:1985:CPg

[Ano85-28] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-11 (7):653, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702069.

Anonymous:1985:CPh

[Ano85-29] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-11(8):823, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702093.

Anonymous:1985:CPi

[Ano85-30] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-11(9):992, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702120.

Anonymous:1985:CPj

[Ano85-31] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-11(10):1252, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701940.

Anonymous:1985:CPk

[Ano85-32] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-11(11):1408, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701958.

Anonymous:1985:CPl

[Ano85-33] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-11(12):1517, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701977.

Anonymous:1985:CPm

[Ano85-34] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-11(12):1517, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701979.

Anonymous:1985:CPPa

[Ano85-35] Anonymous. Call for papers and panel session proposals. IEEE Transactions on Software Engineering, SE-11(2):240, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701996.

Anonymous:1985:CPPb

[Ano85-36] Anonymous. Call for papers and panel session proposals. IEEE Transactions on Software Engineering, SE-11(3):315, March 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702012.

Anonymous:1985:Ea

[Ano85-37] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-11(1):1–2, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701892.

Anonymous:1985:Eb

[Ano85-38] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-11(5):445, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702033.

Anonymous:1985:Ec

[Ano85-39] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-11(10):993, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701913.

Anonymous:1985:ICSb

[Ano85-40] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(1):c2, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701891.

Anonymous:1985:ICSd

[Ano85-41] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(2):c2, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701981.

Anonymous:1985:ICSf

[Ano85-42] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(3):c2, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701998.

Anonymous:1985:ICSh

[Ano85-43] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(4):c2, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702014.

Anonymous:1985:ICSj

[Ano85-44] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(5):c2, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702032.

Anonymous:1985:ICSl

[Ano85-45] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(6):c2, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702045.

Anonymous:1985:ICSn

[Ano85-46] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(7):c2, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702059.

Anonymous:1985:ICSp

[Ano85-47] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(8):c2, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702075.

Anonymous:1985:ICSr

[Ano85-48] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(9):c2, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702098.

Anonymous:1985:ICSt

[Ano85-49] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(10):c2, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701912.

Anonymous:1985:ICSv

[Ano85-50] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(11):c2, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701943.

Anonymous:1985:ICSx

[Ano85-51] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-11(12):c2, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701960.

Anonymous:1985:ICSa

[Ano85-52] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(1): 143–b–143–b, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701909.

Anonymous:1985:ICSc

[Ano85-53] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(2): 240–a–240–a, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701995.

Anonymous:1985:ICSe

[Ano85-54] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(3):315–b–315–b, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702011.

Anonymous:1985:ICSg

[Ano85-55] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(4):443–b–443–b, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702029.

Anonymous:1985:ICSi

[Ano85-56] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(5):496–a–496–a, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702042.

Anonymous:1985:ICSk

[Ano85-57] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(6):572–a–572–a, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702056.

Anonymous:1985:ICSm

[Ano85-58] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(7):655–b–655– b, July 1985. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702072.

Anonymous:1985:ICSo

[Ano85-59] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(8):823–b–823–b, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702095.

Anonymous:1985:ICSq

[Ano85-60] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(9): 992–a–992–a, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702121.

Anonymous:1985:ICSs

[Ano85-61] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(10): 1252–a–1252–a, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701941.

Anonymous:1985:ICSu

[Ano85-62] Anonymous. IEEE Computer Society publications.

IEEE Transactions on Software Engineering, SE-11(11):1408– a–1408–a, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701957.

Anonymous:1985:ICSw

[Ano85-63] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-11(12): 1517–l–1517–l, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701978.

Anonymous:1985:ICFa

[Ano85-64] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-11(1):143, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701908.

Anonymous:1985:ICFb

[Ano85-65] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-11(4):443, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702028.

Anonymous:1985:ICFc

[Ano85-66] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-11(7):655, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702071.

Anonymous:1985:IFY

[Ano85-67] Anonymous. IEEE finding your way. IEEE Transactions on Software Engineering, SE-11(12):1517–j–1517–j, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701976.

Anonymous:1985:IAa

[Ano85-68] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-11(1):143, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701910.

Anonymous:1985:IAb

[Ano85-69] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-11(3):315, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702009.

Anonymous:1985:IAc

[Ano85-70] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-11(6):572, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702057.

Anonymous:1985:LR

[Ano85-71] Anonymous. List of referees. IEEE Transactions on Software Engineering, SE-11(12):1410, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701962.

Anonymous:1985:TWWa

[Ano85-72] Anonymous. Tutorial week Washington 1985. IEEE Transactions on Software Engineering, SE-11(7):654, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702070.

Anonymous:1985:TWWb

[Ano85-73] Anonymous. Tutorial week Washington 1985. IEEE Transactions on Software Engineering, SE-11(8):823–a–823–a, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702094.

Anonymous:1986:AAE

[Ano86a] Anonymous. Ada — applications and environments. IEEE Transactions on Software Engineering, SE-12(2):1, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312951.

Anonymous:1986:AP

[Ano86b] Anonymous. Advance program.

IEEE Transactions on Software Engineering, SE-12(12):1, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313013.

Anonymous:1986:AI

[Ano86c]  Anonymous.  Author in-

dex. IEEE Transactions on Software Engineering, SE-12(12):1–7, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313011.

Anonymous:1986:CPa

[Ano86d]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-12 (1):182, January 1986. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6312931.

Anonymous:1986:CPb

[Ano86e] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-12(2):1, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312952.

Anonymous:1986:CPc

[Ano86f]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-12(2):374, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312950.

Anonymous:1986:CPd

[Ano86g] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-12(3):1, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312895.

Anonymous:1986:CPe

[Ano86h] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-12(3):1, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312894.

Anonymous:1986:CPf

[Ano86i] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-12(3):1, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312892.

Anonymous:1986:CPg

[Ano86j]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-12 (3):502, March 1986. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6312893.

Anonymous:1986:CPh

[Ano86k] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-12(4):1, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312909.

Anonymous:1986:CPi

[Ano86l] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-12(4):586–587, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312908.

Anonymous:1986:CPj

[Ano86m]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-12 (5):676, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        6312962.

Anonymous:1986:CPk

[Ano86n] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-12(6):1, June 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312971.

Anonymous:1986:CPl

[Ano86o] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-12(6):722–723, June 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312970.

Anonymous:1986:CPm

[Ano86p] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-12(7):1, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312981.

Anonymous:1986:CPn

[Ano86q]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-12 (7):799, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6312980.

Anonymous:1986:CPo

[Ano86r] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-12(8):875–876, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312992.

Anonymous:1986:CPp

[Ano86s]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-12 (9):994–995, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313054.

Anonymous:1986:CPq

[Ano86t]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-12(11):1102–1103, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313002.

Anonymous:1986:CPr

[Ano86u]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-12(12):1153–1154, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313012.

Anonymous:1986:C

[Ano86v] Anonymous. Contents. IEEE Transactions on Software Engineering, SE-12(1):1, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312911.

Anonymous:1986:DER

[Ano86w] Anonymous. Data engineering registration form. IEEE Transactions on Software Engineering, SE-12(12):1, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313014.

Anonymous:1986:Ea

[Ano86x] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-12(5):589, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312954.

Anonymous:1986:Eb

[Ano86y] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-12(6):677, June 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312964.

Anonymous:1986:Ec

[Ano86z] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-12(12):1105, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313005.

Anonymous:1986:F

[Ano86-27] Anonymous. Foreword. IEEE Transactions on Software Engineering, SE-12(7):725, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312973.

Anonymous:1986:ICF

[Ano86-28] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-12(1):1–2, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312932.

Anonymous:1986:IFC

[Ano86-29] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, SE-12 (1):c2, January 1986. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6312912.

Anonymous:1987:IIT

[Ano87a] Anonymous. 1987 index IEEE Transactions on Software Engineering vol. SE-13. IEEE Transactions on Software Engineering, SE-13(12):1371, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702187.

Anonymous:1987:AIC

[Ano87b] Anonymous. 9th annual international conference on software engineering. IEEE Transactions on Software Engineering, SE-13(2):296, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702209.

Anonymous:1987:BBSa

[Ano87c] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(3):407–c–407–c, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702230.

Anonymous:1987:BBSb

[Ano87d] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(4):506–d–506–d, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702248.

Anonymous:1987:BBSc

[Ano87e] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(5):607–c–607–c, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702264.

Anonymous:1987:BBSd

[Ano87f] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(8):987–a–987–a, August 1987. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702316.

Anonymous:1987:BBSe

[Ano87g] Anonymous. Beyond the basics: A software publication series for successful computer professionals from the IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(10):1139–a–1139–a, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702154.

Anonymous:1987:ITSa

[Ano87h] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(1):c1, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702123.

Anonymous:1987:ITSb

[Ano87i] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(2):c1, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702191.

Anonymous:1987:ITSc

[Ano87j] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(3):c1, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702212.

Anonymous:1987:ITSd

[Ano87k] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(4):c1, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702231.

Anonymous:1987:ITSe

[Ano87l] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(5):c1, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702249.

Anonymous:1987:ITSf

[Ano87m] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(6):c1, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702265.

Anonymous:1987:ITSg

[Ano87n] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(7):c1, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702285.

Anonymous:1987:ITSh

[Ano87o] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(8):c1, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702302.

Anonymous:1987:ITSi

[Ano87p] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(9):c1, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702319.

Anonymous:1987:ITSj

[Ano87q] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(10):c1, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702143.

Anonymous:1987:ITSk

[Ano87r] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(11):c1, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702157.

Anonymous:1987:ITSl

[Ano87s] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-13(12):c1, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702172.

Anonymous:1987:CPa

[Ano87t]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-13(2):296, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702211.

Anonymous:1987:CPb

[Ano87u] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-13(3):406–407, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702227.

Anonymous:1987:CPc

[Ano87v]  Anonymous.  Call for pa-

pers. IEEE Transactions on Software Engineering, SE-13 (4):506, April 1987. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702244.

Anonymous:1987:CSa

[Ano87w] Anonymous. The computer society. IEEE Transactions on Software Engineering, SE-13(9):1048–b–1048–b, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702329.

Anonymous:1987:CSb

[Ano87x] Anonymous. The computer society. IEEE Transactions on Software Engineering, SE-13(10):1139–c–1139–c, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702156.

Anonymous:1987:CSc

[Ano87y] Anonymous. The computer society. IEEE Transactions on Software Engineering, SE-13(11):1216–b–1216–b, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702171.

Anonymous:1987:CSd

[Ano87z] Anonymous. The computer society. IEEE Transactions on Software Engineering, SE-13(12):1371–k–1371–k, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702190.

Anonymous:1987:E

[Ano87-27] Anonymous. Editorial. IEEE Transactions on Software Engineering, SE-13(12):1217, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702174.

Anonymous:1987:GEN

[Ano87-28] Anonymous. Guest Editors’ note. IEEE Transactions on Software Engineering, SE-13(2):125–128, February 1987. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702193.

Anonymous:1987:ICSb

[Ano87-29] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(1):c2, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702124.

Anonymous:1987:ICSd

[Ano87-30] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(2):c2, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702192.

Anonymous:1987:ICSf

[Ano87-31] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(3):c2, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702213.

Anonymous:1987:ICSh

[Ano87-32] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(4):c2, April 1987.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702232.

Anonymous:1987:ICSj

[Ano87-33] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(5):c2, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702250.

Anonymous:1987:ICSl

[Ano87-34] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(6):c2, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702266.

Anonymous:1987:ICSn

[Ano87-35] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(7):c2, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702286.

Anonymous:1987:ICSp

[Ano87-36] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(8):c2, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702303.

Anonymous:1987:ICSr

[Ano87-37] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(9):c2, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702320.

Anonymous:1987:ICSt

[Ano87-38] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(10):c2, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702144.

Anonymous:1987:ICSv

[Ano87-39] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(11):c2, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702158.

Anonymous:1987:ICSx

[Ano87-40] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-13(12):c2, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702173.

Anonymous:1987:ICSa

[Ano87-41] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(1): 119–f–119–f, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702141.

Anonymous:1987:ICSc

[Ano87-42] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(2): 296–e–296–e, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702210.

Anonymous:1987:ICSe

[Ano87-43] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(3):407–b–407–b, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702229.

Anonymous:1987:ICSg

[Ano87-44] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(4):506–c–506–c, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702247.

Anonymous:1987:ICSi

[Ano87-45] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(5):607–b–607– b, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702263.

Anonymous:1987:ICSk

[Ano87-46] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(6):760–a–760–a, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702283.

Anonymous:1987:ICSm

[Ano87-47] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(7):867–b–867– b, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702300.

Anonymous:1987:ICSo

[Ano87-48] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(8):987–b–987–b, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702317.

Anonymous:1987:ICSq

[Ano87-49] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(9):1048– a–1048–a, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702328.

Anonymous:1987:ICSs

[Ano87-50] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(10): 1139–b–1139–b, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702155.

Anonymous:1987:ICSu

[Ano87-51] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(11):1216– a–1216–a, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702170.

Anonymous:1987:ICSw

[Ano87-52] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-13(12): 1371–j–1371–j, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702189.

Anonymous:1987:ICF

[Ano87-53] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-13(1):119, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702138.

Anonymous:1987:Ib

[Ano87-54] Anonymous. In this issue. IEEE Transactions on Software Engineering, SE-13(4):506, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702245.

Anonymous:1987:Ic

[Ano87-55] Anonymous. In this issue. IEEE Transactions on Software Engineering, SE-13(5):607, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702261.

Anonymous:1987:Id

[Ano87-56] Anonymous. In this issue. IEEE Transactions on Software Engineering, SE-13(7):867, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702301.

Anonymous:1987:If

[Ano87-57] Anonymous. In this issue. IEEE Transactions on Software Engineering, SE-13(8):987, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702318.

Anonymous:1987:Ig

[Ano87-58] Anonymous. In this issue. IEEE Transactions on Software Engineering, SE-13(12):1371, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702188.

Anonymous:1987:IEa

[Ano87-59] Anonymous. The information edge. IEEE Transactions on Software Engineering, SE-13(3):407–a–407–a, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702228.

Anonymous:1987:IEb

[Ano87-60] Anonymous. The information edge. IEEE Transactions on Software Engineering, SE-13(5):607–a–607–a, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702262.

Anonymous:1987:IAa

[Ano87-61] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-13(1):119, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702142.

Anonymous:1987:IAb

[Ano87-62] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-13(6):760, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702284.

Anonymous:1987:LRa

[Ano87-63] Anonymous. List of referees. IEEE Transactions on Software Engineering, SE-13(1):2, January 1987. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702126.

Anonymous:1987:LRb

[Ano87-64] Anonymous. List of referees. IEEE Transactions on Software Engineering, SE-13 (3):302, March 1987. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702215.

Anonymous:1987:LRc

[Ano87-65] Anonymous. List of referees.

IEEE Transactions on Software Engineering, SE-13(8):871, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702305.

Anonymous:1987:MA

[Ano87-66] Anonymous. Membership application. IEEE Transactions on Software Engineering, SE-13(4):506–b–506–b, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702246.

Anonymous:1987:OLIa

[Ano87-67] Anonymous. The one leveraged investment with a guaranteed return. IEEE Transactions on Software Engineering, SE-13(1):119–d–119–e, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702140.

Anonymous:1987:OLIb

[Ano87-68] Anonymous. The one leveraged investment with a guaranteed return. IEEE Transactions on Software Engineering, SE-13(7):867–867–a, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702299.

Anonymous:1987:TIC

[Ano87-69] Anonymous. The third international conference on data engineering. IEEE Transactions on Software Engineering, SE-13(1):119, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702139.

Anonymous:1988:F

[Ano88] Anonymous. Foreword. IEEE Transactions on Software Engineering, 14(4):521, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4674.

Aoe:1989:EDS

[Aoe89a]  J. I. Aoe.  An efficient

digital search algorithm by using a double-array structure. IEEE Transactions on Software Engineering, 15(9): 1066–1077, September 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31365.

Aoe:1989:EIS

[Aoe89b] J. I. Aoe. An efficient implementation of static string pattern matching machines. IEEE Transactions on Software Engineering, 15(8):1010–1016, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31357.

Arthur:1981:DAS

[AR81] J. Arthur and J. Ramanathan.

Design of analyzers for selective program analysis. IEEE Transactions on Software Engineering, SE-7(1):39–51, January/February 1981. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702801.

Ardis:1987:EWI

[Ard87] M. A. Ardis. The evolution of Wang Institute’s Master of Software Engineering Program. IEEE Transactions on Software Engineering, SE-13(11):1149–1155, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702161.

Armenise:1989:SAP

[Arm89] P. Armenise. A structured approach to program optimization. IEEE Transactions on Software Engineering, 15(2):101–108, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=21737.

Adelson:1985:RDE

[AS85] B. Adelson and E. Soloway. The role of domain experience in software design. IEEE Transactions on Software Engineering, SE-11(11):1351–1360, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701951.

Akyildiz:1988:AAL

[AS88] I. F. Akyildiz and A. Sieber. Approximate analysis of load dependent general queueing networks. IEEE Transactions on Software Engineering, 14(11):1537–1545, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=9042.

Andrews:1987:DSD

[ASHP87] G. R. Andrews, R. D. Schlichting, R. Hayes, and T. D. M. Purdin. The design of the saguaro distributed operating system. IEEE Transactions on Software Engineering, SE-13(1):104–118, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702137.

Abu-Sufah:1982:SRW

[ASP82] W. A. Abu-Sufah and D. A.

Padua. Some results on the working set anomalies in numerical programs. IEEE Transactions on Software Engineering, SE-8(2):97–106, March/ April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702917.

Azuma:1985:SHD

[ATOK85] M. Azuma, T. Tabata, Y. Oki, and S. Kamiya. SPD: A humanized documentation technology. IEEE Transactions on Software Engineering, SE-11 (9):945–953, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702113.

Agrawala:1982:ARU

[ATR82] A. K. Agrawala, S. K. Tripathi, and G. Ricart. Adaptive routing using a virtual waiting time technique. IEEE Transactions on Software Engineering, SE-8(1):76–81, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702907.

Akyildiz:1989:CSM

[AV89] I. F. Akyildiz and H. Von Brand. Central server models with multiple job classes, state dependent routing, and rejection blocking. IEEE Transactions on Software Engineering, 15(10):1305–1312, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559784.

Avizienis:1985:VAF

[Avi85] A. Avizienis. The N-version approach to fault-tolerant software. IEEE Transactions on Software Engineering, SE-11(12):1491–1501, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701972.

Aoe:1983:PMR

[AYS83] J. Aoe, Y. Yamamoto, and R. Shimada. A practical method for reducing weak precedence parsers. IEEE Transactions on Software Engineering, SE-9(1):25–30, January/February 1983. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703009.

Aoe:1984:MIS

[AYS84] J. Aoe, Y. Yamamoto, and R. Shimada. A method for improving string pattern matching machines. IEEE Transactions on Software Engineering, SE-10 (1):116–120, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010205.

Bhuyan:1983:PAF

[BA83] L. N. Bhuyan and D. P. Agarval. Performance analysis of FFT algorithms on multiprocessor systems. IEEE Transactions on Software Engineering, SE-9(4):512–521, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703084.

Baber:1981:MRD

[Bab81] R. L. Baber. A method for representing data items of unlimited length in a computer memory. IEEE Transactions on Software Engineering, SE-7(6): 590–593, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702891.

See comments [HH82].

Baber:1982:ARS

[Bab82] R. L. Baber. Author’s reply. IEEE Transactions on Software Engineering, SE-8 (6):621, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702997.

Bagrodia:1989:PSD

[Bag89] R. Bagrodia. Process synchronization: design and performance evaluation of distributed algorithms. IEEE Transactions on Software Engineering, 15(9):1053–1065, September 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=31364.

Baker:1988:EFS

[Bak88]  C. T. Baker.  Effects of

field service on software reliability. IEEE Transactions on Software Engineering, 14(2):254–258, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4642.

Balzer:1981:TIE

[Bal81] R. Balzer. Transformational implementation: An example.

IEEE Transactions on Software Engineering, SE-7(1):3– 14, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702797.

Balzer:1985:YPA

[Bal85] R. Balzer. A 15 year perspective on automatic programming. IEEE Transactions on Software Engineering, SE-11(11):1257–1268, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701945.

Barstow:1985:DSA

[Bar85]  D. R. Barstow.  Domain-

specific automatic programming. IEEE Transactions on Software Engineering, SE-11(11):1321–1336, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701949.

Barrett:1989:FMA

[Bar89] G. Barrett. Formal methods applied to a floatingpoint number system. IEEE Transactions on Software Engineering, 15(5):611–621, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24710.

Basu:1980:NSI

[Bas80a]  S. K. Basu.  A note on

synthesis of inductive assertions. IEEE Transactions on Software Engineering, SE-6(1):32–39, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702692.

Basu:1980:DIP

[Bas80b] S. K. Basu. On development of iterative programs from function specifications. IEEE Transactions on Software Engineering, SE-6(2): 170–182, March/April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702714.

Bastani:1984:PIA

[Bas84]  F. B. Bastani.  Perfor-

mance improvement of abstractions through context dependent transformations. IEEE Transactions on Software Engineering, SE-10(1): 100–116, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010204.

Bass:1985:AUS

[Bas85a] L. J. Bass. An approach to user specification of interactive display interfaces. IEEE Transactions on Software Engineering, SE-11(8):686–698, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702078.

Bassiouni:1985:DCS

[Bas85b] M. A. Bassiouni. Data compression in scientific and statistical databases. IEEE Transactions on Software Engineering, SE-11(10):1047–1058, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701920.

Bastani:1985:EFV

[Bas85c] F. B. Bastani. Experiences with a feedback version development methodology. IEEE Transactions on Software Engineering, SE-11(8):718–723, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702081.

Bastani:1985:UCC

[Bas85d] F. B. Bastani. On the uncertainty in the correctness of computer programs. IEEE Transactions on Software Engineering, SE-11(9):857–864, September 1985. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702105.

Bassiouni:1988:SSD

[Bas88] M. A. Bassiouni. Single-site and distributed optimistic protocols for concurrency control. IEEE Transactions on Software Engineering, 14(8):1071–1080, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7617.

Bidoit:1985:EHF

[BBG^{+}85] M. Bidoit, B. Biebow, M. C. Gaudel, C. Gresse, and G. D. Guiho. Exception handling: Formal specification and systematic program construction. IEEE Transactions on Software Engineering, SE-11(3):242–252, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702000.

Balbo:1986:CQN

[BBG86] G. Balbo, S. C. Bruell, and S. Ghanta. Combining queueing network and generalized stochastic Petri net models for the analysis of some software blocking phenomena. IEEE Transactions on Software Engineering, SE-12(4):561–576, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312904.

Batoory:1988:GED

[BBG^{+}88] D. S. Batoory, J. R. Barnett, J. F. Garza, K. P. Smith, K. Tsukuda, B. C. Twichell, and T. E. Wise. GENESIS: an extensible database management system. IEEE Transactions on Software Engineering, 14(11):1711–1730, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9057.

Bourret:1980:PCE

[BC80] P. Bourret and P. Cros. Presentation and correction of errors in operating system measurements. IEEE Transactions on Software Engineering, SE-6(4):395–398, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702749.

Boksenbaum:1987:CCI

[BCFP87] C. Boksenbaum, M. Cart, J. Ferrie, and J. F. Pons. Concurrent certifications by intervals of timestamps in distributed database systems. IEEE Transactions on Software Engineering, SE-13(4):409–419, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702233.

Bershad:1987:RPC

[BCL^{+}87] B. N. Bershad, D. T. Ching, E. D. Lazowska, J. Sanislo, and M. Schwartz. A remote procedure call facility for interconnecting heterogeneous computer systems. IEEE Transactions on Software Engineering, SE-13(8):880–894, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702307.

Bagrodia:1987:MBA

[BCM87] R. L. Bagrodia, K. M. Chandy, and J. Misra. A messagebased approach to discreteevent simulation. IEEE Transactions on Software Engineering, SE-13(6):654–665, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702271.

Budzinski:1981:CDS

[BD81] R. L. Budzinski and E. S. Davidson. A comparison of dynamic and static virtual memory allocation algorithms. IEEE Transactions on Software Engineering, SE-7 (1):122–131, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702809.

Baroody:1982:IRT

[BD82] A. J. Baroody and D. J. De-Witt. The impact of runtime schema interpretation in a network data model DBMS. IEEE Transactions on Software Engineering, SE-8(2): 123–136, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702920.

Babaoglu:1985:SBN

[BD85] O. Babaoğlu and R. Drummond. Streets of Byzantium: Network architectures for fast reliable broadcasts. IEEE Transactions on Software Engineering, SE-11(6):546–554, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702052.

Bentley:1987:ESD

[BD87] J. L. Bentley and J. A.

Dallen. Exercises in software design. IEEE Transactions on Software Engineering, SE-13(11):1164–1169, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702163.

Balsamo:1989:CTD

[BD89] S. Balsamo and L. Donatiello. On the cycle time distribution in a two-stage cyclic network with blocking. IEEE Transactions on Software Engineering, 15(10):1206–1216, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559769.

Ben-David:1984:ISE

[BDBPLR84] A. Ben-David, M. I. Ben-Porath, J. Z. Loeb, and M. Rich. An industrial software engineering retraining course: Development considerations and lessons learned. IEEE Transactions on Software Engineering, SE-10(6): 748–755, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010303.

Boral:1982:IDM

[BDF^{+}82] H. Boral, D. J. Dewitt, D. Friedland, N. F. Jarrell, and W. K. Wilkinson. Implementation of the database machine DIRECT. IEEE Transactions on Software Engineering, SE-8(6): 533–543, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702986.

Bernard:1989:PDC

[BDHH89] G. Bernard, A. Duda, Y. Haddad, and G. Harrus. Primitives for distributed computing in a heterogeneous local area network environment. IEEE Transactions on Software Engineering, 15(12):1567–1578, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58768.

Budzinski:1981:DAC

[BDMS81] R. L. Budzinski, E. S. Davidson, W. Mayeda, and H. S. Stone. DMIN: An algorithm for computing the optimal dynamic allocation in a virtual memory computer. IEEE Transactions on Software Engineering, SE-7 (1):113–121, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702808.

Braegger:1985:GID

[BDRZ85] R. P. Braegger, A. M. Dudler, J. Rebsamen, and C. A. Zehnder. Gambit: An interactive database design tool for data structures, integrity constraints, and transactions. IEEE Transactions on Software Engineering, SE-11(7):574–583, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702061.

Baiardi:1986:DDC

[BDV86] F. Baiardi, N. De Francesco, and G. Vaglini. Development of a debugger for a concurrent language. IEEE Transactions on Software Engineering, SE-12(4):547–553, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312902.

Bochmann:1989:TAC

[BDZ89] G. V. Bochmann, R. Dssouli, and J. R. Zhao. Trace analysis for conformance and arbitration testing. IEEE Transactions on Software Engineering, 15(11):1347–1356, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=41328.

Bishop:1986:PPD

[BEB^{+}86] P. G. Bishop, D. G. Esp, M. Barnes, P. Humphreys, G. Dahll, and J. Lahti. PODS — a project on diverse software. IEEE Transactions on Software Engineering, SE-12 (9):929–940, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313048.

Beck:1980:GIM

[Bec80] L. L. Beck. A generalized implementation method for relational data sublanguages. IEEE Transactions on Software Engineering, SE-6(2): 152–162, March/April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702712.

Bahaa-El-Din:1989:PAP

[BEDBT89] W. H. Bahaa-El-Din, F. B. Bastani, and J. E. Teng. Performance analysis of periodic and concurrent data structure maintenance strategies for network servers. IEEE Transactions on Software Engineering, 15(12):1526–1536, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58765.

Behrens:1983:MPC

[Beh83]  C. A. Behrens.  Measur-

ing the productivity of computer systems development activities with function points. IEEE Transactions on Software Engineering, SE-9(6): 648–652, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703111.

Benson:1984:CPN

[Ben84]  D. B. Benson.  Counting

paths: Nondeterminism as linear algebra. IEEE Transactions on Software Engineering, SE-10(6):785–794, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010307.

Bendifallah:1987:USM

[Ben87a] S. Bendifallah. Understanding software maintenance work. IEEE Transactions on Software Engineering, SE-13(3):311–323, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702217.

Bengtson:1987:MEO

[Ben87b] N. M. Bengtson. Measuring errors in operational analysis assumptions. IEEE Transactions on Software Engineering, SE-13(7):767–776, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702288.

Bernard:1983:ILC

[Ber83]  G. Bernard.  Interconnec-

tion of local computer networks: Modeling and optimization problems. IEEE Transactions on Software Engineering, SE-9(4):463–470, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703080.

Bersoff:1984:ESC

[Ber84] E. H. Bersoff. Elements of software configuration management. IEEE Transactions on Software Engineering, SE-10 (1):79–87, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010202.

See comments [Ber85c].

Bernstein:1985:LCD

[Ber85a] A. J. Bernstein. A loosely coupled distributed system for reliably storing data. IEEE Transactions on Software Engineering, SE-11(5):446–454, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702034.

Berra:1985:FCC

[Ber85b] P. B. Berra. Foreword computers come and go but data go on forever. IEEE Transactions on Software Engineering, SE-11(7):573, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702060.

Bersoff:1985:C

[Ber85c] E. H. Bersoff. Comments on “Elements of Software Configuration Management”. IEEE Transactions on Software Engineering, SE-11(8):822, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702091.

See [Ber84].

Berry:1987:TFB

[Ber87]  D. M. Berry.  Towards a

formal basis for the formal development method and the ina jo specification language. IEEE Transactions on Software Engineering, SE-13(2):184–201, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702199.

Bevier:1989:KSO

[Bev89]  W. R. Bevier.  Kit: a

study in operating system verification. IEEE Transactions on Software Engineering, 15(11):1382–1396, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41331.

Bloomfield:1986:AFM

[BF86] R. E. Bloomfield and P. K. D. Froome. The application of formal methods to the assessment of high integrity software. IEEE Transactions on Software Engineering, SE-12 (9):988–993, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313053.

Berzins:1985:ADM

[BG85] V. Berzins and M. Gray. Analysis and design in MSG.84: Formalizing functional specifications. IEEE Transactions on Software Engineering, SE-11(8):657–670, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702076.

Bernstein:1983:ACC

[BGL83] P. A. Bernstein, N. Goodman, and Ming-Yee Lai. Analyzing concurrency control algorithms when user and system operations differ. IEEE Transactions on Software Engineering, SE-9(3):233–239, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703050.

Boehm:1984:PVS

[BGS84] B. W. Boehm, T. E. Gray, and T. Seewaldt. Prototyping versus specifying: A multiproject experiment. IEEE Transactions on Software Engineering, SE-10(3):290–303, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010238.

Babad:1980:DES

[BH80] Y. M. Babad and J. A. Hoffer. Data element security and its effects on file segmentation. IEEE Transactions on Software Engineering, SE-6(5): 402–410, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702756.

Brandwajn:1981:SMC

[BH81] A. Brandwajn and J. A. Hernandez. A study of a mechanism for controlling multiprogrammed memory in an interactive system. IEEE Transactions on Software Engineering, SE-7(3):321–331, May/ June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702845.

Basili:1983:ESS

[BH83a] V. R. Basili and D. H. Hutchens. An empirical study of a syntactic complexity family. IEEE Transactions on Software Engineering, SE-9(6): 664–672, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703113.

Bhargava:1983:CMA

[BH83b] B. Bhargava and C. T. Hua.

A causal model for analyzing distributed concurrency control algorithms. IEEE Transactions on Software Engineering, SE-9(4):470–486, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703081.

Bhargava:1982:GER

[Bha82] B. Bhargava. Guest editorial: Reliability issues in distributed systems. IEEE Transactions on Software Engineering, SE-8(3):165–167, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702932.

Black:1987:DAT

[BHJ^{+}87] A. Black, N. Hutchinson, E. Jul, H. Levy, and L. Carter. Distribution and abstract types in Emerald. IEEE Transactions on Software Engineering, SE-13(1):65–76, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702134.

Banerjee:1980:DTQ

[BHN80] J. Banerjee, D. K. Hsiao, and F. K. Ng. Database transformation, query translation, and performance analysis of a new database computer in supporting hierarchical database management. IEEE Transactions on Software Engineering, SE-6(1):91–109, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702698.

Beichter:1984:SSS

[BHP84] F. W. Beichter, O. Herzog, and H. Petzsch. SLAN-4 — a software specification and design language. IEEE Transactions on Software Engineering, SE-10(2):155–162, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010217.

Balsamo:1982:ENT

[BI82] S. Balsamo and G. Iazeolla. An extension of Norton’s theorem for queueing networks. IEEE Transactions on Software Engineering, SE-8(4):298–305, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702953.

Balsamo:1985:PFS

[BI85] S. Balsamo and G. Iazeolla. Product-form synthesis of queueing networks. IEEE Transactions on Software Engineering, SE-11(2):194–199, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1701987.

Birman:1985:IFT

[BJRE85] K. P. Birman, T. A. Joseph, T. Raeuchle, and A. El Abbadi. Implementing faulttolerant distributed objects. IEEE Transactions on Software Engineering, SE-11(6):502–508, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702047.

Banker:1989:SEN

[BK89] R. D. Banker and C. F. Kemerer. Scale economies in new software development. IEEE Transactions on Software Engineering, 15(10):1199–1205, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559768.

Buhr:1989:SCR

[BKHW89] R. J. A. Buhr, G. M. Karam, C. J. Hayes, and C. M. Woodside. Software CAD: a revolutionary approach. IEEE Transactions on Software Engineering, 15(3):235–249, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=21752.

Banerjee:1988:CDC

[BKKG88] J. Banerjee, W. Kim, S. J. Kim, and J. F. Garza. Clustering a DAG for CAD databases. IEEE Transactions on Software Engineering, 14(11):1684–1699, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=9055.

Brilliant:1989:CCP

[BKL89] S. S. Brilliant, J. C. Knight, and N. G. Leveson. The consistent comparison problem in Nversion software. IEEE Transactions on Software Engineering, 15(11):1481–1485, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=41339.

Batini:1984:MDS

[BL84] C. Batini and M. Lenzerini. A methodology for data schema integration in the entity relationship model. IEEE Transactions on Software Engineering, SE-10(6): 650–664, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010294.

Bologna:1986:FRS

[BL86] S. Bologna and N. G. Leveson. Foreword reliability and safety in real-time systems. IEEE Transactions on Software Engineering, SE-12 (9):877–878, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313043.

Baer:1989:IQP

[BL89a] J. L. Baer and Y. B. Lin. Improving quicksort performance with a codeword data structure. IEEE Transactions on Software Engineering, 15(5):622–631, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24711.

Bisant:1989:TPI

[BL89b] D. B. Bisant and J. R. Lyle. A two-person inspection method to improve prog ramming productivity. IEEE Transactions on Software Engineering, 15(10):1294–1304, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559782.

Blikle:1981:DCS

[Bli81] A. J. Blikle. On the development of correct specified programs. IEEE Transactions on Software Engineering, SE-7(5): 519–527, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702878.

Blum:1987:PDI

[Blu87] B. I. Blum. A paradigm for developing information systems. IEEE Transactions on Software Engineering, SE-13(4):432–439, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702235.

Basili:1982:UDP

[BM82] V. R. Basili and H. D. Mills. Understanding and documenting programs. IEEE Transactions on Software Engineering, SE-8(3):270–283, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702942.

Boyle:1984:PRT

[BM84] J. M. Boyle and M. N. Muralidharan. Program reusability through program transformation. IEEE Transactions on Software Engineering, SE-10(5):574–588, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=
5010281.

Bruno:1986:PTP

[BM86] G. Bruno and G. Marchetto.

Process-translatable Petri nets for the rapid prototyping of process control systems. IEEE Transactions on Software Engineering, SE-12(2):346–357, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6312948.

Brown:1989:CMD

[BMD89] D. B. Brown, S. Maghsoodloo, and W. H. Deason. A cost model for determining the optimal number of software test cases. IEEE Transactions on Software Engineering, 15(2):218–221, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21747.

Burton:1987:SEO

[BMOW87] F. W. Burton, D. Maurer, H. G.

Oberhauser, and R. Wilhelm. A space-efficient optimization of call-by-need. IEEE Transactions on Software Engineering, SE-13(6):636–642, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702269.

Barnard:1986:RPD

[BMP86] H. J. Barnard, R. F. Metz, and A. L. Price. A recommended practice for describing software designs: IEEE standards project 1016. IEEE Transactions on Software Engineering, SE-12(2):258–263, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312941.

Bauer:1989:FPC

[BMPP89] F. L. Bauer, B. Moller, H. Partsch, and P. Pepper. Formal program construction by transformations-computer-aided, intuition-guided programming. IEEE Transactions on Software Engineering, 15(2):165–180, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=21743.

Basili:1980:CAF

[BN80a] V. R. Basili and R. E. Noonan. A comparison of the axiomatic and functional models of structured programming. IEEE Transactions on Software Engineering, SE-6(5): 454–465, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702762.

Birrell:1980:UFS

[BN80b] A. D. Birrell and R. M. Needham. A universal file server. IEEE Transactions on Software Engineering, SE-6(5): 450–453, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702761.

Batini:1986:LAD

[BNT86] C. Batini, E. Nardelli, and R. Tamassia. A layout algorithm for data flow diagrams. IEEE Transactions on Software Engineering, SE-12(4):538–546, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312901.

Bobrow:1985:IPA

[Bob85] D. G. Bobrow. If Prolog is the answer, what is the question? or what it takes to support AI programming paradigms. IEEE Transactions on Software Engineering, SE-11(11):1401–1408, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701956.

Boehm:1981:ESS

[Boe81] B. W. Boehm. An experiment in small-scale application software engineering.

IEEE Transactions on Software Engineering, SE-7(5): 482–493, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702874.

Boehm:1984:SEE

[Boe84] B. W. Boehm. Software engineering economics. IEEE Transactions on Software Engineering, SE-10(1):4–21, January/February 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010193.

Bokhari:1981:STA

[Bok81] S. H. Bokhari. A shortest tree algorithm for optimal assignments across space and time in a distributed processor system. IEEE Transactions on Software Engineering, SE-7(6): 583–589, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702890.

Booch:1986:OOD

[Boo86] G. Booch. Object-oriented development. IEEE Transactions on Software Engineering, SE-12(2):211–221, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312937.

Broy:1981:PDF

[BP81] M. Broy and P. Pepper. Program development as a formal activity. IEEE Transactions on Software Engineering, SE-7(1):14–22, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702798.

Beck:1983:SSE

[BP83] L. L. Beck and T. E. Perkins. A survey of software engineering practice: Tools, methods, and results. IEEE Transactions on Software Engineering, SE-9(5): 541–561, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703095.

Branstad:1984:SEP

[BP84a] M. Branstad and P. B. Powell.

Software engineering project standards. IEEE Transactions on Software Engineering, SE-10(1):73–78, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010201.

Buckley:1984:SQA

[BP84b] F. J. Buckley and R. Poston. Software quality assurance. IEEE Transactions on Software Engineering, SE-10 (1):36–41, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010196.

Boehm:1988:UCS

[BP88] B. W. Boehm and P. N. Papaccio. Understanding and controlling software costs. IEEE Transactions on Software Engineering, 14(10):1462–1477, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6191.

Basili:1981:CEQ

[BR81] V. R. Basili and R. W. Reiter. A controlled experiment quantitatively comparing software development approaches. IEEE Transactions on Software Engineering, SE-7(3):299–320, May/ June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702844.

Biswas:1987:ADF

[BR87a] S. Biswas and V. Rajaraman. An algorithm to decide feasibility of linear integer constraints occurring in decision tables. IEEE Transactions on Software Engineering, SE-13(12):1340–1347, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702183.

Burns:1987:TCC

[BR87b] J. E. Burns and E. L. Robertson. Two complementary course sequences on the design and implementation of software products. IEEE Transactions on Software Engineering, SE-13(11):1170–1175, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702164.

Basili:1988:TPT

[BR88] V. R. Basili and H. D. Rombach. The TAME project: towards improvement-oriented software environments. IEEE Transactions on Software Engineering, 14(6):758–773, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6156.

Bhargava:1989:RDD

[BR89a] B. Bhargava and J. Riedl.

The raid distributed database system. IEEE Transactions on Software Engineering, 15(6):726–736, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24726.

Boehm:1989:TWS

[BR89b] B. W. Boehm and R. Ross. Theory-W software project management principles and examples. IEEE Transactions on Software Engineering, 15(7):902–916, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=29489.

Bradley:1988:GSO

[Bra88] J. Bradley. A group-select operation for relational algebra and implications for database machine design. IEEE Transactions on Software Engineering, 14(1):126–129, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4630.

Brender:1980:GB

[Bre80] R. F. Brender. Generation of blisses. IEEE Transactions on Software Engineering, SE-6(6): 553–563, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702782.

Bryant:1981:HLL

[Bry81] R. M. Bryant. On homogeneity and on-line–off-line behavior in M/G/1 queueing systems. IEEE Transactions on Software Engineering, SE-7(3):291–299, May/ June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702843.

Bentley:1980:ASC

[BS80] J. L. Bentley and M. Shaw.

An Alphard specification of a correct and efficient transformation on data structures. IEEE Transactions on Software Engineering, SE-6(6): 572–584, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702784.

Barstow:1981:GEP

[BS81] D. R. Barstow and H. E.

Shrobe. Guest editorial: Programming environments. IEEE Transactions on Software Engineering, SE-7(5): 449–450, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702870.

Bazzichi:1982:AGC

[BS82] F. Bazzichi and I. Spadafora.

An automatic generator for compiler testing. IEEE Transactions on Software Engineering, SE-8(4):343–353, July/ August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702957.

Bryan:1984:MSV

[BS84] W. Bryan and S. Siegel. Making software visible, operational, and maintainable in a small project environment. IEEE Transactions on Software Engineering, SE-10(1): 59–67, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010199.

Blumer:1986:MVA

[BS86] T. P. Blumer and D. P. Sidhu. Mechanical verification and automatic implementation of communication protocols. IEEE Transactions on Software Engineering, SE-12(8):827–843, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312985.

Basili:1987:CES

[BS87] V. R. Basili and R. W. Selby. Comparing the effectiveness of software testing strategies. IEEE Transactions on Software Engineering, SE-13(12):1278–1296, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702179.

Basili:1986:ESE

[BSH86] V. R. Basili, R. W. Selby, and D. H. Hutchens. Experimentation in software engineering. IEEE Transactions on Software Engineering, SE-12(7):733–743, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312975.

Basili:1983:MAD

[BSP83] V. R. Basili, R. W. Selby, and T. Phillips. Metric analysis and data validation across Fortran projects. IEEE Transactions on Software Engineering, SE-9(6): 652–663, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703112.

Brindle:1989:DAT

[BTM89] A. F. Brindle, R. N. Taylor, and D. F. Martin. A debugger for Ada tasking. IEEE Transactions on Software Engineering, 15(3):293–304, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21757.

Belkhouche:1986:DIA

[BU86] B. Belkhouche and J. E. Urban.

Direct implementation of abstract data types from abstract specifications. IEEE Transactions on Software Engineering, SE-12(5):649–661, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312960.

Bunt:1987:F

[Bun87] R. B. Bunt. Foreword. IEEE Transactions on Software Engineering, SE-13(3):362, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702222.

Booth:1980:PAD

[BW80] T. L. Booth and C. A. Wiecek. Performance abstract data types as a tool in software performance analysis and design. IEEE Transactions on Software Engineering, SE-6(2):138–151, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702711.

Basili:1984:MCV

[BW84] V. R. Basili and D. M. Weiss. A methodology for collecting valid software engineering data. IEEE Transactions on Software Engineering, SE-10(6): 728–738, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010301.

Billington:1988:PHL

[BWWH88] J. Billington, G. R. Wheeler, and M. C. Wilbur-Ham. PRO-TEAN: a high-level Petri net tool for the specification and verification of communication protocols. IEEE Transactions on Software Engineering, 14(3):301–316, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4651.

Bastani:1988:CIF

[BYC88] F. B. Bastani, I. L. Yen, and I. R. Chen. A class of inherently fault tolerant distributed programs. IEEE Transactions on Software Engineering, 14(10):1432–1442, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6188.

Baker:1980:CMC

[BZ80] A. L. Baker and S. H. Zweben.

A comparison of measures of control flow complexity. IEEE Transactions on Software Engineering, SE-6(6): 506–512, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702775.

Bergland:1986:GEP

[BZ86] G. D. Bergland and P. Zave. Guest Editors’ prologue special issue on software design methods. IEEE Transactions on Software Engineering, SE-12(2):185–191, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312934.

Chou:1982:LBD

[CA82] T. C. K. Chou and J. A. Abraham. Load balancing in distributed systems. IEEE Transactions on Software Engineering, SE-8(4):401–412, July/ August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702962.

Curry:1984:ETX

[CA84] G. A. Curry and R. M. Ayers. Experience with traits in the Xerox Star workstation. IEEE Transactions on Software Engineering, SE-10(5): 519–527, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010276.

Carlson:1985:TNG

[CA85] C. R. Carlson and A. K.

Arora. Toward the next generation of data modeling tools. IEEE Transactions on Software Engineering, SE-11(9):966–970, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702116.

Cameron:1986:OJ

[Cam86] J. R. Cameron. An overview of JSD. IEEE Transactions on Software Engineering, SE-12(2):222–240, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312938.

Carey:1987:IPO

[Car87]  M. J. Carey.  Improving

the performance of an optimistic concurrency control algorithm through timestamps and versions. IEEE Transactions on Software Engineering, SE-13(6):746–751, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702279.

Cavano:1985:THC

[Cav85] J. P. Cavano. Toward high confidence software. IEEE Transactions on Software Engineering, SE-11(12):1449–1455, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701967.

Chang:1980:MSD

[CC80] Shi-Kuo Chang and Wu-Haung Cheng. A methodology for structured database decomposition. IEEE Transactions on Software Engineering, SE-6(2):205–218, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702717.

Chang:1985:TVO

[CC85a] Shi-Kuo Chang and Wu-Lung Chan. Transformation and verification of office procedures. IEEE Transactions on Software Engineering, SE-11(8):724–734, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702082.

Chor:1985:SER

[CC85b] B. Chor and B. A. Coan. A simple and efficient randomized Byzantine agreement algorithm. IEEE Transactions on Software Engineering, SE-11(6):531–539, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702050.

Chetto:1989:SRE

[CC89a] H. Chetto and M. Chetto. Some results of the earliest deadline scheduling algorithm. IEEE Transactions on Software Engineering, 15(10):1261–1269, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559777.

Chiou:1989:SBU

[CC89b] Guang-Huei Chiou and Wen-Tsuen Chen. Secure broadcasting using the secure lock. IEEE Transactions on Software Engineering, 15(8):929–934, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31350.

Card:1986:ESS

[CCA86] D. N. Card, V. E. Church, and W. W. Agresti. An empirical study of software design practices. IEEE Transactions on Software Engineering, SE-12(2):264–271, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312942.

Cameron:1988:IMP

[CCG^{+}88] E. J. Cameron, D. M. Cohen, B. Gopinath, W. M. Keese, L. Ness, P. Uppaluru, and J. R. Vollaro. The IC* model of parallel computation and programming environment. IEEE Transactions on Software Engineering, 14(3):317–326, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4652.

Ceri:1988:SPR

[CCRDL88] S. Ceri, S. Crespi-Reghizzi, A. Di Maio, and L. A. Lavazza. Software prototyping by relational techniques: experiences with program construction systems. IEEE Transactions on Software Engineering, 14(11):1597–1609, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9048.

Claybrook:1985:DDV

[CCW85] B. G. Claybrook, A. M. Claybrook, and J. Williams. Defining database views as data abstractions. IEEE Transactions on Software Engineering, SE-11(1):3–14, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701893.

Chang:1984:PAC

[CDL84] C. C. Chang, M. W. Du, and R. C. T. Lee. Performance analyses of cartesian product files and random files. IEEE Transactions on Software Engineering, SE-10(1):88–99, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010203.

Currit:1986:CRS

[CDM86] P. A. Currit, M. Dyer, and H. D. Mills. Certifying the reliability of software. IEEE Transactions on Software Engineering, SE-12(1):3–11, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312914.

See correction [CDM89].

Curritt:1989:CCR

[CDM89] P. A. Curritt, M. Dyer, and H. D. Mills. Correction to ‘Certifying the reliability of software’ (Jan. 1986 3–11). IEEE Transactions on Software Engineering, 15(3):362–??, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21765.

See [CDM86].

Cornell:1986:MCT

[CDY86] D. W. Cornell, D. M. Dias, and P. S. Yu. On multisystem coupling through function request shipping. IEEE Transactions on Software Engineering, SE-12(10):1006–1017, October 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313017.

Chang:1980:QPE

[CF80] Ning-San Chang and King-Sun Fu. Query-by-pictorial-example. IEEE Transactions on Software Engineering, SE-6(6): 519–524, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702777.

Chang:1981:EKD

[CF81] Jo-Mei Chang and King-Sun Fu. Extended K-d tree database organization: A dynamic multiattribute clustering method. IEEE Transactions on Software Engineering, SE-7(3):284–290, May/ June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702842.

Christodoulakis:1984:DCM

[CF84] S. Christodoulakis and C. Faloutsos. Design considerations for a message file server. IEEE Transactions on Software Engineering, SE-10(2): 201–210, March/April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010223.

Ceri:1985:TSR

[CG85a] S. Ceri and G. Gottlob. Translating SQL into relational algebra: Optimization, semantics, and equivalence of SQL queries. IEEE Transactions on Software Engineering, SE-11(4):324–345, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702016.

Chan:1985:IDR

[CG85b] A. Chan and R. Gray. Implementing distributed read-only transactions. IEEE Transactions on Software Engineering, SE-11(2):205–212, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701989.

Comer:1985:FHL

[CG85c] D. E. Comer and N. H. Gehani.

Flex: A high-level language for specifying customized microprocessors. IEEE Transactions on Software Engineering, SE-11(4):387–396, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702021.

Conway:1986:DAC

[CG86] A. E. Conway and N. D. Georganas. Decomposition and aggregation by class in closed queueing networks. IEEE Transactions on Software Engineering, SE-12(10): 1025–1040, October 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313019.

Coutant:1983:MPB

[CGH83] C. A. Coutant, R. E. Griswold, and D. R. Hanson. Measuring the performance and behavior of icon programs. IEEE Transactions on Software Engineering, SE-9(1):93–103, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703016.

Coffman:1981:ONC

[CGP81] E. G. Coffman, E. Gelenbe, and B. Plateau. Optimization of the number of copies in a distributed data base. IEEE Transactions on Software Engineering, SE-7(1):78– 84, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702804.

Card:1987:ESE

[CGP87] D. N. Card, F. E. Mc Garry, and G. T. Page. Evaluating software engineering technologies. IEEE Transactions on Software Engineering, SE-13(7):845–851, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702295.

Cmelik:1989:EMP

[CGR89] R. F. Cmelik, N. H. Gehani, and W. D. Roome. Experience with multiple processor versions of Concurrent C. IEEE Transactions on Software Engineering, 15(3):335–344, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21761.

Ceri:1989:EDA

[CGW89] S. Ceri, G. Gottlob, and G. Wiederhold. Efficient database access from Prolog. IEEE Transactions on Software Engineering, 15 (2):153–164, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21742.

Chusho:1981:PAP

[CH81] T. Chusho and T. Hayashi.

Performance analyses of paging algorithms for compilation of a highly modularized program. IEEE Transactions on Software Engineering, SE-7(2):248–254, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702831.

Cooper:1988:PAC

[CH88] R. C. B. Cooper and K. G. Hamilton. Preserving abstraction in concurrent programming. IEEE Transactions on Software Engineering, 14(2):258–263, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4643.

Chandrasekaran:1980:GES

[Cha80] B. Chandrasekaran. Guest editorial: Special collection on program testing. IEEE Transactions on Software Engineering, SE-6(3):233–235, May/ June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702724.

Chang:1982:EAD

[Cha82] E. J. H. Chang. Echo algorithms: Depth parallel operations on general graphs. IEEE Transactions on Software Engineering, SE-8(4): 391–401, July/August 1982.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702961.

Cheung:1980:UOS

[Che80]  R. C. Cheung.  A user-

oriented software reliability model. IEEE Transactions on Software Engineering, SE-6 (2):118–125, March/April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702709.

Cheung:1983:GTT

[Che83] To-Yat Cheung. Graph traversal techniques and the maximum flow problem in distributed computation. IEEE Transactions on Software Engineering, SE-9(4):504–512, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703083.

Cheatham:1984:RTP

[Che84] T. E. Cheatham. Reusability through program transformations. IEEE Transactions on Software Engineering, SE-10(5):589–594, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010282.

Cheung:1986:PMP

[Che86] T. Y. Cheung. On the projection method for protocol verification. IEEE Transactions on Software Engineering, SE-12(11):1088–1089, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312998.

See [Hir90].

Chellappa:1987:NPP

[Che87] M. Chellappa. Nontraversible paths in a program. IEEE Transactions on Software Engineering, SE-13(6):751–756, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702280.

Chi:1985:FSU

[Chi85] U. H. Chi. Formal specification of user interfaces: A comparison and evaluation of four axiomatic approaches. IEEE Transactions on Software Engineering, SE-11(8):671–685, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702077.

See comments [Ale88].

Cabrera:1988:UPC

[CHKH88] L. F. Cabrera, E. Hunter, M. J. Karels, and D. A. Hosher. User-process communication performance in networks of computers. IEEE Transactions on Software Engineering, 14(1):38–53, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4621.

Chowdhury:1989:MRD

[Cho89] S. Chowdhury. The mean resequencing delay for M/H_{K} /∞ systems. IEEE Transactions on Software Engineering, 15(12):1633–1638, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58774.

Coleman:1981:MSD

[CHP81] D. Coleman, J. W. Hughes, and M. S. Powell. A method for the syntax directed design of multiprograms. IEEE Transactions on Software Engineering, SE-7(2):189–196, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702825.

Clarke:1982:CLD

[CHR82] L. A. Clarke, J. Hassell, and D. J. Richardson. A close look at domain testing. IEEE Transactions on Software Engineering, SE-8(4):380–390, July/ August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702960.

Chusho:1987:TDS

[Chu87] T. Chusho. Test data selection and quality estimation based on the concept of essential branches for path testing. IEEE Transactions on Software Engineering, SE-13(5):509–517, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702251.

Cidon:1989:EDK

[Cid89] I. Cidon. An efficient distributed knot detection algorithm. IEEE Transactions on Software Engineering, 15(5):644–649, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24714.

Cheng:1988:NMI

[CIK88] Y. Cheng, S. S. Iyengara, and R. L. Kashyap. A new method of image compression using irreducible covers of maximal rectangles. IEEE Transactions on Software Engineering, 14(5):651–658, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6142.

Chang:1985:IPL

[CJL^{+}85] Shi-Kuo Chang, E. Jungert, S. Levialdi, G. Tortora, and T. Ichikawa. An image processing language with icon-assisted navigation. IEEE Transactions on Software Engineering, SE-11(8):811–819, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702089.

Cidon:1987:LDD

[CJS87] I. Cidon, J. M. Jaffe, and M. Sidi. Local distributed deadlock detection by cycle detection and clustering. IEEE Transactions on Software Engineering, SE-13(1):3–14, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702127.

Cohen:1985:ESP

[CK85] J. Cohen and S. Kolodner. Estimating the speedup in parallel parsing. IEEE Transactions on Software Engineering, SE-11(1):114–124, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701903.

Casavant:1988:ERS

[CK88a] T. L. Casavant and J. G.

Kuhl. Effects of response and stability on scheduling in distributed computing systems. IEEE Transactions on Software Engineering, 14(11): 1578–1588, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9046.

Casavant:1988:TSG

[CK88b] T. L. Casavant and J. G. Kuhl. A taxonomy of scheduling in general-purpose distributed computing systems. IEEE Transactions on Software Engineering, 14(2):141–154, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4634.

Cellary:1988:OCO

[CKM88] W. Cellary, Z. Krolikowski, and T. Morzy. Other comments on “Optimization algorithms for distributed queries” by P. M. G. Apers. IEEE Transactions on Software Engineering, 14(4):439–441, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4666. See

[AHY83].

Coffman:1985:AOF

[CKS85] E. G. Coffman, T. T. Kadota, and L. A. Shepp. On the asymptotic optimality of first-fit storage allocation. IEEE Transactions on Software Engineering, SE-11 (2):235–239, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701993.

Choudhary:1989:CMP

[CKST89a] A. N. Choudhary, W. H.

Kohler, J. A. Stankovic, and D. Towsley. Correction to “A modified priority based probe algorithm for distributed deadlock detection and resolution’ (A. N. Choudhary et al.). IEEE Transactions on Software Engineering, 15(12):1644, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58776.

See [CKST89b].

Choudhary:1989:MPB

[CKST89b] A. N. Choudhary, W. H.

Kohler, J. A. Stankovic, and D. Towsley. A modified priority based probe algorithm for distributed deadlock detection and resolution. IEEE Transactions on Software Engineering, 15(1):10–17, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21721.

See correction [CKST89a].

Cook:1983:STA

[CL83] R. P. Cook and T. J. Leblanc. A symbol table abstraction to implement languages with explicit scope control. IEEE Transactions on Software Engineering, SE-9(1):8–12, January/February 1983. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1703006.

Chen:1985:OAP

[CL85] A. L. P. Chen and V. O. K. Li. An optimal algorithm for processing distributed star queries. IEEE Transactions on Software Engineering, SE-11 (10):1097–1107, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701925.

Chow:1988:PIP

[CL88] C. H. Chow and S. S. Lam. PROSPEC: an interactive programming environment for designing and verifying communication protocols. IEEE Transactions on Software Engineering, 14(3):327–338, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4653.

Chen:1989:OJF

[CL89] J. S. J. Chen and V. O. K.

Li. Optimizing joins in fragmented database systems on a broadcast local network. IEEE Transactions on Software Engineering, 15(1):26–38, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21723.

Claybrook:1982:SMS

[Cla82] B. G. Claybrook. A specification method for specifying data and procedural abstractions. IEEE Transactions on Software Engineering, SE-8(5): 449–459, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702975.

Clark:1988:PA

[Cla88] K. L. Clark. PARLOG and its applications. IEEE Transactions on Software Engineering, 14(12):1792–1804, December 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9064.

Chang:1982:SGC

[CLD82] C. C. Chang, R. C. T. Lee, and M. W. Du. Symbolic gray code as a perfect multiattribute hashing scheme for partial match queries. IEEE Transactions on Software Engineering, SE-8(3):235–249, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702940.

Liu:1989:FMS

[cLH89] Lung chun Liu and E. Horowitz. A formal model for software project management. IEEE Transactions on Software Engineering, 15(10):1280–1293, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=559781.

Cheng:1984:UVH

[CLP84] T. T. Cheng, E. D. Lock, and N. S. Prywes. Use of very high level languages and program generation by management professionals. IEEE Transactions on Software Engineering, SE-10(5):552–563, September/October 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5010279.

Choobineh:1988:EDD

[CMNK88] J. Choobineh, M. V. Mannino, J. F. Nunamaker, and B. R. Konsynski. An expert database design system based on analysis of forms. IEEE Transactions on Software Engineering, 14(2):242–253, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4641.

Ceri:1983:DDL

[CNW83] S. Ceri, S. Navathe, and G. Wiederhold. Distribution design of logical database schemas. IEEE Transactions on Software Engineering, SE-9(4):487–504, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703082.

Chin:1982:AIC

[CO82] F. Y. Chin and G. Ozsoyoglu. Auditing and inference control in statistical databases. IEEE Transactions on Software Engineering, SE-8(6): 574–582, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702991.

Comer:1981:PPD

[Com81] D. Comer. Principles of program design induced from experience with small public programs. IEEE Transactions on Software Engineering, SE-7(2):169–174, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702823.

Cook:1980:MLD

[Coo80] R. P. Cook. *MOD — a language for distributed programming. IEEE Transactions on Software Engineering, SE-6(6): 563–571, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702783.

Cooper:1984:SDM

[Coo84]  J. Cooper.  Software de-

velopment management planning. IEEE Transactions on Software Engineering, SE-10 (1):22–26, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010194.

Cottam:1984:RDS

[Cot84] I. D. Cottam. The rigorous development of a system version control program. IEEE Transactions on Software Engineering, SE-10(2):143–154, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010216.

Coulter:1983:SSC

[Cou83] N. S. Coulter. Software science and cognitive psychology. IEEE Transactions on Software Engineering, SE-9(2): 166–171, March/April 1983.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703033.

Chin:1987:ETN

[CP87] Y. H. Chin and Weng-Ling Peng. An evaluation of two new inference control methods. IEEE Transactions on Software Engineering, SE-13(12):1329–1339, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702182.

Clarke:1989:FED

[CPRZ89] L. A. Clarke, A. Podgurski, D. J. Richardson, and S. J. Zeil. A formal evaluation of data flow path selection criteria. IEEE Transactions on Software Engineering, 15(11): 1318–1332, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41326.

Coolahan:1983:TRT

[CR83] J. E. Coolahan and N. Roussopoulos. Timing requirements for time-driven systems using augmented Petri nets. IEEE Transactions on Software Engineering, SE-9(5): 603–616, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703100.

Campbell:1986:ERA

[CR86] R. H. Campbell and B. Randell. Error recovery in asynchronous systems. IEEE Transactions on Software Engineering, SE-12(8):811–826, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312984.

Chin:1987:IBM

[CR87] F. Y. Chin and K. V. S. Ramarao. An information-based model for failure-handling in distributed database systems. IEEE Transactions on Software Engineering, SE-13(4):420–431, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702234.

Cristian:1984:CRP

[Cri84] F. Cristian. Correct and robust programs. IEEE Transactions on Software Engineering, SE-10(2):163–174, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010218.

Cristian:1985:RAF

[Cri85] F. Cristian. A rigorous approach to fault-tolerant programming. IEEE Transactions on Software Engineering, SE-11(1):23–31, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701895.

Carlson:1983:SSD

[CRW83] E. D. Carlson, J. R. Rhyne, and D. L. Weller. Software structure for display management systems. IEEE Transactions on Software Engineering, SE-9(4):385–394, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703073.

Corrigan:1984:CGP

[CS84] N. B. Corrigan and J. D.

Starkey. A concurrent general purpose operator interface. IEEE Transactions on Software Engineering, SE-10(6): 738–748, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010302.

Cabe:1985:STA

[CS85a] T. J. Mc Cabe and G. G. Schulmeyer. System testing aided by structured analysis: A practical experience. IEEE Transactions on Software Engineering, SE-11(9):917–921, September 1985. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702109.

Cline:1985:AAS

[CS85b] C. L. Cline and H. J. Siegel. Augmenting Ada for SIMD parallel processing. IEEE Transactions on Software Engineering, SE-11(9):970–977, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702117.

Cherniavsky:1987:RTA

[CS87a] J. C. Cherniavsky and C. H. Smith. A recursion theoretic approach to program testing. IEEE Transactions on Software Engineering, SE-13(7):777–784, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702289.

Cristian:1987:FSI

[CS87b] F. Cristian and D. Skeen. Foreword special issue on distributed systems. IEEE Transactions on Software Engineering, SE-13(1):1–2, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702125.

Chung:1989:CDC

[CS89a] A. Chung and D. P. Sidhu.

On conditions for defining a closed cover to verify progress for communicating finite state machines. IEEE Transactions on Software Engineering, 15(11):1491–1494, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41341.

Ciaccia:1989:OSR

[CS89b] P. Ciaccia and M. R. Scalas.

Optimization strategies for relational queries. IEEE Transactions on Software Engineering, 15(10):1217–1235, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559773.

Cune:1985:RSR

[CTDS85] B. P. Mc Cune, R. M.

Tong, J. S. Dean, and D. G. Shapiro. RUBRIC: A system for rule-based information retrieval. IEEE Transactions on Software Engineering, SE-11(9):939–945, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702112.

Chen:1989:RPS

[CTT89] S. K. Chen, W. T. Tsai, and M. B. Thuraisingham. Recovery point selection on a reverse binary tree task model. IEEE Transactions on Software Engineering, 15(8):963–976, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31353.

Chang:1989:VLC

[CTYY89] S. K. Chang, M. J. Tauber, B. Yu, and J. S. Yu. A visual language compiler. IEEE Transactions on Software Engineering, 15(5):506–525, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24700.

Curtis:1983:SMG

[Cur83]  B. Curtis.  Software met-

rics: Guest Editor’s introduction. IEEE Transactions on Software Engineering, SE-9(6): 637–638, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703109.

Clark:1981:CVF

[CV81] K. L. Clark and M. H. Van Emden. Consequence verification of flowcharts. IEEE Transactions on Software Engineering, SE-7(1):52–60, January/February 1981. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702802.

Ciminiera:1989:AMM

[CV89] L. Ciminiera and A. Valenzano. Authentication mechanisms in microprocessor-based local area networks. IEEE Transactions on Software Engineering, 15(5):654–658, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24716.

Celentano:1980:SCP

[CVGM80] A. Celentano, P. D. Vigna, C. Ghezzi, and D. Mandrioli. Separate compilation and partial specification in pascal. IEEE Transactions on Software Engineering, SE-6(4): 320–328, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702739.

Chen:1983:FSV

[CY83] Bo-Shoe Chen and R. T. Yeh.

Formal specification and verification of distributed systems. IEEE Transactions on Software Engineering, SE-9(6): 710–722, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703116.

Cornell:1989:OSA

[CY89] D. W. Cornell and P. S. Yu. On optimal site assignment for relations in the distributed database environment. IEEE Transactions on Software Engineering, 15(8):1004–1009, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=31356.

Chang:1988:IID

[CYDA88] S. K. Chang, C. W. Yan, D. C. Dimitroff, and T. Arndt. An intelligent image database system. IEEE Transactions on Software Engineering, 14(5):681–688, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6147.

Concepcion:1988:DFF

[CZ88] A. I. Concepcion and B. P. Zeigler. DEVS formalism: a framework for hierarchical model development. IEEE Transactions on Software Engineering, 14(2):228–241, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4640.

Denning:1987:VMD

[DAH^{+}87] D. E. Denning, S. G. Akl, M. Heckman, T. F. Lunt, M. Morgenstern, P. G. Neumann, and R. R. Schell. Views for multilevel database security. IEEE Transactions on Software Engineering, SE-13 (2):129–140, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702194.

Dalal:1987:DAC

[Dal87] Y. K. Dalal. A distributed algorithm for constructing minimal spanning trees. IEEE Transactions on Software Engineering, SE-13(3):398–405, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702226.

Damm:1988:ML

[Dam88] W. Damm. A microprogramming logic. IEEE Transactions on Software Engineering, 14(5):559–574, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6134.

Dasarathy:1985:TCR

[Das85] B. Dasarathy. Timing constraints of real-time systems: Constructs for expressing them, methods of validating them. IEEE Transactions on Software Engineering, SE-11(1):80–86, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701900.

Das:1987:RIM

[DAS^{+}87] S. K. Das, V. K. Agrawal, D. Sarkar, L. M. Patnaik, and P. S. Goel. Reflexive Incidence Matrx (RIM) representation of Petri nets. IEEE Transactions on Software Engineering, SE-13(6):643–653, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702270.

Deen:1987:DID

[DAT87] S. M. Deen, R. R. Amin, and M. C. Taylor. Data integration in distributed databases. IEEE Transactions on Software Engineering, SE-13(7):860–864, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702297.

Davies:1980:ABE

[Dav80] A. C. Davies. The analogy between electrical networks and flowcharts. IEEE Transactions on Software Engineering, SE-6(4):391–394, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702748.

Davidson:1987:NCP

[Dav87] C. M. Davidson. A note on concurrent programming control. IEEE Transactions on Software Engineering, SE-13(7):865–866, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702298.

Davidson:1988:QR

[Dav88]  C. M. Davidson.  Quick-

sort revisited. IEEE Transactions on Software Engineering, 14(10):1480–1481, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6193.

Davcev:1989:DVS

[Dav89] D. Davcev. A dynamic voting scheme in distributed systems. IEEE Transactions on Software Engineering, 15(1):93–97, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21731.

Dubois:1982:PSI

[DB82] M. Dubois and F. A. Briggs.

Performance of synchronized iterative processes in multiprocessor systems. IEEE Transactions on Software Engineering, SE-8(4):419–431, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702964.

Dunlop:1984:HDL

[DB84] D. D. Dunlop and V. R. Basili. A heuristic for deriving loop functions. IEEE Transactions on Software Engineering, SE-10(3):275–285, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010236.

Doerflinger:1985:MSD

[DB85] C. W. Doerflinger and V. R. Basili. Monitoring software development through dynamic variables. IEEE Transactions on Software Engineering, SE-11(9):978–985, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702118.

Davis:1988:SCA

[DBC88] A. M. Davis, E. H. Bersoff, and E. R. Comer. A strategy for comparing alternative software development life cycle models. IEEE Transactions on Software Engineering, 14(10):1453–1461, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6190.

Dugan:1989:SPN

[DC89] J. Bechta Dugan and G. Ciardo.

Stochastic Petri net analysis of a replicated file system. IEEE Transactions on Software Engineering, 15(4):394–401, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=16600.

Dannenberg:1982:FPV

[DE82] R. B. Dannenberg and G. W.

Ernst. Formal program verification using symbolic execution. IEEE Transactions on Software Engineering, SE-8(1):43–52, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702904.

Deak:1981:TDP

[Dea81] E. Deak. A transformational derivation of a parsing algorithm in a high-level language. IEEE Transactions on Software Engineering, SE-7(1):23–31, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702799.

Denning:1980:WSP

[Den80]  P. J. Denning.  Work-

ing sets past and present. IEEE Transactions on Software Engineering, SE-6(1):64– 84, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702696.

Denning:1987:IDM

[Den87] D. E. Denning. An intrusiondetection model. IEEE Transactions on Software Engineering, SE-13(2):222–232, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702202.

Desai:1989:PCA

[Des89] B. C. Desai. Performance of a composite attribute and join index. IEEE Transactions on Software Engineering, 15(2):142–152, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21741.

DeWitt:1987:CMD

[DFS87] D. J. De Witt, R. Finkel, and M. Solomon. The crystal multicomputer: Design and implementation experience. IEEE Transactions on Software Engineering, SE-13(8):953–966, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702313.

Dayal:1984:VDG

[DH84] U. Dayal and H. Y. Hwang.

View definition and generalization for database integration in a multidatabase system. IEEE Transactions on Software Engineering, SE-10(6): 628–645, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010292.

Demurjian:1988:TBU

[DH88] S. A. Demurjian and D. K.

Hsiao. Towards a better understanding of data models through the multilingual database system. IEEE Transactions on Software Engineering, 14(7):946–958, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=42737.

Dhillon:1986:ICRa

[Dhi86a] B. Dhillon. IASTED conference on reliability and quality control. IEEE Transactions on Software Engineering, SE-12(9):996, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313055.

Dhillon:1986:ICRb

[Dhi86b] B. Dhillon. IASTED conference on reliability and quality control. IEEE Transactions on Software Engineering, SE-12(11):1104, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313003.

Du:1983:SNP

[DHJS83] M. W. Du, T. M. Hsieh, K. F. Jea, and D. W. Shieh. The study of a new perfect hash scheme. IEEE Transactions on Software Engineering, SE-9(3):305–313, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703058.

Devarakonda:1989:PPR

[DI89] M. V. Devarakonda and R. K. Iyer. Predictability of process resource usage: a measurementbased study on UNIX. IEEE Transactions on Software Engineering, 15(12):1579–1586, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58769.

Dias:1989:ICC

[DIRY89] D. M. Dias, B. R. Iyer, J. T. Robinson, and P. S. Yu. Integrated concurrencycoherency controls for multisystem data sharing. IEEE Transactions on Software Engineering, 15(4):437–448, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=16604.

Dhar:1988:DDR

[DJ88] V. Dhar and M. Jarke. Dependence directed reasoning and learning in systems maintenance support. IEEE Transactions on Software Engineering, 14(2):211–227, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4639.

Dolk:1984:KRM

[DK84] D. R. Dolk and B. R. Konsynski. Knowledge representation for model management systems. IEEE Transactions on Software Engineering, SE-10(6): 619–628, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010291.

Davis:1988:SAC

[DL88a] J. S. Davis and R. J. LeBlanc.

A study of the applicability of complexity measures. IEEE Transactions on Software Engineering, 14(9): 1366–1372, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6179.

Dittrich:1988:VSE

[DL88b] K. R. Dittrich and R. A. Lorie. Version support for engineering database systems. IEEE Transactions on Software Engineering, 14(4):429–437, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4664.

Dershowitz:1981:IRP

[DM81] N. Dershowitz and Z. Manna. Inference rules for program annotation. IEEE Transactions on Software Engineering, SE-7(2):207–222, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702827.

Dertouzos:1989:MOS

[DM89] M. L. Dertouzos and A. K. Mok. Multiprocessor online scheduling of hard-real-time tasks. IEEE Transactions on Software Engineering, 15(12):1497–1506, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=58762.

Duran:1984:ERT

[DN84] J. W. Duran and S. C. Ntafos. An evaluation of random testing. IEEE Transactions on Software Engineering, SE-10(4):438–444, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010257.

Draper:1985:SEU

[DN85] S. W. Draper and D. A. Norman. Software engineering for user interfaces. IEEE Transactions on Software Engineering, SE-11(3):252–258, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702001.

Downs:1985:AMS

[Dow85] T. Downs. An approach to the modeling of software testing with some applications. IEEE Transactions on Software Engineering, SE-11(4):375–386, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702020.

Downs:1986:EAM

[Dow86]  T. Downs.  Extensions to

an approach to the modeling of software testing with some performance comparisons. IEEE Transactions on Software Engineering, SE-12 (9):979–987, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313052.

Doyle:1985:ES

[Doy85] J. Doyle. Expert systems and the “myth” of symbolic reasoning. IEEE Transactions on Software Engineering, SE-11(11):1386–1390, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701954.

Dromey:1988:SPD

[Dro88] R. G. Dromey. Systematic program development. IEEE Transactions on Software Engineering, 14(1):12–29, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4619.

DeJonge:1987:TAM

[DTV87] W. De Jonge, A. S. Tanenbaum, and R. P. Van De Riet. Two access methods using compact binary trees. IEEE Transactions on Software Engineering, SE-13(7):799–810, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702291.

Du:1985:FDP

[Du85]  Hung-Chang Du.  On the

file design problem for partial match retrieval. IEEE Transactions on Software Engineering, SE-11(2):213–222, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701990.

Dunham:1986:ESR

[Dun86] J. R. Dunham. Experiments in software reliability: Life-critical applications. IEEE Transactions on Software Engineering, SE-12(1):110–123, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312925.

Francesco:1988:DTS

[DV88] N. De Francesco and G. Vaglini.

Description for a tool specifying and prototyping concurrent programs. IEEE Transactions on Software Engineering, 14(11):1554–1564, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9044.

Duran:1981:CRS

[DW81] J. W. Duran and J. J.

Wiorkowski. Capture-recapture sampling for estimating software error content. IEEE Transactions on Software Engineering, SE-7(1):147–148, January/February 1981. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702812.

Dogac:1989:GES

[DYS89] A. Dogac, B. Yuruten, and S. Spaccapietra. A generalized expert system for database design. IEEE Transactions on Software Engineering, 15(4):479–491, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=16607.

Edel:1988:TGP

[Ede88] M. Edel. The Tinkertoy graphical programming environment. IEEE Transactions on Software Engineering, 14(8):1110–1115, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7621.

Evangelist:1989:MII

[EFK89] M. Evangelist, N. Francez, and S. Katz. Multiparty interactions for interprocess communication and synchronization. IEEE Transactions on Software Engineering, 15(11): 1417–1426, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=41333.

Estrin:1986:SSA

[EFRV86] G. Estrin, R. S. Fenchel, R. R. Razouk, and M. K. Vernon. SARA (System ARchitects Apprentice): Modeling, analysis, and simulation support for design of concurrent systems. IEEE Transactions on Software Engineering, SE-12(2):293–311, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312945.

Eggert:1981:RCI

[Egg81] P. R. Eggert. Runtime checking for ISO Standard Pascal. IEEE Transactions on Software Engineering, SE-7(4): 447–448, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702865.

Eswaran:1981:CFA

[EHS81] K. P. Eswaran, V. C. Hamacher, and G. S. Shedler. Collisionfree access control for computer communication bus networks. IEEE Transactions on Software Engineering, SE-7(6): 574–582, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702889.

Eich:1988:GDL

[Eic88]  M. H. Eich.  Graph di-

rected locking. IEEE Transactions on Software Engineering, 14(2):133–140, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4633.

Eckhardt:1985:TBA

[EL85] D. E. Eckhardt and L. D. Lee. A theoretical basis for the analysis of multiversion software subject to coincident errors. IEEE Transactions on Software Engineering, SE-11(12):1511–1517, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701974.

Erman:1988:AEE

[ELHR88] L. D. Erman, J. S. Lark, and F. Hayes-Roth. ABE: an environment for engineering intelligent systems. IEEE Transactions on Software Engineering, 14(12):1758–1770, December 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=9062.

Ellozy:1981:DLI

[Ell81]  H. A. Ellozy.  The de-

termination of loop invariants for programs with arrays. IEEE Transactions on Software Engineering, SE-7(2): 197–206, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702826.

Eager:1986:ALS

[ELZ86] D. L. Eager, E. D. Lazowska, and J. Zahorjan. Adaptive load sharing in homogeneous distributed systems. IEEE Transactions on Software Engineering, SE-12(5):662–675, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312961.

Ephremides:1982:AHA

[EM82] A. Ephremides and O. A.

Mowafi. Analysis of a hybrid access scheme for buffered users-probabilistic time division. IEEE Transactions on Software Engineering, SE-8 (1):52–61, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702905.

Elmagarmid:1988:DDD

[ESL88] A. K. Elmagarmid, N. Soundararajan, and M. T. Liu. A distributed deadlock detection and resolution algorithm and its correctness proof. IEEE Transactions on Software Engineering, 14(10):1443–1452, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6189.

Estrin:1987:CIN

[Est87] D. Estrin. Controls for interorganization networks. IEEE Transactions on Software Engineering, SE-13(2):249–261, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702204.

Etkin:1989:DLC

[EZ89] J. Etkin and J. A. Zinky. Development life cycle of computer networks: the executable model approach. IEEE Transactions on Software Engineering, 15(9):1078–1089, September 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=31366.

Fagan:1986:ASI

[Fag86] M. E. Fagan. Advances in software inspections. IEEE Transactions on Software Engineering, SE-12(7):744–751, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312976.

Fairley:1987:GEI

[Fai87] R. E. Fairley. Guest Editor’s introduction. IEEE Transactions on Software Engineering, SE-13(11):1141–1142, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702159.

Faloutsos:1988:GCP

[Fal88]  Christos Faloutsos.  Gray

codes for partial match and range queries. IEEE Transactions on Software Engineering, 14(10):1381–1393, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6184.

Fernandez-Baca:1989:AMP

[FB89] D. Fernandez-Baca. Allocating modules to processors in a distributed system. IEEE Transactions on Software Engineering, 15(11):1427–1436, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41334.

Feather:1982:PSA

[Fea82] M. S. Feather. Program specification applied to a text formatter. IEEE Transactions on Software Engineering, SE-8(5): 490–498, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702977.

Feather:1989:CSC

[Fea89] M. S. Feather. Constructing specifications by combining parallel elaborations. IEEE Transactions on Software Engineering, 15(2):198–208, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21745.

Ferguson:1983:WPS

[Fer83] M. J. Ferguson. Weighted processor sharing-results for hyperexponential servers. IEEE Transactions on Software Engineering, SE-9(4):531–535, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703087.

Ferguson:1984:MNM

[Fer84] M. J. Ferguson. Multiaccess in a nonqueueing mailbox environment. IEEE Transactions on Software Engineering, SE-10(3):237–243, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010232.

Ferrari:1986:CIP

[Fer86] D. Ferrari. Considerations on the insularity of performance evaluation. IEEE Transactions on Software Engineering, SE-12(6):678–683, June 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312965.

Fischer:1982:GSD

[FGL82] M. J. Fischer, N. D. Griffeth, and N. A. Lynch. Global states of a distributed system. IEEE Transactions on Software Engineering, SE-8 (3):198–202, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702936.

Fuggetta:1989:SCR

[FGM89] A. Fuggetta, C. Ghezzi, and D. Mandrioli. Some consideration on real-time behavior of concurrent programs. IEEE Transactions on Software Engineering, 15(3):356–359, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21763.

Fickas:1985:ATD

[Fic85] S. F. Fickas. Automating the transformational development of software. IEEE Transactions on Software Engineering, SE-11(11):1268–1277, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701946.

Fisher:1988:OGM

[Fis88] G. Fisher. An overview of a graphical multilanguage applications environment. IEEE Transactions on Software Engineering, 14(6):774–786, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6157.

Faiman:1980:OPC

[FK80] R. N. Faiman and A. A. Koretesoja. An optimizing Pascal compiler. IEEE Transactions on Software Engineering, SE-6(6):512–519, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702776.

Farmer:1985:SDC

[FKM85] D. B. Farmer, R. King, and D. A. Myers. The semantic database constructor. IEEE Transactions on Software Engineering, SE-11(7):583–591, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702062.

Fischer:1980:IRT

[FL80] C. N. Fischer and R. J.

LeBlanc. The implementation of run-time diagnostics in pascal. IEEE Transactions on Software Engineering, SE-6(4):313–319, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702738.

Felgentreu:1987:OSS

[FLS87] K. Felgentreu, W. M. Lippe, and F. H. Simon. Optimizing static scope Lisp by repetitive interpretation of recursive function calls. IEEE Transactions on Software Engineering, SE-13(6):628–635, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702268.

Ford:1985:ADT

[FM85] R. Ford and K. Miller. Abstract data type development and implementation: An example. IEEE Transactions on Software Engineering, SE-11 (10):1033–1037, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701917.

Feldman:1989:VDT

[FM89] M. B. Feldman and M. L.

Moran. Validating a demonstration tool for graphicsassisted debugging of Ada concurrent programs. IEEE Transactions on Software Engineering, 15(3):305–313, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21758.

Florin:1989:NSE

[FN89] G. Florin and S. Natkin. Necessary and sufficient ergodicity condition for open synchronized queueing networks. IEEE Transactions on Software Engineering, 15(4):367–380, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=16598.

Farrag:1987:TGC

[FO87] A. A. Farrag and M. T. Ozsu. Towards a general concurrency control algorithm for database systems. IEEE Transactions on Software Engineering, SE-13(10):1073–1079, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702147.

Fontenot:1989:SCM

[Fon89] M. L. Fontenot. Software congestion, mobile servers, and the hyperbolic model. IEEE Transactions on Software Engineering, 15(8):947–962, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31352.

Foo:1987:ASS

[Foo87] N. Y. Foo. Algebraic specifications as solutions of implementation equations. IEEE Transactions on Software Engineering, SE-13(12):1364–1369, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702185.

Foster:1980:EST

[Fos80] K. A. Foster. Error Sensitive Test Cases Analysis (ESTCA). IEEE Transactions on Software Engineering, SE-6(3):258–264, May/ June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702727.

Francez:1982:ADT

[FR82] N. Francez and M. Rodeh.

Achieving distributed termination without freezing. IEEE Transactions on Software Engineering, SE-8(3): 287–292, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702944.

Freeman:1987:CAD

[Fre87a] P. Freeman. A conceptual analysis of the draco approach to constructing software systems. IEEE Transactions on Software Engineering, SE-13(7):830–844, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702294.

Freeman:1987:EES

[Fre87b] P. Freeman. Essential elements of software engineering education revisited. IEEE Transactions on Software Engineering, SE-13(11):1143–1148, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702160.

Fryer:1985:RAC

[Fry85] M. O. Fryer. Risk assessment of computer controlled systems. IEEE Transactions on Software Engineering, SE-11(1):125–129, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701904.

Finkel:1989:ECS

[FSAC89] R. A. Finkel, M. L. Scott, Y. Artsy, and H. Y. Chang. Experience with charlotte: simplicity and function in a distributed operating system.

IEEE Transactions on Software Engineering, 15(6):676–685, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24721.

Frankl:1988:AFD

[FW88] P. G. Frankl and E. J. Weyuker.

An applicable family of data flow testing criteria. IEEE Transactions on Software Engineering, 14(10):1483–1498, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6194.

Fuchs:1987:CDL

[FWA87] W. K. Fuchs, Kun-Lung Wu, and J. A. Abraham. Companson and diagnosis of large replicated files. IEEE Transactions on Software Engineering, SE-13(1):15–22, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702128.

Ferrari:1983:VVI

[FyY83] D. Ferrari and Yiu yoo Yih.

VSWS: The variable-interval sampled working set policy. IEEE Transactions on Software Engineering, SE-9 (3):299–305, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703057.

Felician:1989:VHT

[FZ89] L. Felician and G. Zalateu. Validating Halstead’s theory for Pascal programs. IEEE Transactions on Software Engineering, 15(12):1630–1632, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=58773.

Gaffney:1984:ENF

[Gaf84] J. E. Gaffney. Estimating the number of faults in code. IEEE Transactions on Software Engineering, SE-10(4): 459–464, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010260.

See comments and corrections [Lip86].

Gaffney:1986:ISD

[Gaf86] J. E. Gaffney. The impact on software development costs of using HOL’s. IEEE Transactions on Software Engineering, SE-12(3):496–499, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312890.

Gait:1985:AAH

[Gai85] J. Gait. An aspect of aesthetics in human–computer communications: Pretty windows. IEEE Transactions on Software Engineering, SE-11(8):714–717, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702080.

Garg:1985:APS

[Gar85] K. Garg. An approach to performance specification of communication protocols using timed Petri nets. IEEE Transactions on Software Engineering, SE-11(10):1216–1225, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701937.

Goel:1985:FSR

[GB85] A. L. Goel and F. B. Bastani. Foreword software reliability. IEEE Transactions on Software Engineering, SE-11(12):1409–1410, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701961.

Goel:1986:FSR

[GB86] A. L. Goel and F. B. Bastani. Foreword software reliability. IEEE Transactions on Software Engineering, SE-12(1):1–2, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312913.

Gligor:1987:DIS

[GCC^{+}87] V. D. Gligor, C. S. Chandersekaran, R. S. Chapman, L. J. Dotterer, M. S. Hetch, Wen-Der Jiang, A. Johri, G. L. Luckenbaugh, and N. Vasudevan. Design and implementation of secure xenix. IEEE Transactions on Software Engineering, SE-13(2):208–221, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702201.

Gligor:1987:NST

[GCJ^{+}87] V. D. Gligor, C. S. Chandersekaran, Wen-Der Jiang, A. Johri, G. L. Luckenbaugh, and L. E. Reich. A new security testing method and its application to the Secure Xenix kernel. IEEE Transactions on Software Engineering, SE-13(2):169–183, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702198.

Gehani:1982:DUM

[Geh82] N. H. Gehani. Databases and units of measure. IEEE Transactions on Software Engineering, SE-8(6):605–611, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702994.

Gehani:1984:BSP

[Geh84] N. H. Gehani. Broadcasting Sequential Processes (BSP). IEEE Transactions on Software Engineering, SE-10(4): 343–351, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010247.

German:1984:MDB

[Ger84] S. M. German. Monitoring for deadlock and blocking in Ada tasking. IEEE Transactions on Software Engineering, SE-10(6):764–777, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010305.

Gordon:1988:HTE

[GF88] A. J. Gordon and R. A. Finkel.

Handling timing errors in distributed programs. IEEE Transactions on Software Engineering, 14(10):1525–1535, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6197.

Gannon:1987:TM

[GHM87] J. D. Gannon, R. G. Hamlet, and H. D. Mills. Theory of modules. IEEE Transactions on Software Engineering, SE-13(7):820–829, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702293.

Ghosh:1985:ASD

[Gho85] S. P. Ghosh. An application of statistical databases in manufacturing testing. IEEE Transactions on Software Engineering, SE-11(7):591–598, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702063.

Ghosh:1986:SRT

[Gho86]  S. P. Ghosh.  Statistical

relational tables for statistical database management. IEEE Transactions on Software Engineering, SE-12(12): 1106–1116, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313006.

Ghosh:1989:NOR

[Gho89] S. P. Ghosh. Numerical operations on a relational database. IEEE Transactions on Software Engineering, 15(5):600–610, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24709.

Gibbons:1987:SGM

[Gib87]  P. B. Gibbons.  A stub

generator for multilanguage RPC in heterogeneous environments. IEEE Transactions on Software Engineering, SE-13(1):77–87, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702135.

Gieser:1982:HMM

[Gie82] J. L. Gieser. On horizontally microprogrammed microarchitecture description techniques. IEEE Transactions on Software Engineering, SE-8(5): 513–525, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702979.

Girling:1987:CCL

[Gir87] C. G. Girling. Covert channels in LAN’s. IEEE Transactions on Software Engineering, SE-13(2):292–296, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702208.

Getz:1983:VHL

[GKS83] S. L. Getz, G. Kalligiannis, and S. R. Schach. A very high-level interactive graphical trace for the Pascal heap. IEEE Transactions on Software Engineering, SE-9(2):179–185, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703035.

Gouda:1985:PLT

[GL85] M. G. Gouda and Hui-Seng Lee. Proving liveness and termination of systolic arrays using communicating finite state machines. IEEE Transactions on Software Engineering, SE-11(10):1240–1251, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701939.

Glass:1981:PSE

[Gla81] R. L. Glass. Persistent software errors. IEEE Transactions on Software Engineering, SE-7(2):162–168, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702822.

Gligor:1984:NDS

[Gli84] V. D. Gligor. A note on denial-of-service in operating systems. IEEE Transactions on Software Engineering, SE-10(3):320–324, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010241.

Gardarin:1982:PCD

[GM82] G. Gardarin and M. Melkanoff.

Proving consistency of database transactions written in extended pascal. IEEE Transactions on Software Engineering, SE-8(4):440–446, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702968.

Griffyth:1985:PMD

[GM85] N. Griffyth and J. A. Miller.

Performance modeling of database recovery protocols. IEEE Transactions on Software Engineering, SE-11(6):564–572, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702054.

Gurbaxani:1987:SHD

[GM87] V. Gurbaxani and H. Mendelson. Software and hardware in data processing budgets. IEEE Transactions on Software Engineering, SE-13(9): 1010–1017, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702323.

Garcia-Molina:1984:DDC

[GMGK84] H. Garcia-Molina, F. Germano, and W. H. Kohler. Debugging a distributed computing system. IEEE Transactions on Software Engineering, SE-10(2):210–219, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010224.

Garcia-Molina:1988:AHA

[GMK88] H. Garcia-Molina and B. Kogan. Achieving high availability in distributed databases. IEEE Transactions on Software Engineering, 14(7):886–896, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=42732.

Gabrielian:1985:QFA

[GMT85] A. Gabrielian, L. P. McNamee, and D. J. Trawick. The qualified function approach to analysis of program behavior and performance. IEEE Transactions on Software Engineering, SE-11(8):758–773, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702085.

Gligor:1985:F

[GN85] V. D. Gligor and P. A. B.

Ng. Foreword. IEEE Transactions on Software Engineering, SE-11(6):497–501, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702046.

Goel:1980:SD

[Goe80] A. L. Goel. A summary of the discussion on “An Analysis of Competing Software Reliability Models”. IEEE Transactions on Software Engineering, SE-6(5):501–502, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=

1702767. See [SW78].

Goel:1985:SRM

[Goe85] A. L. Goel. Software reliability models: Assumptions, limitations, and applicability. IEEE Transactions on Software Engineering, SE-11(12):1411–1423, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701963.

Goguen:1984:PP

[Gog84] J. A. Goguen. Parameterized programming. IEEE Transactions on Software Engineering, SE-10(5):528–543, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010277.

Goldbogen:1981:PFM

[Gol81] G. C. Goldbogen. PRIM: A fast matrix transpose method. IEEE Transactions on Software Engineering, SE-7(2): 255–257, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702832.

Goldberg:1986:KBP

[Gol86] A. T. Goldberg. Knowledgebased programming: A survey of program design and construction techniques. IEEE Transactions on Software Engineering, SE-12(7):752–768, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312977.

Goodwin:1981:WPE

[Goo81]  J. W. Goodwin.  Why

programming environments need dynamic data types. IEEE Transactions on Software Engineering, SE-7(5): 451–457, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702871.

Gourlay:1983:MFI

[Gou83] J. S. Gourlay. A mathematical framework for the investigation of testing. IEEE Transactions on Software Engineering, SE-9(6):686–709, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703115.

Gouda:1984:CCV

[Gou84] M. G. Gouda. Closed covers:

To verify progress for communicating finite state machines. IEEE Transactions on Software Engineering, SE-10(6): 846–855, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010313.

Gehani:1988:RFC

[GR88] N. H. Gehani and W. D.

Roome. Rendezvous facilities: Concurrent C and the Ada language. IEEE Transactions on Software Engineering, 14(11):1546–1553, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9043.

Gray:1986:ADC

[Gra86] J. N. Gray. An approach to decentralized computer systems. IEEE Transactions on Software Engineering, SE-12(6):684–692, June 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312966.

Gray:1988:FSS

[Gra88] D. Gray. The formal specification of a small bookshop information system. IEEE Transactions on Software Engineering, 14(2):263–272, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4644.

Gannon:1978:OAF

[GS78] T. F. Gannon and S. D. Shapiro. An optimal approach to fault tolerant software systems design. IEEE Transactions on Software Engineering, SE-4(5):390–409, September/October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=

1702554. See comments [LLS81].

Gligor:1980:DDD

[GS80] V. D. Gligor and S. H. Shattuck. On deadlock detection in distributed systems.

IEEE Transactions on Software Engineering, SE-6(5): 435–440, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702759.

Gannon:1981:CCL

[GS81] T. F. Gannon and S. D.

Shapiro. Comments on the critique of lee, lloyd, and shrivastava. IEEE Transactions on Software Engineering, SE-7(6): 610–611, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702894.

Giacalone:1988:IEF

[GS88] A. Giacalone and S. A. Smolka.

Integrated environments for formally well-founded design and simulation of concurrent systems. IEEE Transactions on Software Engineering, 14(6):787–802, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6158.

Gudes:1980:DCB

[Gud80] E. Gudes. The design of a cryptography based secure file system. IEEE Transactions on Software Engineering, SE-6(5): 411–420, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702757.

Guttag:1980:NTA

[Gut80] J. Guttag. Notes on type abstraction (version 2). IEEE Transactions on Software Engineering, SE-6(1):13–23, January/February 1980. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702690.

Gupta:1989:CMC

[GWW89] J. P. Gupta, S. C. Winter, and D. R. Wilson. CTDNet — a mechanism for the concurrent execution of lambda graphs. IEEE Transactions on Software Engineering, 15(11): 1357–1367, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=41329.

Gusella:1989:ACS

[GZ89] R. Gusella and S. Zatti. The accuracy of the clock synchronization achieved by TEMPO in Berkeley UNIX 4.3BSD. IEEE Transactions on Software Engineering, 15(7):847–853, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=29484.

Haase:1981:RTB

[Haa81] V. H. Haase. Real-time behavior of programs. IEEE Transactions on Software Engineering, SE-7(5):494–501, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702875.

Hac:1986:DSQ

[Hać86]  A. Hać.  A decomposition

solution to a queueing network model of a distributed file system with dynamic locking. IEEE Transactions on Software Engineering, SE-12(4):521–530, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312899.

Hac:1989:DAP

[Hac89] A. Hac. A distributed algorithm for performance improvement through file replication, file migration, and process migration. IEEE Transactions on Software Engineering, 15(11):1459–1470, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41337.

Hager:1989:SCR

[Hag89] J. A. Hager. Software cost reduction methods in practice. IEEE Transactions on Software Engineering, 15(12): 1638–1644, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58775.

Hardgrave:1980:APB

[Har80]  W. T. Hardgrave.  Ambi-

guity in processing Boolean queries on TDMS tree structures: A study of four different philosophies. IEEE Transactions on Software Engineering, SE-6(4):357–372, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702744.

Harrus:1985:MBB

[Har85] G. Harrus. A model for the basic block protocol of the Cambridge Ring. IEEE Transactions on Software Engineering, SE-11(1):130–136, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701905.

Hartley:1988:ASP

[Har88a] S. J. Hartley. An analysis of some problems in managing virtual memory systems with fast secondary storage devices. IEEE Transactions on Software Engineering, 14(8):1176–1187, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7627.

Hartley:1988:CTP

[Har88b] S. J. Hartley. Compile-time program restructuring in multiprogrammed virtual memory systems. IEEE Transactions on Software Engineering, 14(11):1640–1644, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9051.

Hartzman:1989:DDD

[Har89] C. S. Hartzman. The delay due to dynamic two-phase locking. IEEE Transactions on Software Engineering, 15(1):72–82, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21728.

Hawryszkiewycz:1983:SDM

[Haw83] I. T. Hawryszkiewycz. A semantic design method. IEEE Transactions on Software Engineering, SE-9(4):373–384, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703072.

Hayes:1985:AFS

[Hay85] I. J. Hayes. Applying formal specification to software development in industry. IEEE Transactions on Software Engineering, SE-11 (2):169–178, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701984.

Hayes:1986:SDM

[Hay86]  I. J. Hayes.  Specifica-

tion directed module testing. IEEE Transactions on Software Engineering, SE-12 (1):124–133, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312926.

Hutchens:1985:SSA

[HB85] D. H. Hutchens and V. R. Basili. System structure analysis: Clustering with data bindings. IEEE Transactions on Software Engineering, SE-11(8):749–757, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702084.

Herndon:1988:RBL

[HB88] R. M. Herndon and V. A. Berzins. The realizable benefits of a language prototyping language. IEEE Transactions on Software Engineering, 14(6):803–809, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6159.

Hufnagel:1989:PPV

[HB89] S. P. Hufnagel and J. C.

Browne. Performance properties of vertically partitioned object-oriented systems. IEEE Transactions on Software Engineering, 15(8):935–946, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31351.

Hawthorn:1982:PAA

[HD82] P. B. Hawthorn and D. J.

DeWitt. Performance analysis of alternative database machine architectures. IEEE Transactions on Software Engineering, SE-8(1):61–75, January/February 1982. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702906.

Halici:1989:CCD

[HD89] U. Halici and A. Dogac. Concurrency control in distributed databases through time intervals and short-term locks. IEEE Transactions on Software Engineering, 15(8):994–1003, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31355.

Heninger:1980:SSR

[Hen80]  K. L. Heninger.  Specify-

ing software requirements for complex systems: New techniques and their application. IEEE Transactions on Software Engineering, SE-6(1):2– 13, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702689.

Henderson:1986:FPF

[Hen86]  P. Henderson.  Functional

programming, formal specification, and rapid prototyping. IEEE Transactions on Software Engineering, SE-12 (2):241–250, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312939.

Hsu:1989:MLS

[HFG89] Wei-Chung Hsu, C. N. Fischer, and J. R. Goodman. On the minimization of loads/stores in local register allocation. IEEE Transactions on Software Engineering, 15(10):1252–1260, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=559775.

Horspool:1982:C

[HH82] R. N. Horspool and E. C. R. Hehner. Comments on “A Method for Representing Data Items of Unlimited Length in a Computer Memory”. IEEE Transactions on Software Engineering, SE-8(6): 620–621, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702996.

See [Bab81].

Hecht:1986:SRS

[HH86] H. Hecht and M. Hecht. Software reliability in the system context. IEEE Transactions on Software Engineering, SE-12(1):51–58, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312919.

Hailpern:1989:CTF

[HHR89] B. Hailpern, T. Huynh, and G. Revesz. Comparing two functional programming systems. IEEE Transactions on Software Engineering, 15(5): 532–542, May 1989. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
24702.

Hirakawa:1990:CRP

[Hir90] Y. Hirakawa. Comments, with reply, on “On the projection method for protocol verification” by T.-Y. Cheung. IEEE Transactions on Software Engineering, 16(3):370–371, March 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=48933.

See [Che86].

Henry:1981:SSM

[HK81] S. Henry and D. Kafura. Software structure metrics based on information flow. IEEE Transactions on Software Engineering, SE-7(5): 510–518, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702877.

Hatzopoulos:1984:OSM

[HK84] M. Hatzopoulos and J. G. Kollias. On the optimal selection of multilist database structures. IEEE Transactions on Software Engineering, SE-10(6): 681–687, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010296.

Hudson:1988:CPD

[HK88a] S. E. Hudson and R. King. The cactis project: database support for software environments. IEEE Transactions on Software Engineering, 14(6):709–719, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6152.

Hudson:1988:SFH

[HK88b] S. E. Hudson and R. King. Semantic feedback in the Higgens UIMS. IEEE Transactions on Software Engineering, 14 (8):1188–1206, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7628.

Haigh:1987:EUT

[HKMY87] J. T. Haigh, R. A. Kemmerer, J. Mchugh, and W. D. Young. An experience using two covert channel analysis techniques on a real system design. IEEE Transactions on Software Engineering, SE-13(2):157–168, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702197.

Heidelberger:1988:PCM

[HL88] P. Heidelberger and M. S. Lakshmi. A performance comparison of multimicro and mainframe database architectures. IEEE Transactions on Software Engineering, 14(4):522–531, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4675.

Heitmeyer:1983:ARS

[HM83] C. L. Heitmeyer and J. D. McLean. Abstract requirements specification: A new approach and its application. IEEE Transactions on Software Engineering, SE-9(5): 580–589, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703098.

Horowitz:1984:EVR

[HM84] E. Horowitz and J. B. Munson. An expansive view of reusable software. IEEE Transactions on Software Engineering, SE-10(5):477–487, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        5010270.

Herlihy:1989:TBO

[HM89] M. P. Herlihy and M. S. McKendry. Timestamp-based orphan elimination. IEEE Transactions on Software Engineering, 15(7):825–831, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=29482.

Higashino:1984:ASH

[HMS^{+}84] T. Higashino, M. Mori, Y. Sugiyama, K. Taniguchi, and T. Kasami. An algebraic specification of HDLC procedures and its verification. IEEE Transactions on Software Engineering, SE-10(6):825–836, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010311.

Habermann:1986:GSD

[HN86] A. N. Habermann and D. Notkin.

Gandalf: Software development environments. IEEE Transactions on Software Engineering, SE-12(12):1117–1127, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313007.

Hayward:1983:SAA

[HO83] V. Hayward and A. Osorio. A system to automatically analyze assembled programs. IEEE Transactions on Software Engineering, SE-9(2):210–213, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703039.

Holzmann:1987:APV

[Hol87] G. J. Holzmann. Automated protocol validation in argos: Assertion proving and scatter searching. IEEE Transactions on Software Engineering, SE-13(6):683–696, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702274.

Halpern:1987:MCA

[HOPW87] J. D. Halpern, S. Owre, N. Proctor, and W. F. Wilson. Muse — a computer assisted verification system. IEEE Transactions on Software Engineering, SE-13 (2):151–156, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702196.

Howden:1980:FPT

[How80] W. E. Howden. Functional program testing. IEEE Transactions on Software Engineering, SE-6(2):162–169, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702713.

Howden:1982:WMT

[How82] W. E. Howden. Weak mutation testing and completeness of test sets. IEEE Transactions on Software Engineering, SE-8(4):371–379, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702959.

Howes:1984:MSD

[How84]  N. R. Howes.  Managing

software development projects for maximum productivity. IEEE Transactions on Software Engineering, SE-10(1): 27–35, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010195.

Howden:1985:F

[How85] W. E. Howden. Foreword.

IEEE Transactions on Software Engineering, SE-11(3): 241, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1701999.

Howden:1986:FAP

[How86] W. E. Howden. A functional approach to program testing and analysis. IEEE Transactions on Software Engineering, SE-12(10):997–1005, October 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313016.

Hennell:1980:SAN

[HP80a] M. A. Hennell and J. A. Prudom. A static analysis of the NAG Library. IEEE Transactions on Software Engineering, SE-6(4):329–333, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702740.

Hsia:1980:FDP

[HP80b] Pei Hsia and F. E. Petry. A framework for discipline in programming. IEEE Transactions on Software Engineering, SE-6(2):226–232, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702719.

Harrus:1982:QAR

[HP82] G. Harrus and B. Plateau. Queueing analysis of a reordering issue. IEEE Transactions on Software Engineering, SE-8(2):113–123, March/ April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702919.

Holley:1981:QDF

[HR81] L. H. Holley and B. K. Rosen. Qualified data flow problems. IEEE Transactions on Software Engineering, SE-7(1):60–78, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702803.

Ho:1982:PDD

[HR82] G. S. Ho and C. V. Ramamoorthy. Protocols for deadlock detection in distributed database systems. IEEE Transactions on Software Engineering, SE-8(6): 554–557, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702988.

See comments [JV83].

Hong:1982:MDP

[HS82] Yang-Chang Hong and S. Y. W.

Su. A mechanism for database protection in cellular-logic devices. IEEE Transactions on Software Engineering, SE-8(6): 583–596, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702992.

Hayes:1987:FML

[HS87] R. Hayes and R. D. Schlichting. Facilitating mixed language programming in distributed systems. IEEE Transactions on Software Engineering, SE-13(12):1254–1264, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702177.

Hoffman:1988:TSM

[HS88] D. Hoffman and R. Snodgrass.

Trace specifications: methodology and models. IEEE Transactions on Software Engineering, 14(9):1243–1252, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6168.

Haas:1989:SPN

[HS89] P. J. Haas and G. S. Shedler. Stochastic Petri net representation of discrete event simulations. IEEE Transactions on Software Engineering, 15(4):381–393, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=16599.

Sun:1989:PIQ

[hSKN89] Xian he Sun, N. N. Kamel, and L. M. Ni. Processing implication on queries. IEEE Transactions on Software Engineering, 15(10):1168–1175, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559764.

Hillyer:1986:NVP

[HSN86a] B. K. Hillyer, D. E. Shaw, and A. Nigam. NON-VON’s performance on certain database benchmarks. IEEE Transactions on Software Engineering, SE-12(4):577–583, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312905.

Hura:1986:SDA

[HSN86b] G. S. Hura, H. Singh, and N. K. Nanda. Some design aspects of databases through Petri net modeling. IEEE Transactions on Software Engineering, SE-12(4):505–510, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312897.

Huang:1980:NVR

[Hua80] J. C. Huang. A new verification rule and its applications. IEEE Transactions on Software Engineering, SE-6(5): 480–484, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702764.

Huang:1985:MSP

[Hua85] J. P. Huang. Modeling of software partition for distributed real-time applications. IEEE Transactions on Software Engineering, SE-11 (10):1113–1126, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701927.

Holliday:1987:GTP

[HV87] M. A. Holliday and M. K. Vernon. A generalized timed Petri net model for performance analysis. IEEE Transactions on Software Engineering, SE-13(12):1297–1310, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702180.

Horowitz:1986:SSD

[HW86a] E. Horowitz and R. C. Williamson. SODOS: A software documentation support environment — its use. IEEE Transactions on Software Engineering, SE-12(11):1076–1087, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312997.

Horqwitz:1986:SSD

[HW86b] E. Horqwitz and R. C. Williamson. SODOS: A software documentation support environment — its definition. IEEE Transactions on Software Engineering, SE-12(8):849–859, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312987.

Howes:1989:MAO

[HW89] N. R. Howes and A. C. Weaver. Measurements of Ada overhead in OSI-style communications systems. IEEE Transactions on Software Engineering, 15(12):1507–1517, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58763.

See comments [Kar90].

Haigh:1987:ENV

[HY87] J. T. Haigh and W. D.

Young. Extending the noninterference version of MLS for SAT. IEEE Transactions on Software Engineering, SE-13(2):141–150, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702195.

Herath:1988:DCM

[HYSY88] J. Herath, Y. Yamaguchi, N. Saito, and T. Yuba. Dataflow computing models, languages, and machines for intelligence computations. IEEE Transactions on Software Engineering, 14(12):1805–1828, December 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9065.

Hall:1986:CPS

[HZ86] W. E. Hall and S. H. Zweben.

The cloze procedure and software comprehensibility measurement. IEEE Transactions on Software Engineering, SE-12(5):608–623, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312957.

Ichikawa:1986:ARD

[IH86] T. Ichikawa and M. Hirakawa. ARES: A relational database with the capability of performing flexible interpretation of queries. IEEE Transactions on Software Engineering, SE-12(5):624–634, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312958.

Ibaraki:1988:CTS

[IKK88] T. Ibaraki, T. Kameda, and N. Katoh. Cautious transaction schedulers for database concurrency control. IEEE Transactions on Software Engineering, 14(7):997–1009, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=42740.

Iannino:1984:CSR

[IMOL84] A. Iannino, J. D. Musa, K. Okumoto, and B. Littlewood. Criteria for software reliability model comparisons. IEEE Transactions on Software Engineering, SE-10(6): 687–691, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010297.

Iyer:1985:ESW

[IR85] R. K. Iyer and D. J. Rossetti. Effect of system workload on operating system reliability: A study on IBM 3081. IEEE Transactions on Software Engineering, SE-11(12):1438–1448, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701966.

Iyer:1989:SPD

[IS89] V. R. Iyer and H. A. Sholl.

Sofiware partitioning for distributed, sequential, pipelined applications. IEEE Transactions on Software Engineering, 15(10):1270–1279, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559779.

Ip:1983:SOS

[ISR83] M. Y. L. Ip, L. V. Saxton, and V. V. Raghavan. On the selection of an optimal set of indexes. IEEE Transactions on Software Engineering, SE-9(2):135–143, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703030.

Indurkhya:1986:OPR

[ISXC86] B. Indurkhya, H. S. Stone, and L. Xi-Cheng. Optimal partitioning of randomly generated distributed programs. IEEE Transactions on Software Engineering, SE-12(3):483–495, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312889.

Iyer:1985:HRS

[IV85] R. K. Iyer and P. Velardi. Hardware-related software errors: Measurement and analysis. IEEE Transactions on Software Engineering, SE-11 (2):223–231, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701991.

Jachner:1984:DFA

[JA84] J. Jachner and V. K. Agarwal. Data flow anomaly detection. IEEE Transactions on Software Engineering, SE-10(4):432–437, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010256.

Jain:1987:CMD

[Jai87] H. K. Jain. A comprehensive model for the design of distributed computer systems. IEEE Transactions on Software Engineering, SE-13 (10):1092–1104, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702149.

Jajodia:1987:E

[Jaj87] S. Jajodia. An extension of “Representative Instances and γ-Acyclic Relational Schemes”. IEEE Transactions on Software Engineering, SE-13(9): 1047–1048, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702327.

See [JN84].

Jalote:1989:FRN

[Jal89a] P. Jalote. Functional refinement and nested objects for object-oriented design. IEEE Transactions on Software Engineering, 15(3):264–270, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21754.

Jalote:1989:ROB

[Jal89b] P. Jalote. Resilient objects in broadband networks. IEEE Transactions on Software Engineering, 15(1):68–72, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21727.

Jalote:1989:TCS

[Jal89c] P. Jalote. Testing the completeness of specifications. IEEE Transactions on Software Engineering, 15(5):526–531, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24701.

Jayaraman:1988:SE

[Jay88]  B. Jayaraman.  Semantics

of EqL. IEEE Transactions on Software Engineering, 14(4):472–480, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4670.

Jalote:1986:AAF

[JC86] P. Jalote and R. H. Campbell. Atomic actions for faulttolerance using CSP. IEEE Transactions on Software Engineering, SE-12(1):59–68, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312920.

Joseph:1988:PHL

[JC88] T. Joseph and A. F. Cardenas. PICQUERY: a high level query language for pictorial database management. IEEE Transactions on Software Engineering, 14(5):630–638, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6140.

Jeffery:1987:TSC

[Jef87] D. R. Jeffery. Time-sensitive cost models in the commercial MIS environment. IEEE Transactions on Software Engineering, SE-13(7):852–859, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702296.

Jewell:1985:BEB

[Jew85] W. S. Jewell. Bayesian extensions to a basic model of software reliability. IEEE Transactions on Software Engineering, SE-11(12):1465–1471, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701969.

Jayaraman:1989:ELI

[JG89] B. Jayaraman and G. Gupta.

EQL: the language and its implementation. IEEE Transactions on Software Engineering, 15(6):771–779, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24730.

Jha:1989:APP

[JKC89] R. Jha, J. M. Kamrad, and D. T. Cornhill. Ada program partitioning language: a notion for distributing Ada programs. IEEE Transactions on Software Engineering, 15(3):271–280, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21755.

Jenq:1988:QNM

[JKT88] B. C. Jenq, W. H. Kohler, and D. Towsley. A queueing network model for a distributed database testbed system. IEEE Transactions on Software Engineering, 14(7):908–921, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=42734.

Jain:1987:MVR

[JL87] P. Jain and S. S. Lam. Modeling and verification of realtime protocols for broadcast networks. IEEE Transactions on Software Engineering, SE-13(8):924–937, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702311.

Jajodia:1983:SPP

[JLN83] S. Jajodia, J. Liu, and P. A.

Ng. A scheme of parallel processing for MIMD systems. IEEE Transactions on Software Engineering, SE-9(4): 436–445, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703078.

Jahanian:1986:SAT

[JM86] F. Jahanian and A. K. L.

Mok. Safety analysis of timing properties in real-time systems. IEEE Transactions on Software Engineering, SE-12 (9):890–904, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313045.

Jajodia:1989:PCC

[JM89] S. Jajodia and D. Mutchler. A pessimistic consistency control algorithm for replicated files which achieves high availability. IEEE Transactions on Software Engineering, 15(1):39–46, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21724.

Jard:1988:DVP

[JMG88] C. Jard, J. F. Monin, and R. Groz. Development of veda, a prototyping tool for distributed algorithms. IEEE Transactions on Software Engineering, 14(3):339–352, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4654.

Jajodia:1984:RIG

[JN84] S. Jajodia and P. A. Ng. Representative instances and γ-acyclic relational schemes. IEEE Transactions on Software Engineering, SE-10(6): 614–618, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010290.

See extension [Jaj87].

Jajodia:1983:PEE

[JNS83] S. Jajodia, P. A. Ng, and F. N. Springsteel. The problem of equivalence for entity-relationship diagrams. IEEE Transactions on Software Engineering, SE-9(5): 617–630, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703101.

Joe:1989:SIG

[Joe89]  H. Joe.  Statistical infer-

ence for general-order-statistics and nonhomogeneous-poisson-process software reliability models. IEEE Transactions on Software Engineering, 15(11): 1485–1490, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41340.

Jomier:1981:MMC

[Jom81] G. Jomier. A mathematical model for the comparison of static and dynamic memory allocation in a paged system. IEEE Transactions on Software Engineering, SE-7(4): 375–385, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702858.

Jones:1984:RPS

[Jon84]  T. C. Jones.  Reusabil-

ity in programming: A survey of the state of the art. IEEE Transactions on Software Engineering, SE-10(5): 488–494, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010271.

Jacob:1989:SME

[JP89] R. T. Jacob and I. P. Page. Synthesis of mutual exclusion solutions based on binary semaphores. IEEE Transactions on Software Engineering, 15(5):560–568, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24705.

Johnson:1985:PKB

[JS85] W. L. Johnson and E. Soloway. PROUST: Knowledge-based program understanding. IEEE Transactions on Software Engineering, SE-11(3):267–275, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702003.

Jajodia:1987:CUI

[JS87] S. Jajodia and F. N. Springsteel. Construction of universal instances for loop-free network databases using a joinlike operation. IEEE Transactions on Software Engineering, SE-13(7):811–819, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702292.

Jajodia:1987:F

[JT87] S. Jajodia and S. K. Tripathy. Foreword. IEEE Transactions on Software Engineering, SE-13(8):869–870, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702304.

Jarke:1985:FEN

[JTS^{+}85] M. Jarke, J. A. Tuner, E. A.

Stohr, Y. Vassiliou, N. H. White, and K. Michielsen. A field evaluation of natural language for data retrieval. IEEE Transactions on Software Engineering, SE-11(1):97–114, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701902.

Jagannathan:1983:C

[JV83] J. R. Jagannathan and R. Vasudevan. Comments on “Protocols for Deadlock Detection in Distributed Database Systems”. IEEE Transactions on Software Engineering, SE-9(3):371, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703066.

See [HR82].

Jensen:1985:ESS

[JV85] H. A. Jensen and K. Vairavan. An experimental study of software metrics for realtime software. IEEE Transactions on Software Engineering, SE-11(2):231–234, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701992.

Jacobs:1983:GQE

[JW83] B. E. Jacobs and C. A. Walczak. A generalized queryby-example data manipulation language based on database logic. IEEE Transactions on Software Engineering, SE-9(1):40–57, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703011.

Kasturi:1988:IEI

[KA88] R. Kasturi and J. Alemany. Information extraction from images of paper-based maps. IEEE Transactions on Software Engineering, 14(5):671–675, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6145.

Kameda:1984:PNC

[Kam84]  H. Kameda.  A property

of normalization constants for closed queueing networks. IEEE Transactions on Software Engineering, SE-10(6): 856–857, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010314.

Kameda:1986:GCC

[Kam86]  H. Kameda.  The gains

from computer communication. IEEE Transactions on Software Engineering, SE-12(11):1049–1055, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312994.

Kant:1985:UAA

[Kan85] E. Kant. Understanding and automating algorithm design. IEEE Transactions on Software Engineering, SE-11(11): 1361–1374, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701952.

Karasik:1985:ETT

[Kar85] M. S. Karasik. Environmental testing techniques for software certification. IEEE Transactions on Software Engineering, SE-11(9):934–938, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702111.

Karam:1990:CMA

[Kar90] G. M. Karam. Comments on “Measurements of Ada overhead in OSI-style communications systems”. IEEE Transactions on Software Engineering, 16(12):1435–1439, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=62440.

See [HW89].

Kaudel:1990:CAP

[Kau90] F. Kaudel. Comments on “Allocating programs containing branches and loops within a multiple processor system” by D. Towsley. IEEE Transactions on Software Engineering, 16(4):471–??, April 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=54298.

See [Tow86].

Kant:1981:RPI

[KB81] E. Kant and D. R. Barstow. The refinement paradigm: The interaction of coding and efficiency knowledge in program synthesis. IEEE Transactions on Software Engineering, SE-7(5):458–471, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702872.

Ketabchi:1988:MMC

[KB88] M. A. Ketabchi and V. Berzins. Mathemtical model of composite objects and its application for organizing engineering databases. IEEE Transactions on Software Engineering, 14(1):71–84, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4624.

Kavi:1987:IBP

[KBB87] K. M. Kavi, B. P. Buckles, and U. N. Bhat. Isomorphisms between Petri nets and dataflow graphs. IEEE Transactions on Software Engineering, SE-13(10):1127–1134, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702152.

Khoshafian:1985:ESS

[KBD85] S. N. Khoshafian, D. M. Bates, and D. J. De Witt. Efficient support of statistical operations. IEEE Transactions on Software Engineering, SE-11(10):1058–1070, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701921.

Khayat:1985:C

[KBFL85] M. G. Khayat, W. S. Breger, M. Freiling, and T. G. Lewis. Correction to “A Concurrency Measure”. IEEE Transactions on Software Engineering, SE-11(8):822, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702092.

See [Kha84].

Kim:1988:OIC

[KCB88] W. Kim, H. T. Chou, and J. Banerjee. Operations and implementation of complex objects. IEEE Transactions on Software Engineering, 14(7):985–996, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=42739.

Kao:1988:PQR

[KCL88] M. Kao, N. Cercone, and W. S. Luk. Providing quality responses with natural language interfaces: the null value problem. IEEE Transactions on Software Engineering, 14(7):959–984, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=42738.

Kearns:1983:LRH

[KD83] J. P. Kearns and S. De-Fazio. Locality of reference in hierarchical database systems. IEEE Transactions on Software Engineering, SE-9(2): 128–134, March/April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703029.

Keeman:1987:ESE

[Kee87] W. M. Mc Keeman. Experience with a software engineering project course. IEEE Transactions on Software Engineering, SE-13(11):1182–1192, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702166.

Kemmerer:1985:TFS

[Kem85] R. A. Kemmerer. Testing formal specifications to detect design errors. IEEE Transactions on Software Engineering, SE-11(1):32–43, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701896.

Kernighan:1984:USS

[Ker84] B. W. Kernighan. The Unix system and software reusability. IEEE Transactions on Software Engineering, SE-10(5): 513–518, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010275.

Kessels:1981:SPC

[Kes81]  J. L. W. Kessels.  The

soma: A programming construct for distributed processing. IEEE Transactions on Software Engineering, SE-7(5): 502–509, September/October 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702876.

Kessels:1988:PLP

[Kes88] J. L. W. Kessels. PHILAN: a LAN providing a reliable message service for distributed processing. IEEE Transactions on Software Engineering, 14(10):1424–1431, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6187.

Kent:1988:OSR

[KGM88] J. Kent and H. Garcia-Molina. Optimizing shadow recovery algorithms. IEEE Transactions on Software Engineering, 14(2):155–168, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4635.

Kinsley:1988:EDU

[KH88] K. C. Kinsley and C. E. Hughes. Evaluating database update schemes: a methodology and its applications to distributive systems. IEEE Transactions on Software Engineering, 14(8):1081–1089, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7618.

Khayat:1984:CM

[Kha84] M. G. Khayat. A concurrency measure. IEEE Transactions on Software Engineering, SE-10(6):804–810, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=

5010309. See correction [KBFL85].

Kumar:1986:DPR

[KHR86] V. K. P. Kumar, S. Hariri, and C. S. Raghavendra. Distributed program reliability analysis. IEEE Transactions on Software Engineering, SE-12(1):42–50, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312918.

Kim:1982:AMC

[Kim82]  K. H. Kim.  Approaches

to mechanization of the conversation scheme based on monitors. IEEE Transactions on Software Engineering, SE-8(3):189–197, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702935.

Kimura:1986:SEA

[Kim86]  G. D. Kimura.  A struc-

ture editor for abstract document objects. IEEE Transactions on Software Engineering, SE-12(3):417–435, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312883.

Kim:1988:PTC

[Kim88] K. H. Kim. Programmertransparent coordination of recovering concurrent processes: philosophy and rules for efficient implementation. IEEE Transactions on Software Engineering, 14(6):810–821, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6160.

Kim:1989:AEE

[Kim89] K. H. Kim. An approach to experimental evaluation of realtime fault-tolerant distributed computing schemes. IEEE Transactions on Software Engineering, 15(6):715–725, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24725.

King:1980:PCI

[Kin80] J. C. King. Program correctness: on inductive assertion methods. IEEE Transactions on Software Engineering, SE-6(5):465–479, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702763.

Koch:1983:ORT

[KK83] H. S. Koch and P. Kubat.

Optimal release time of computer software. IEEE Transactions on Software Engineering, SE-9(3):323–327, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703060.

Karimi:1988:ASD

[KK88] J. Karimi and B. R. Konsynski. An automated software design assistant. IEEE Transactions on Software Engineering, 14(2):194–210, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4638.

Kritzinger:1980:ISO

[KKT80a] P. S. Kritzinger, A. E. Krzesinski, and P. Teunissen. Incorporating system overhead in queuing network models. IEEE Transactions on Software Engineering, SE-6(4): 381–390, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702746.

Kritzinger:1980:QMT

[KKT80b] P. S. Kritzinger, A. E. Krzesinski, and P. Teunissen. A queuing model of a time-sliced priority-driven task dispatching algorithm. IEEE Transactions on Software Engineering, SE-6(2):219–225, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702718.

Katz:1984:DSV

[KL84] R. H. Katz and T. J. Lehman. Database support for versions and alternatives of large design files. IEEE Transactions on Software Engineering, SE-10(2):191–200, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010222.

Knight:1986:EEA

[KL86] J. C. Knight and N. G. Leveson. An experimental evaluation of the assumption of independence in multiversion programming. IEEE Transactions on Software Engineering, SE-12(1):96–109, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312924.

Kain:1987:ACC

[KL87] R. Y. Kain and C. E.

Landwehr. On access checking in capability-based systems. IEEE Transactions on Software Engineering, SE-13 (2):202–207, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702200.

Kifer:1988:SIL

[KL88] M. Kifer and E. L. Lozinskii.

SYGRAF: implementing logic programs in a database style. IEEE Transactions on Software Engineering, 14(7):922–935, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=42735.

Kalantari:1983:DSA

[KM83] I. Kalantari and G. McDonald. A data structure and an algorithm for the nearest point problem. IEEE Transactions on Software Engineering, SE-9(5): 631–634, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703102.

Kramer:1985:DCD

[KM85] J. Kramer and J. Magee. Dynamic configuration for distributed systems. IEEE Transactions on Software Engineering, SE-11(4):424–436, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702024.

Kamath:1987:IFS

[KM87] Y. H. Kamath and M. M. Matthews. Implementation of an FP-Shell. IEEE Transactions on Software Engineering, SE-13(5):532–539, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702253.

Keirsey:1985:AVC

[KMB^{+}85] D. Keirsey, J. Mitchell, B. Bullock, T. Nussmeir, and D. Y. Tseng. Autonomous vehicle control using al techniques. IEEE Transactions on Software Engineering, SE-11 (9):986–992, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702119.

Kearns:1982:PEC

[KMS82] J. P. Kearns, C. J. Meier, and M. L. Soffa. The performance evaluation of control implementations. IEEE Transactions on Software Engineering, SE-8(2):89–96, March/ April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702916.

Kashyap:1983:NSM

[KO83] R. L. Kashyap and B. J.

Oommen. The noisy substring matching problem. IEEE Transactions on Software Engineering, SE-9(3):365–370, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703065.

Korel:1988:PPE

[Kor88] B. Korel. PELAS-program error-locating assistant system. IEEE Transactions on Software Engineering, 14(9): 1253–1260, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6169.

Kontos:1982:IMG

[KP82] J. Kontos and G. K. Papakonstantinou. The interpretation of meta grammars describing syntax-directed interpreters using an attribute grammar interpreter. IEEE Transactions on Software Engineering, SE-8(4):435–436, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702966.

Kafura:1987:USC

[KR87] D. Kafura and G. R. Reddy. The use of software complexity metrics in software maintenance. IEEE Transactions on Software Engineering, SE-13(3):335–343, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702219.

Kluge:1983:PNM

[KS83] W. E. Kluge and H. Schlutter. Petri net models for the evaluation of applicative programs based on λexpressions. IEEE Transactions on Software Engineering, SE-9(4):415–427, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703076.

Kligerman:1986:RTE

[KS86] E. Kligerman and A. D. Stoyenko. Real-time euclid: A language for reliable realtime systems. IEEE Transactions on Software Engineering, SE-12(9):941–949, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313049.

Katz:1987:CDA

[KS87] S. Katz and O. Shmueli. Cooperative distributed algorithms for dynamic cycle prevention. IEEE Transactions on Software Engineering, SE-13(5):540–552, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702254.

Kumar:1989:PCO

[KS89] A. Kumar and M. Stonebraker. Performance considerations for an operating system transaction manager. IEEE Transactions on Software Engineering, 15(6):705–714, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24724.

Karam:1989:CRA

[KSB89] G. M. Karam, C. M. Stanczyk, and G. W. Bond. Critical races in Ada programs. IEEE Transactions on Software Engineering, 15(11):1471–1480, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/

                                        stamp.jsp?arnumber=41338.

See comments [MO90].

Koo:1987:CRR

[KT87] R. Koo and S. Toueg. Checkpointing and rollback-recovery for distributed systems. IEEE Transactions on Software Engineering, SE-13(1):23–31, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702129.

Knight:1987:IUA

[KU87] J. C. Knight and J. I. A. Urquhart. On the implementation and use of Ada on faulttolerant distributed systems. IEEE Transactions on Software Engineering, SE-13(5):553–563, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702255.

Kundu:1982:CLM

[Kun82] S. Kundu. On a class of linear maps for data compression. IEEE Transactions on Software Engineering, SE-8(5): 530–532, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702981.

Kundu:1986:CRT

[Kun86] S. Kundu. The call-return tree and its application to program performance analysis. IEEE Transactions on Software Engineering, SE-12(11): 1096–1098, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313000.

Kung:1989:CMC

[Kun89] C. H. Kung. Conceptual modeling in the context of development. IEEE Transactions on Software Engineering, 15(10):1176–1187, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559766.

Kusiak:1989:ESO

[Kus89] A. Kusiak. Expert systems and optimization. IEEE Transactions on Software Engineering, 15(8):1017–1020, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31358.

Kwong:1982:NMC

[KW82] Yat-Sang Kwong and D. Wood.

A new method for concurrency in B-trees. IEEE Transactions on Software Engineering, SE-8(3):211–222, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702938.

Keller:1985:UER

[KW85a] A. M. Keller and M. W. Wilkins. On the use of an extended relational model to handle changing incomplete information. IEEE Transactions on Software Engineering, SE-11(7):620–633, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702066.

Konstam:1985:SSA

[KW85b] A. H. Konstam and D. E. Wood. Software science applied to APL. IEEE Transactions on Software Engineering, SE-11(10):994–1000, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701914.

Kruskal:1985:AIS

[KW85c] C. P. Kruskal and A. Weiss. Allocating independent subtasks on parallel processors. IEEE Transactions on Software Engineering, SE-11(10): 1001–1016, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701915.

Kakuda:1988:AEA

[KWN88] Y. Kakuda, Y. Wakahara, and M. Norigoe. An acyclic expansion algorithm for fast protocol validation. IEEE Transactions on Software Engineering, 14(8):1059–1070, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7616.

Lint:1981:CID

[LA81] B. Lint and T. Agerwala. Communication issues in the design and analysis of parallel algorithms. IEEE Transactions on Software Engineering, SE-7(2):174–188, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702824.

Larsen:1983:CHT

[LA83] R. L. Larsen and A. K.

Agrawala. Control of a heterogeneous two-server exponential queueing system. IEEE Transactions on Software Engineering, SE-9(4):522–526, July/ August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703085.

Lakshman:1986:EDC

[LA86] T. V. Lakshman and A. K.

Agrawala. Efficient decentralized consensus protocols.

IEEE Transactions on Software Engineering, SE-12(5):600–607, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312956.

Lam:1988:CPC

[Lam88a] S. S. Lam. Correction to ‘Protocol conversion’. IEEE Transactions on Software Engineering, 14(9):1376–??, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6181. See

[Lam88b].

Lam:1988:PC

[Lam88b] S. S. Lam. Protocol conversion. IEEE Transactions on Software Engineering, 14(3):353–362, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4655. See

correction [Lam88a].

Laprie:1984:DES

[Lap84]  J. C. Laprie.  Depend-

ability evaluation of software systems in operation. IEEE Transactions on Software Engineering, SE-10(6): 701–714, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010299.

Lausen:1988:MAB

[Lau88] G. Lausen. Modeling and analysis of the behavior of information systems. IEEE Transactions on Software Engineering, 14(11):1610–1620, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9049.

Lilien:1984:SBV

[LB84] L. Lilien and B. Bhargava.

A scheme for batch verification of integrity assertions in a database system. IEEE Transactions on Software Engineering, SE-10(6): 664–680, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010295.

Lilien:1985:DIB

[LB85] L. Lilien and B. Bhargava.

Database integrity block construct: Concepts and design issues. IEEE Transactions on Software Engineering, SE-11(9):865–885, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702106.

Leu:1987:MTP

[LB87] Pei-Jyun Leu and B. Bhargava. Multidimensional timestamp protocols for concurrency control. IEEE Transactions on Software Engineering, SE-13(12):1238–1253, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702176.

Leu:1988:CTP

[LB88] P. J. Leu and B. Bhargava. Clarification of two phase locking in concurrent transaction processing. IEEE Transactions on Software Engineering, 14(1):122–125, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4629.

Luqi:1988:PLR

[LBY88] Luqi, V. Berzins, and R. Yeh. A prototyping language for realtime software. IEEE Transactions on Software Engineering, 14(10):1409–1423, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6186.

Leung:1984:ELB

[LC84] C. H. C. Leung and Q. H. Choo. On the execution of large batch programs in unreliable computing systems. IEEE Transactions on Software Engineering, SE-10(4):444–450, July/ August 1984. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010258.

Liestman:1986:FTS

[LC86] A. L. Liestman and R. H.

Campbell. A fault-tolerant scheduling problem. IEEE Transactions on Software Engineering, SE-12(11):1089–1095, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6312999.

Li:1987:ESS

[LC87] H. F. Li and W. K. Cheung.

An empirical study of software metrics. IEEE Transactions on Software Engineering, SE-13(6):697–708, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702275.

Lew:1988:SCI

[LDF88] K. S. Lew, T. S. Dillon, and K. E. Forward. Software complexity and its impact on software reliability. IEEE Transactions on Software Engineering, 14(11):1645–1655, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9052.

LeBoudec:1987:IIT

[Le 87] J. Y. Le Boudec. Interinput and interoutput time distribution in classical productform networks. IEEE Transactions on Software Engineering, SE-13(6):756–759, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702281.

Lee:1987:ITAa

[Lee87a] T. T. Lee. An informationtheoretic analysis of relational databases — part i: Data dependencies and information metric. IEEE Transactions on Software Engineering, SE-13(10):1049–1061, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702145.

Lee:1987:ITAb

[Lee87b] T. T. Lee. An informationtheoretic analysis of relational databases — Part II: Information structures of database schemas. IEEE Transactions on Software Engineering, SE-13(10):1062–1072, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702146.

Lehman:1989:ECT

[Leh89]  J. A. Lehman.  An em-

pirical comparison of textual and graphical data structure documentation for Cobol programs. IEEE Transactions on Software Engineering, 15(9):1131–1135, September 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31370.

Leinbaugh:1980:GRT

[Lei80]  D. W. Leinbaugh.  Guar-

anteed response times in a hard-real-time environment. IEEE Transactions on Software Engineering, SE-6(1):85– 91, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702697.

Leinbaugh:1984:SHL

[Lei84] D. W. Leinbaugh. Selectors:

High-level resource schedulers. IEEE Transactions on Software Engineering, SE-10(6): 810–825, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010310.

Lenders:1989:DCS

[Len89] P. M. Lenders. Distributed computing with single readsingle write variables. IEEE Transactions on Software Engineering, 15(5):569–574, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24706.

Leung:1983:ASS

[Leu83] C. H. C. Leung. Analysis of secondary storage fragmentation. IEEE Transactions on Software Engineering, SE-9(1):87–93, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703015.

Leung:1986:DSF

[Leu86] C. H. C. Leung. Dynamic storage fragmentation and file deterioration. IEEE Transactions on Software Engineering, SE-12(3):436–441, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312884.

Levy:1986:MME

[Lev86] L. S. Levy. A metaprogramming method and its economic justification. IEEE Transactions on Software Engineering, SE-12(2):272–277, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312943.

Lau:1983:PRM

[LF83] E. J. Lau and D. Ferrari. Program restructuring in a multilevel virtual memory. IEEE Transactions on Software Engineering, SE-9 (1):69–79, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703013.

LeBlanc:1984:SSC

[LF84] R. J. Le Blanc and C. N.

Fischer. A simple separate compilation mechanism for block-structured languages. IEEE Transactions on Software Engineering, SE-10(3): 221–227, May/June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010230.

Lanergan:1984:SER

[LG84] R. G. Lanergan and C. A.

Grasso. Software engineering with reusable designs and code. IEEE Transactions on Software Engineering, SE-10(5): 498–501, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010273.

Lin:1985:ARP

[LG85] Kwei-Jay Lin and J. D. Gannon. Atomic remote procedure call. IEEE Transactions on Software Engineering, SE-11(10):1126–1135, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701928.

Lamb:1983:SPV

[LH83a] D. A. Lamb and P. N. Hilfinger. Simulation of procedure variables using Ada tasks. IEEE Transactions on Software Engineering, SE-9(1):13–15, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703007.

Leveson:1983:ASS

[LH83b] N. G. Leveson and P. R. Harvey. Analyzing software safety. IEEE Transactions on Software Engineering, SE-9(5): 569–579, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703097.

Laih:1989:DSK

[LHL89] Chi-Sung Laih, Lein Harn, and Jau-Yien Lee. On the design of a single-key-lock mechanism based on Newton’s interpolating polynomial. IEEE Transactions on Software Engineering, 15(9):1135–1137, September 1989. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31371.

Lichtman:1986:GCC

[Lic86]  Z. L. Lichtman.  Genera-

tion and consistency checking of design and program structures. IEEE Transactions on Software Engineering, SE-12(1):172–181, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312930.

Lingzhong:1989:C

[Lin89] Liu Lingzhong. Comments on “A distributed scheme for detecting communication deadlocks” by N. Natarajan. IEEE Transactions on Software Engineering, 15(7):926–??, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=29492.

See [Nat86].

Lipow:1982:NFP

[Lip82] M. Lipow. Number of faults per line of code. IEEE Transactions on Software Engineering, SE-8(4):437–439, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702967.

See comments [Ste86].

Lipow:1986:CEN

[Lip86] M. Lipow. Comments on “Estimating the number of faults in code” and two corrections to published data. IEEE Transactions on Software Engineering, SE-12(4):584–585, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312907.

See [Gaf84].

Liskov:1982:LSD

[Lis82]  B. Liskov.  On linguistic

support for distributed programs. IEEE Transactions on Software Engineering, SE-8(3):203–210, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702937.

Littlewood:1980:TSR

[Lit80] B. Littlewood. Theories of software reliability: How good are they and how can they be improved? IEEE Transactions on Software Engineering, SE-6(5): 489–500, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702766.

Laski:1983:DFO

[LK83] J. W. Laski and B. Korel. A data flow oriented program testing strategy. IEEE Transactions on Software Engineering, SE-9(3):347–354, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703063.

Lin:1987:GML

[LK87] F. C. H. Lin and R. M. Keller.

The gradient model load balancing method. IEEE Transactions on Software Engineering, SE-13(1):32–38, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702130.

Lee:1988:RAA

[LL88] T. T. Lee and M. Y. Lai. A relational algebraic approach to protocol verification. IEEE Transactions on Software Engineering, 14(2):184–193, February 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4637.

Lazzerini:1989:AME

[LL89a] B. Lazzerini and L. Lopriore. Abstraction mechanisms for event control in program debugging. IEEE Transactions on Software Engineering, 15(7):890–901, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=29488.

Lin:1989:CAS

[LL89b] C. Y. Lin and R. R. Levary. Computer-aided software development process design. IEEE Transactions on Software Engineering, 15(9):1025–1037, September 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        31362.

Luk:1989:AES

[LL89c] W. S. Luk and F. Ling. An analytic/empirical study of distributed sorting on a local area network. IEEE Transactions on Software Engineering, 15(5):575–586, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24707.

Lee:1981:SCC

[LLS81] P. A. Lee, J. L. Lloyd, and S. K. Shrivastava. Some critical comments on the paper “An Optimal Approach to Fault Tolerant Software Systems Design” by Gannon and Shapiro. IEEE Transactions on Software Engineering, SE-7(6): 608–610, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/

stamp.jsp?arnumber=1702893.

See [GS78].

Lockman:1982:UTR

[LM82] A. Lockman and N. Minsky. Unidirectional transport of rights and take–grant control. IEEE Transactions on Software Engineering, SE-8(6): 597–604, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702993.

Litvintchouk:1984:DAS

[LM84] S. D. Litvintchouk and A. S.

Matsumoto. Design of Ada systems yielding reusable components: An approach using structured algebraic specification. IEEE Transactions on Software Engineering, SE-10(5): 544–551, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010278.

Leventhal:1987:CTU

[LM87] L. M. Leventhal and B. T. Mynatt. Components of typical undergraduate software engineering courses: Results from a survey. IEEE Transactions on Software Engineering, SE-13(11):1193–1198, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702167.

Littlewood:1989:CMC

[LM89] B. Littlewood and D. R. Miller. Conceptual modeling of coincident failures in multiversion software. IEEE Transactions on Software Engineering, 15(12):1596–1614, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=58771.

Larson:1989:TAE

[LNE89] J. A. Larson, S. B. Navathe, and R. Elmasri. A theory of attributed equivalence in databases with application to schema integration. IEEE Transactions on Software Engineering, 15(4):449–463, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=16605.

Lantz:1985:ESD

[LNT85] K. A. Lantz, W. I. Nowicki, and A. M. Theimer. An empirical study of distributed application performance. IEEE Transactions on Software Engineering, SE-11(10):1162–1174, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701932.

Lomet:1980:SPD

[Lom80] D. B. Lomet. Subsystems of processes with deadlock avoidance. IEEE Transactions on Software Engineering, SE-6(3):297–304, May/ June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702732.

Lomet:1985:MPS

[Lom85] D. B. Lomet. Making pointers safe in system programming languages. IEEE Transactions on Software Engineering, SE-11(1):87–96, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701901.

Labetoulle:1980:IMN

[LP80] J. Labetoulle and G. Pujolle.

Isolation method in a network of queues. IEEE Transactions on Software Engineering, SE-6(4):373–381, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702745.

Lavi:1987:ICS

[LPBD87] J. Z. Lavi, M. I. B. Porat, and A. Ben-David. IAI corporate software engineering training and education program. IEEE Transactions on Software Engineering, SE-13(11):1207–1216, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702169.

Linton:1989:MPP

[LQ89] M. A. Linton and R. W. Quong. A macroscopic profile of program compilation and linking. IEEE Transactions on Software Engineering, 15(4):427–436, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=16603.

Leung:1980:AFS

[LR80] W. H. Leung and C. V. Ramamoorthy. An approach to formal specification of control modules. IEEE Transactions on Software Engineering, SE-6(5): 485–489, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702765.

Leung:1987:CPS

[LR87] H. K. N. Leung and H. K. Reghbati. Comments on program slicing. IEEE Transactions on Software Engineering, SE-13(12):1370–1371, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702186.

Logrippo:1983:FSP

[LS83] L. Logrippo and D. R. Skuce.

File structures, program structures, and attributed grammars. IEEE Transactions on Software Engineering, SE-9(3):260–266, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703053.

Lam:1984:PVP

[LS84] S. S. Lam and A. U. Shankar.

Protocol verification via projections. IEEE Transactions on Software Engineering, SE-10(4):325–342, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010246.

Lundstrom:1985:FAD

[LS85] S. F. Lundstrom and E. E.

Swartzlander. Foreword advances in distributed computing systems. IEEE Transactions on Software Engineering, SE-11(10):1092–1096, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701924.

Leveson:1987:SAU

[LS87] N. G. Leveson and J. L. Stolzy. Safety analysis using Petri nets. IEEE Transactions on Software Engineering, SE-13(3):386–397, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702225.

Lam:1988:PDD

[LSK88] H. Lam, S. Y. W. Su, and N. R. Koganti. A physical database design evaluation system for CODASYL databases. IEEE Transactions on Software Engineering, 14(7):1010–1022, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=42741.

Lakshmanan:1989:EDP

[LTC89] K. B. Lakshmanan, K. Thulasiraman, and M. A. Comeau. An efficient distributed protocol for finding shortest paths in networks with negative weights. IEEE Transactions on Software Engineering, 15(5):639–644, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24713.

Luccio:1985:VMR

[Luc85] F. Luccio. Variations on a method for representing data items of unlimited length. IEEE Transactions on Software Engineering, SE-11(4):439–441, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702026.

Ludewig:1983:ESP

[Lud83] J. Ludewig. ESPRESO — a system for process control software specification. IEEE Transactions on Software Engineering, SE-9(4):427–436, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703077.

Lind:1989:EIS

[LV89] R. K. Lind and K. Vairavan. An experimental investigation of software metrics and their relationship to software development effort. IEEE Transactions on Software Engineering, 15(5):649–653, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24715.

Lakhanai:1984:C

[LW84] G. D. Lakhanai and J. S.

Wang. Comments on “Optimization Algorithms for Distributed Queries”. IEEE Transactions on Software Engineering, SE-10(4):464–465, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010261.

See [AHY83].

Lai:1989:DJO

[LWL89] M. Y. Lai, W. K. Wilkinson, and V. Lanin. On distributing JASMIN’s optimistic multiversioning page manager. IEEE Transactions on Software Engineering, 15(6):696–704, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24723.

Ledgard:1980:EHE

[LWSS80] H. F. Ledgard, J. A. Whiteside, W. Seymour, and A. Singer. An experiment on human engineering of interactive software. IEEE Transactions on Software Engineering, SE-6(6): 602–604, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702787.

Leinbaugh:1986:GRT

[LY86] D. W. Leinbaugh and M. R. Yamini. Guaranteed response times in a distributed hard-real-time environment. IEEE Transactions on Software Engineering, SE-12(12):1139–1144, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313009.

Mourad:1987:RIM

[MA87] S. Mourad and D. Andrews.

On the reliability of the IBM MVS/XA operating system. IEEE Transactions on Software Engineering, SE-13 (10):1135–1139, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702153.

MacEwen:1982:C

[Mac82]  G. H. MacEwen.  Correc-

tion to “Information Flow Certification Using an Intermediate Code Program Representation”. IEEE Transactions on Software Engineering, SE-8(4):446, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702969.

See [MM81].

MacEwen:1983:DSS

[Mac83] G. H. MacEwen. The design for a secure system based on program analysis. IEEE Transactions on Software Engineering, SE-9(3):289–299, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703056.

Madhavji:1988:FBP

[Mad88] N. H. Madhavji. Fragtypes: a basis for programming environments. IEEE Transactions on Software Engineering, 14(1):85–97, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4625.

Magott:1987:NNC

[Mag87] J. Magott. New NP-complete problems in performance evaluation of concurrent systems using Petri nets. IEEE Transactions on Software Engineering, SE-13(5):578–581, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702257.

Manber:1984:CMB

[Man84] U. Manber. Concurrent maintenance of binary search trees. IEEE Transactions on Software Engineering, SE-10(6): 777–784, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010306.

Mancini:1986:MRM

[Man86]  L. Mancini.  Modular re-

dundancy in a message passing system. IEEE Transactions on Software Engineering, SE-12(1):79–86, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312922.

Marselos:1985:FTC

[Mar85] N. L. Marselos. Foreword technical conferences: A window to the future. IEEE Transactions on Software Engineering, SE-11(9):830–831, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702101.

Matsumoto:1984:SEP

[Mat84] Y. Matsumoto. Some experiences in promoting reusable software: Presentation in higher abstract levels. IEEE Transactions on Software Engineering, SE-10(5):502–513, September/October 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010274.

Mathis:1986:LP

[Mat86]  R. F. Mathis.  The last

10 percent. IEEE Transactions on Software Engineering, SE-12(6):705–712, June 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312968.

Mathur:1987:MBS

[Mat87] R. N. Mathur. Methodology for business system development. IEEE Transactions on Software Engineering, SE-13(5):593–601, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702259.

May:1989:PAP

[May89] C. May. The parallel assignment problem redefined. IEEE Transactions on Software Engineering, 15(6):821–824, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24735.

Menand:1983:MMA

[MB83] J. R. Menand and M. Becker. Modeling a multiprocessor architecture. IEEE Transactions on Software Engineering, SE-9(2):201–210, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703038.

Marsan:1989:EEP

[MBB^{+}89] M. Ajmone Marsan, G. Balbo, A. Bobbio, G. Chiola, G. Conte, and A. Cumani. The effect of execution policies on the semantics and analysis of stochastic Petri nets. IEEE Transactions on Software Engineering, 15(7):832–846, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=29483.

Mills:1989:MPF

[MBGH89] H. D. Mills, V. R. Basili, J. D. Gannon, and R. G. Hamlet. Mathematical principles for a first course in software engineering. IEEE Transactions on Software Engineering, 15(5):550–559, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24704.

Misra:1981:PNP

[MC81] J. Misra and K. M. Chandy.

Proofs of networks of processes. IEEE Transactions on Software Engineering, SE-7(4): 417–426, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702862.

See corrigendum [Oss82].

McKendry:1984:ILS

[MC84] M. S. McKendry and R. H.

Campbell. Implementing language support in high-level languages. IEEE Transactions on Software Engineering, SE-10(3):227–236, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010231.

Millen:1987:IPS

[MCF87] J. K. Millen, S. C. Clark, and S. B. Freedman. The interrogator: Protocol secuity analysis. IEEE Transactions on Software Engineering, SE-13(2):274–288, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702206.

McGill:1984:SES

[McG84] J. P. McGill. The software engineering shortage: A third choice. IEEE Transactions on Software Engineering, SE-10(1):42–49, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010197.

Milutinovic:1988:SST

[MCH88] V. M. Milutinovic, J. J. Crnkovic, and C. E. Houstis. A simulation study of two distributed task allocation procedures. IEEE Transactions on Software Engineering, 14(1):54–61, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4622.

Mckendry:1985:OAV

[Mck85] M. S. Mckendry. Ordering actions for visibility. IEEE Transactions on Software Engineering, SE-11(6):509–519, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702048.

Mikkilineni:1988:PNB

[MCS88] K. P. Mikkilineni, Y. C. Chow, and S. Y. W. Su. Petrinet-based modeling and evaluation of pipelined processing of concurrent database queries. IEEE Transactions on Software Engineering, 14(11): 1656–1667, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9053.

Martin:1986:OSG

[MD86] E. W. Martin and R. A. De Millo. Operational survivability in gracefully degrading distributed processing systems. IEEE Transactions on Software Engineering, SE-12(6):693–704, June 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312967.

Mueller:1988:HRC

[MDSW88] R. A. Mueller, M. R. Duda, P. H. Sweany, and J. S.

Walicki. Horizon: a retargetable compiler for horizontal microarchitectures. IEEE Transactions on Software Engineering, 14(5):575–583, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6135.

Meandzija:1986:FMC

[Mea86] B. Meandzija. A formal method for composing a network command language. IEEE Transactions on Software Engineering, SE-12(8):860–865, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312988.

Meandzija:1988:AUM

[Mea88] B. Meandzija. Archetype: a unified method for the design and implementation of protocol architectures. IEEE Transactions on Software Engineering, 14(6):822–837, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6161.

Mendelson:1982:AEH

[Men82] H. Mendelson. Analysis of extendible hashing. IEEE Transactions on Software Engineering, SE-8(6):611–619, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702995.

Mohan:1985:LCN

[MFKS85] C. Mohan, D. Fussell, Z. M.

Kedem, and A. Silberschatz. Lock conversion in non-two-phase locking protocols. IEEE Transactions on Software Engineering, SE-11(1):15–22, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701894.

McMullin:1983:CTF

[MG83] P. R. McMullin and J. D.

Gannon. Combining testing with formal specifications: A case study. IEEE Transactions on Software Engineering, SE-9(3):328–335, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703061.

Moitra:1986:DPA

[MI86] A. Moitra and S. S. Iyengar. Derivation of a parallel algorithm for balancing binary trees. IEEE Transactions on Software Engineering, SE-12(3):442–449, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312885.

Minoura:1989:DTA

[MI89] T. Minoura and S. S. Iyengar. Data and time abstraction techniques for analyzing multilevel concurrent systems. IEEE Transactions on Software Engineering, 15(1):47–59, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=21725.

Moitra:1988:MDS

[MIBY88] A. Moitra, S. S. Iyengar, F. B. Bastani, and I. L. Yen. Multilevel data structures: models and performance. IEEE Transactions on Software Engineering, 14(6):858–867, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6164.

Mili:1985:TTF

[Mil85] A. Mili. Towards a theory of forward error recovery. IEEE Transactions on Software Engineering, SE-11(8):735–748, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702083.

Miller:1986:EOS

[Mil86] D. R. Miller. Exponential order statistic models of software reliability growth. IEEE Transactions on Software Engineering, SE-12(1):12–24, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312915.

Milutinovic:1987:SSV

[Mil87] V. Milutinovic. A simulation study of the verticalmigration microprocessor architecture. IEEE Transactions on Software Engineering, SE-13(12):1265–1277, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702178.

Muro:1985:EFR

[MIMH85] S. Muro, T. Ibaraki, H. Miyajima, and T. Hasegawa. Evaluation of the file redundancy in distributed database systems. IEEE Transactions on Software Engineering, SE-11 (2):199–205, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701988.

Masuyama:1987:SSS

[MINH87] S. Masuyama, T. Ibaraki, S. Nishio, and T. Hasegawa. Shortest semijoin schedule for a local area distributed database system. IEEE Transactions on Software Engineering, SE-13(5):602–606, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702260.

Mohan:1988:OOK

[MK88] L. Mohan and R. L. Kashyap. An object-oriented knowledge representation for spatial information. IEEE Transactions on Software Engineering, 14(5):675–681, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6146.

Manolopoulos:1989:ECP

[MK89] Y. Manolopoulos and J. G. Kollias. Expressions for completely and partly unsuccessful batched search of sequential and tree-structured files. IEEE Transactions on Software Engineering, 15(6):794–799, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24732.

Mamrak:1989:CSS

[MKNS89] S. A. Mamrak, M. S. Kaelbling, C. K. Nicholas, and M. Share. Chameleon: a system for solving the data-translation problem. IEEE Transactions on Software Engineering, 15(9):1090–1108, September 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=31367.

Magee:1989:CDS

[MKS89] J. Magee, J. Kramer, and M. Sloman. Constructing distributed systems in conic. IEEE Transactions on Software Engineering, 15(6):663–675, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24720.

Ma:1981:DMC

[ML81] P. Y. R. Ma and T. G.

Lewis. On the design of a microcode compiler for a machine-independent high-level language. IEEE Transactions on Software Engineering, SE-7(3):261–274, May/ June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702839.

Mills:1986:DSP

[ML86] H. D. Mills and R. C. Linger.

Data structured programming: Program design without arrays and pointers. IEEE Transactions on Software Engineering, SE-12(2):192–197, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312935.

Merrett:1989:ISP

[ML89] T. H. Merrett and N. Laliberte. Including scalars in a programming language based on the relational algebra. IEEE Transactions on Software Engineering, 15(11):1437–1443, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=41335.

Mennie:1981:IFC

[MM81] A. L. Mennie and G. H. Macewen. Information flow certification using an intermediate code program representation. IEEE Transactions on Software Engineering, SE-7(6): 594–607, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702892.

See correction [Mac82].

Maekawa:1982:PAA

[MM82] M. Maekawa and Y. Morimoto. Performance adjustment of an APL interpreter. IEEE Transactions on Software Engineering, SE-8(4):331–343, July/ August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702956.

Meadows:1987:MSA

[MM87] C. Meadows and D. Mutchler. Matching secrets in the absence of a continuously available trusted authority. IEEE Transactions on Software Engineering, SE-13(2):289–292, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702207.

Matsumura:1987:ASM

[MMA87] K. Matsumura, H. Mizutani, and M. Arai. An application of structural modeling to software requirements analysis and design. IEEE Transactions on Software Engineering, SE-13(4):461–471, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702237.

Medina-Mora:1981:IPE

[MMF81] R. Medina-Mora and P. H.

Feiler. An incremental programming environment. IEEE Transactions on Software Engineering, SE-7(5):472–482, September/October 1981. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702873.

Matthews:1988:MCT

[MMS88] R. S. Matthews, K. H. Muralidhar, and S. Sparks. MAP 2.1 conformance testing tools. IEEE Transactions on Software Engineering, 14(3):363–374, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4656.

Malhotra:1989:ERP

[MMT^{+}89] A. Malhotra, H. M. Markowitz, Y. Tsalalikhin, D. P. Pazel, and L. M. Burns. An entityrelationship programming language. IEEE Transactions on Software Engineering, 15(9): 1120–1130, September 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=31369.

Mittermeir:1987:SBF

[MO87] R. T. Mittermeir and M. Oppitz. Software bases for the flexible composition of application systems. IEEE Transactions on Software Engineering, SE-13(4):440–460, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702236.

McNamee:1990:CCR

[MO90] C. M. McNamee and R. A. Olsson. Comments on “Critical races in Ada programs”. IEEE Transactions on Software Engineering, 16(12):1439–??, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/

stamp.jsp?arnumber=62452.

See [KSB89].

Moher:1988:PPV

[Moh88] T. G. Moher. PROVIDE: a process visualization and debugging environment. IEEE Transactions on Software Engineering, 14(6):849–857, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6163.

Moitra:1982:DIA

[Moi82] A. Moitra. Direct implementation of algebraic specification of abstract data types. IEEE Transactions on Software Engineering, SE-8(1):12– 20, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702901.

Molloy:1985:DTS

[Mol85] M. K. Molloy. Discrete time stochastic Petri nets. IEEE Transactions on Software Engineering, SE-11(4):417–423, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702023.

Mostow:1985:FWA

[Mos85] J. Mostow. Foreword what is AI? And what does it have to do with software engineering? IEEE Transactions on Software Engineering, SE-11(11):1253–1256, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701944.

Motro:1987:SVI

[Mot87] A. Motro. Superviews: Virtual integration of multiple databases. IEEE Transactions on Software Engineering, SE-13(7):785–798, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702290.

Marie:1983:SSP

[MP83] R. A. Marie and J. M. Pellaumail. Steady-state probabilities for a queue with a general service distribution and state-dependent arrivals. IEEE Transactions on Software Engineering, SE-9 (1):109–113, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703018.

Martella:1985:AMD

[MPS85] G. Martella, B. Pernici, and F. A. Schreiber. An availability model for distributed transaction systems. IEEE Transactions on Software Engineering, SE-11(5):483–491, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702039.

Morgan:1987:ISS

[MR87] E. T. Morgan and R. R.

Razouk. Interactive statespace analysis of concurrent systems. IEEE Transactions on Software Engineering, SE-13(10):1080–1091, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702148.

Maier:1985:RRU

[MRS85] D. Maier, D. Rozenshtein, and J. Stein. Representing roles in universal scheme interfaces. IEEE Transactions on Software Engineering, SE-11(7):644–652, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702068.

Morgan:1984:SUF

[MS84] C. Morgan and B. Sufrin.

Specification of the UNIX filing system. IEEE Transactions on Software Engineering, SE-10(2):128–142, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010215.

MacQueen:1985:CPS

[MS85] D. B. MacQueen and D. T. Sannella. Completeness of proof systems for equational specifications. IEEE Transactions on Software Engineering, SE-11(5):454–461, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702035.

Moore:1987:CSK

[MS87] J. H. Moore and G. J. Simmons. Cycle structure of the DES for keys having palindromic (or antipalindromic) sequences of round keys. IEEE Transactions on Software Engineering, SE-13(2):262–273, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702205.

Mikkilineni:1988:ERJ

[MS88] Krishna P. Mikkilineni and Stanley Y. W. Su. An evaluation of relational join algorithms in a pipelined query processing environment. IEEE Transactions on Software Engineering, 14(6):838–848, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6162.

Maggiolo-Schettini:1988:WST

[MSNT88] A. Maggiolo-Schettini, M. Napoli, and G. Tortora. Web structures: a tool for representing and manipulating programs. IEEE Transactions on Software Engineering, 14(11): 1621–1639, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9050.

Murata:1989:DAS

[MSS89] T. Murata, B. Shenker, and S. M. Shatz. Detection of Ada static deadlocks using Petri net invariants. IEEE Transactions on Software Engineering, 15(3):314–326, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21759.

Munoz:1988:ASP

[Mun88]  C. U. Munoz.  An ap-

proach to software product testing. IEEE Transactions on Software Engineering, 14(11): 1589–1596, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9047.

Murata:1980:SDF

[Mur80]  T. Murata.  Synthesis of

decision-free concurrent systems for prescribed resources and performance. IEEE Transactions on Software Engineering, SE-6(6):525–530, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702778.

Musser:1980:ADT

[Mus80] D. R. Musser. Abstract data type specification in the affirm system. IEEE Transactions on Software Engineering, SE-6(1):24–32, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702691.

Musa:1983:GEC

[Mus83]  J. D. Musa.  Guest edi-

torial: COMPSAC ’81 special section. IEEE Transactions on Software Engineering, SE-9(2):118–119, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703027.

Minoura:1982:RET

[MW82] T. Minoura and G. Wiederhold. Resilient extended true-copy token scheme for a distributed database system. IEEE Transactions on Software Engineering, SE-8(3):173–189, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702934.

Moldovan:1988:HKB

[MW88] D. I. Moldovan and C. I.

Wu. A hierarchical knowledge based system for airplane classification. IEEE Transactions on Software Engineering, 14(12):1829–1834, December 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9066.

Mao:1980:CPL

[MY80a] T. W. Mao and R. T. Yeh.

Communication port: A language concept for concurrent programming. IEEE Transactions on Software Engineering, SE-6(2):194–204, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702716.

Mekly:1980:SDR

[MY80b] L. J. Mekly and S. S. Yau.

Software design representation using abstract process networks. IEEE Transactions on Software Engineering, SE-6(5): 420–435, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702758.

Murata:1988:PTN

[MZ88] T. Murata and D. Zhang. A predicate-transition net model for parallel interpretation of logic programs. IEEE Transactions on Software Engineering, 14(4):481–497, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4671.

Natarajan:1985:CSP

[Nat85] N. Natarajan. Communication and synchronization primitives for distributed programs. IEEE Transactions on Software Engineering, SE-11(4):396–416, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702022.

Natarajan:1986:DSD

[Nat86] N. Natarajan. A distributed scheme for detecting communication deadlocks. IEEE Transactions on Software Engineering, SE-12(4):531–537, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312900.

See comments [Lin89].

Neighbors:1984:DAC

[Nei84] J. M. Neighbors. The draco approach to constructing software from reusable components.

IEEE Transactions on Software Engineering, SE-10(5): 564–574, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010280.

Nessett:1987:FAD

[Nes87] D. M. Nessett. Factors affecting distributed system security. IEEE Transactions on Software Engineering, SE-13(2):233–248, February 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702203.

Neumann:1986:HDC

[Neu86] P. G. Neumann. On hierarchical design of computer systems for critical applications. IEEE Transactions on Software Engineering, SE-12 (9):905–920, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313046.

Ng:1981:FAE

[Ng81] P. A. Ng. Further analysis of the entity-relationship approach to database design. IEEE Transactions on Software Engineering, SE-7(1):85– 99, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702805.

Ni:1981:PMS

[NH81] L. M. Ni and Kai Hwang. Performance modeling of sharedresource array processors. IEEE Transactions on Software Engineering, SE-7(4):386–394, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702859.

Ni:1985:OLB

[NH85] L. M. Ni and Kai Hwang. Optimal load balancing in a multiple processor system with many job classes. IEEE Transactions on Software Engineering, SE-11(5):491–496, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702040.

See correction [NH86].

Ni:1986:COL

[NH86] L. M. Ni and K. Hwang. Correction to “Optimal load balancing in a multiple processor system with many job classes”. IEEE Transactions on Software Engineering, SE-12(3):500, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312891.

See [NH85].

Nehmer:1987:KCI

[NHM^{+}87] J. Nehmer, D. Haban, F. Mattern, D. Wybranietz, and H. D. Rombach. Key concepts of the INCAS Multicomputer Project. IEEE Transactions on Software Engineering, SE-13(8):913–923, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702310.

Nelson:1983:CPN

[NHS83] R. A. Nelson, L. M. Haibt, and P. B. Sheridan. Casting Petri nets into programs. IEEE Transactions on Software Engineering, SE-9(5): 590–602, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703099.

Nicol:1989:OPR

[Nic89] D. M. Nicol. Optimal partitioning of random programs across two processors. IEEE Transactions on Software Engineering, 15(2):134–141, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21740.

Nishigaki:1983:EKC

[Nis83] T. Nishigaki. Experiments on the knee criterion in a multiprogrammed computer system. IEEE Transactions on Software Engineering, SE-9 (1):79–86, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703014.

Nicola:1987:QAF

[NKT87] V. F. Nicola, V. G. Kulkarni, and K. S. Trivedi. Queueing analysis of fault-tolerant computer systems. IEEE Transactions on Software Engineering, SE-13(3):363–375, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702223.

Negrini:1983:SPD

[NS83] R. M. Negrini and M. Sami. Some properties derived from structural analysis of program graph models. IEEE Transactions on Software Engineering, SE-9(2):172–178, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703034.

Narayanaswamy:1987:MCE

[NS87] K. Narayanaswamy and W. Scacchi. Maintaining configurations of evolving software systems. IEEE Transactions on Software Engineering, SE-13(3):324–334, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702218.

Nosek:1988:UVI

[NS88] J. T. Nosek and R. B. Schwartz.

User validation of information system requirements: some empirical results. IEEE Transactions on Software Engineering, 14(9):1372–1375, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6180.

Neches:1985:EME

[NSM85] R. Neches, W. R. Swartout, and J. D. Moore. Enhanced maintenance and explanation of expert systems through explicit models of their development. IEEE Transactions on Software Engineering, SE-11(11):1337–1351, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701950.

Ntafos:1984:RET

[Nta84] S. C. Ntafos. On required element testing. IEEE Transactions on Software Engineering, SE-10(6):795–803, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010308.

Ntafos:1988:CSS

[Nta88] S. C. Ntafos. A comparison of some structural testing strategies. IEEE Transactions on Software Engineering, 14(6):868–874, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6165.

Nelson:1988:PAP

[NTT88] R. Nelson, D. Towsley, and A. N. Tantawi. Performance analysis of parallel processing systems. IEEE Transactions on Software Engineering, 14(4):532–540, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4676.

Ni:1989:DTP

[NW89] L. M. Ni and C. F. E. Wu. Design tradeoffs for process scheduling in shared memory multiprocessor systems. IEEE Transactions on Software Engineering, 15(3):327–334, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=21760.

Ni:1985:DDA

[NXG85] L. M. Ni, Chong-Wei Xu, and T. B. Gendreau. A distributed drafting algorithm for load balancing. IEEE Transactions on Software Engineering, SE-11(10):1153–1161, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701931.

Oberndorf:1988:CAP

[Obe88] P. A. Oberndorf. The Common Ada Programming Support Environment (APSE) Interface Set (CAIS). IEEE Transactions on Software Engineering, 14(6):742–748, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6154.

Ozsoyoglu:1982:ESS

[OC82] G. Ozsoyoglu and F. Y. Chin.

Enhancing the security of statistical databases with a question-answering system and a kernel design. IEEE Transactions on Software Engineering, SE-8(3):223–234, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702939.

Oldehoeft:1984:DRM

[OJ84] A. E. Oldehoeft and S. F.

Jennings. Dataflow resource managers and their synthesis from open path expressions. IEEE Transactions on Software Engineering, SE-10 (3):244–257, May/June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010233.

Okumoto:1985:SMS

[Oku85] K. Okumoto. A statistical method for software quality control. IEEE Transactions on Software Engineering, SE-11(12):1424–1430, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701964.

Oldehoeft:1983:PGE

[Old83] R. R. Oldehoeft. Program graphs and execution behavior. IEEE Transactions on Software Engineering, SE-9 (1):103–108, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703017.

Orenstein:1988:PSD

[OM88] J. A. Orenstein and F. A. Manola. PROBE spatial data modeling and query processing in an image database application. IEEE Transactions on Software Engineering, 14(5):611–629, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6139.

Omiecinski:1989:HJP

[Omi89] Edward R. Omiecinski. Heuristics for join processing using nonclustered indexes. IEEE Transactions on Software Engineering, 15(1):18–25, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21722.

Ozsoyoglu:1985:SDQ

[OO85] G. Ozsoyoglu and Z. M. Ozsoyoglu. Statistical database query languages. IEEE Transactions on Software Engineering, SE-11(10):1071–1081, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701922.

Onvural:1989:ATA

[OP89] R. O. Onvural and H. G. Perros. Approximate throughput analysis of cyclic queueing networks with finite buffers. IEEE Transactions on Software Engineering, 15(6):800–808, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24733.

Orman:1988:FDD

[Orm88] L. Orman. Functional development of database applications. IEEE Transactions on Software Engineering, 14(9):1280–1292, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6172.

Ossefort:1982:C

[Oss82]  M. Ossefort.  Corrigendum

to “Proofs of Networks of Processes”. IEEE Transactions on Software Engineering, SE-8(2):160, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702923.

See [MC81].

Ossher:1989:CSS

[Oss89] H. Ossher. A case study in structure specification: a grid description of scribe. IEEE Transactions on Software Engineering, 15(11): 1397–1416, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41332.

Osterweil:1983:TES

[Ost83] L. J. Osterweil. Toolpack — an experimental software development environment research project. IEEE Transactions on Software Engineering, SE-9(6): 673–685, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703114.

Ottenstein:1989:SFR

[Ott89] K. J. Ottenstein. A simplified framework for reduction in strength. IEEE Transactions on Software Engineering, 15(1):86–92, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21730.

Ozsoyoglu:1989:RCS

[OW89] G. Ozsoyoglu and H. Wang.

A relational calculus with set operators, its safety, and equivalent graphical languages. IEEE Transactions on Software Engineering, 15(9):1038– 1052, September 1989. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
31363.

Ozsu:1985:MAD

[Ozs85] M. T. Ozsu. Modeling and analysis of distributed database concurrency control algorithms using an extended Petri net formalism. IEEE Transactions on Software Engineering, SE-11(10):1225–1240, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701938.

Perros:1986:AAO

[PA86] H. G. Perros and T. Altiok. Approximate analysis of open networks of queues with blocking: Tandem configurations. IEEE Transactions on Software Engineering, SE-12(3):450–461, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312886.

Pagan:1983:DNA

[Pag83] F. G. Pagan. A diagrammatic notation for abstract syntax and abstract structured objects. IEEE Transactions on Software Engineering, SE-9(3):280–289, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703055.

Pagan:1987:PSC

[Pag87] F. G. Pagan. Program structure charts for applicative languages. IEEE Transactions on Software Engineering, SE-13(4):490–493, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702240.

Papakonstantinou:1981:IAG

[Pap81] G. K. Papakonstantinou. An interpreter of attribute grammars and its application to waveform analysis. IEEE Transactions on Software Engineering, SE-7(3):279–283, May/ June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702841.

Parr:1980:ARC

[Par80]  F. N. Parr.  An alterna-

tive to the Rayleigh curve model for software development effort. IEEE Transactions on Software Engineering, SE-6(3):291–296, May/ June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702731.

Pashtan:1985:OSM

[Pas85] A. Pashtan. Operating system models in a Concurrent Pascal environment: Complexity and performance considerations. IEEE Transactions on Software Engineering, SE-11(1):136–141, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701906.

Pun:1987:PST

[PB87] K. H. Pun and G. G. Belford.

Performance study of two phase locking in single-site database systems. IEEE Transactions on Software Engineering, SE-13(12):1311–1328, December 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702181.

Peachey:1985:SEO

[PBC85] J. B. Peachey, R. B. Bunt, and C. J. Colbourn. Some empirical observations on program behavior with applications to program restructuring. IEEE Transactions on Software Engineering, SE-11 (2):188–193, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701986.

Parnas:1986:CRD

[PC86a] D. L. Parnas and P. C. Clements. Correction to “A rational design process: How and why to fake it”. IEEE Transactions on Software Engineering, SE-12(8):874, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312991.

See [PC86b].

Parnas:1986:RDP

[PC86b] D. L. Parnas and P. C. Clements. A rational design process: How and why to fake it. IEEE Transactions on Software Engineering, SE-12(2):251–257, February 1986. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312940.

See correction [PC86a].

Perrott:1985:CAV

[PCMP85] R. H. Perrott, D. Crookes, P. Milligan, and W. R. M. Purdy. A compiler for an array and vector processing language. IEEE Transactions on Software Engineering, SE-11(5):471–478, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702037.

Parnas:1985:MSC

[PCW85] D. L. Parnas, P. C. Clements, and D. M. Weiss. The modular structure of complex systems. IEEE Transactions on Software Engineering, SE-11(3):259–266, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702002.

Peercy:1981:SME

[Pee81] D. E. Peercy. A software maintainability evaluation methodology. IEEE Transactions on Software Engineering, SE-7(4): 343–351, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702852.

Perros:1981:SEO

[Per81] H. G. Perros. A symmetrical exponential open queue network with blocking and feedback. IEEE Transactions on Software Engineering, SE-7(4):395–402, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702860.

Perrizo:1984:MPD

[Per84] W. Perrizo. A method for processing distributed database queries. IEEE Transactions on Software Engineering, SE-10(4):466–471, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010262.

Perry:1985:RBA

[Per85]  K. J. Perry.  Randomized

Byzantine agreement. IEEE Transactions on Software Engineering, SE-11(6):539–546, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702051.

Phillips:1984:SDT

[Phi84] N. C. K. Phillips. Safe data type specifications. IEEE Transactions on Software Engineering, SE-10(3):285–289, May/ June 1984. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010237.

Patnaik:1986:LLF

[PI86] L. M. Patnaik and K. V. Iyer.

Load-leveling in fault-tolerant distributed computing systems. IEEE Transactions on Software Engineering, SE-12(4):554–560, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312903.

Plattner:1984:RTE

[Pla84] B. Plattner. Real-time execution monitoring. IEEE Transactions on Software Engineering, SE-10(6):756–764, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010304.

Prather:1987:PPS

[PM87] R. E. Prather and J. P. Myers. The path prefix software testing strategy. IEEE Transactions on Software Engineering, SE-13(7):761–766, July 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702287.

Peterka:1989:PPA

[PM89] G. Peterka and T. Murata.

Proof procedure and answer extraction in Petri net model of logic programs. IEEE Transactions on Software Engineering, 15(2):209–217, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21746.

Pu:1988:RRO

[PNP88] C. Pu, J. D. Noe, and A. Proudfoot. Regeneration of replicated objects: a technique and its Eden implementation. IEEE Transactions on Software Engineering, 14(7):936–945, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=42736.

Polster:1986:RST

[Pol86] F. J. Polster. Reuse of software through generation of partial systems. IEEE Transactions on Software Engineering, SE-12(3):402–416, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312882.

Prywes:1983:CNS

[PP83] N. S. Prywes and A. Pnueli. Compilation of nonprocedural specifications into computer programs. IEEE Transactions on Software Engineering, SE-9(3):267–279, May/ June 1983. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703054.

Parker:1983:DMI

[PPR^{+}83] D. S. Parker, G. J. Popek, G. Rudisin, A. Stoughton, B. J. Walker, E. Walton, J. M. Chow, D. Edwards, S. Kiser, and C. Kline. Detection of mutual inconsistency in distributed systems. IEEE Transactions on Software Engineering, SE-9(3):240–247, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703051.

See comments [Ram87].

Panzieri:1985:IUD

[PR85] F. Panzieri and B. Randell. Interfacing UNIX to data communications networks. IEEE Transactions on Software Engineering, SE-11(10): 1016–1032, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701916.

Probert:1982:OIS

[Pro82]  R. L. Probert.  Optimal

insertion of software probes in well-delimited programs. IEEE Transactions on Software Engineering, SE-8(1):34– 42, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702903.

Parent:1985:AGE

[PS85] C. Parent and S. Spaccapietra. An algebra for a general entity-relationship model. IEEE Transactions on Software Engineering, SE-11(7):634–643, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702067.

Purushothaman:1987:RAP

[PS87] S. Purushothaman and P. A. Subrahmanyam. Reasoning about probabilistic behavior in concurrent systems. IEEE Transactions on Software Engineering, SE-13(6):740–745, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702278.

Panzieri:1988:RRP

[PS88] F. Panzieri and S. K. Shrivastava. Rajdoot: a remote procedure call mechanism supporting orphan detection and killing. IEEE Transactions on Software Engineering, 14(1):30–37, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4620.

Pollock:1989:IVI

[PS89] L. L. Pollock and M. L. Soffa.

An incremental version of iterative data flow analysis. IEEE Transactions on Software Engineering, 15(12):1537–1549, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58766.

Perry:1986:DAP

[PT86] K. J. Perry and S. Toueg. Distributed agreement in the presence of processor and communication faults. IEEE Transactions on Software Engineering, SE-12(3):477–482, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312888.

Pramanik:1988:OJQ

[PV88] S. Pramanik and D. Vineyard. Optimizing join queries in distributed databases. IEEE Transactions on Software Engineering, 14(9):1319–1326, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6175.

Qadah:1988:JAS

[QI88] G. Z. Qadah and K. B. Irani.

The join algorithms on a shared-memory multiprocessor database machine. IEEE Transactions on Software Engineering, 14(11):1668–1683, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9054.

Rajlich:1987:RMA

[Raj87] V. Rajlich. Refinement methodology for Ada. IEEE Transactions on Software Engineering, SE-13(4):472–478, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702238.

Ramamritham:1985:SCR

[Ram85] K. Ramamritham. Synthesizing code for resource controllers. IEEE Transactions on Software Engineering, SE-11(8):774–783, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702086.

Ramarao:1987:C

[Ram87] K. V. S. Ramarao. Comments on “Detection of Mutual Inconsistency in Distributed Systems”. IEEE Transactions on Software Engineering, SE-13(6):759–760, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/

stamp.jsp?arnumber=1702282.

See [PPR^{+}83].

Ravn:1980:DM

[Rav80] A. P. Ravn. Device monitors. IEEE Transactions on Software Engineering, SE-6 (1):49–53, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702694.

Ramamoorthy:1982:SRS

[RB82a] C. V. Ramamoorthy and F. B.

Bastani. Software reliability — status and perspectives. IEEE Transactions on Software Engineering, SE-8(4): 354–371, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702958.

Rowe:1982:LNB

[RB82b] L. A. Rowe and K. P. Birman. A local network based on the UNIX operating system. IEEE Transactions on Software Engineering, SE-8(2): 137–146, March/April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702921.

Ramakrishnan:1983:PDP

[RB83] I. V. Ramakrishnan and J. C.

Browne. A paradigm for the design of parallel algorithms with applications. IEEE Transactions on Software Engineering, SE-9(4):411–415, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703075.

Ramsey:1989:EES

[RB89] C. L. Ramsey and V. R. Basili. An evaluation of expert systems for software engineering management. IEEE Transactions on Software Engineering, 15(6):747–759, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24728.

Richardson:1985:PAM

[RC85] D. J. Richardson and L. A. Clarke. Partition analysis: A method combining testing and verification. IEEE Transactions on Software Engineering, SE-11(12):1477–1490, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701971.

Ramamoorthy:1985:IAP

[RDU85] C. V. Ramamoorthy, S. T. Dong, and Y. Usuda. An implementation of an Automated Protocol Synthesizer (APS) and its application to the X.21 protocol. IEEE Transactions on Software Engineering, SE-11(9):886–908, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702107.

Ramakrishnan:1989:PAM

[RE89] K. K. Ramakrishnan and J. S.

Emer. Performance analysis of mass storage service alternatives for distributed systems. IEEE Transactions on Software Engineering, 15 (2):120–133, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21739.

Redwine:1983:EAS

[Red83] S. T. Redwine. An engineering approach to software test data design. IEEE Transactions on Software Engineering, SE-9(2):191–200, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703037.

Reisner:1981:FGH

[Rei81] P. Reisner. Formal grammar and human factors design of an interactive graphics system. IEEE Transactions on Software Engineering, SE-7(2):229–240, March/ April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702829.

Reiss:1985:PPD

[Rei85] S. P. Reiss. PECAN: Program development systems that support multiple views. IEEE Transactions on Software Engineering, SE-11(3):276–285, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702004.

Reiss:1987:ACP

[Rei87] S. P. Reiss. Automatic compiler production: The front end. IEEE Transactions on Software Engineering, SE-13(6):609–627, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702267.

Reuter:1980:FTO

[Reu80] A. Reuter. A fast transactionoriented logging scheme for undo recovery. IEEE Transactions on Software Engineering, SE-6(4):348–356, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702743.

Reuss:1982:MIS

[Reu82] J. L. Reuss. Macro implementation of a structured assembly language. IEEE Transactions on Software Engineering, SE-8(3):284–287, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702943.

Roussopoulos:1988:EPD

[RFS88] N. Roussopoulos, C. Faloutsos, and T. Sellis. An efficient pictorial database system for PSQL. IEEE Transactions on Software Engineering, 14(5):639–650, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6141.

Ramamoorthy:1986:PL

[RGP86] C. V. Ramamoorthy, V. Garg, and A. Prakash. Programming in the large. IEEE Transactions on Software Engineering, SE-12(7):769–783, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312978.

Ramamoorthy:1988:SRG

[RGP88] C. V. Ramamoorthy, V. Garg, and A. Prakash. Support for reusability in genesis. IEEE Transactions on Software Engineering, 14(8):1145–1154, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7625.

Ramamoorthy:1980:PEA

[RH80] C. V. Ramamoorthy and G. S. Ho. Performance evaluation of asynchronous concurrent systems using Petri nets. IEEE Transactions on Software Engineering, SE-6(5): 440–449, September/October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702760.

Raghavendra:1985:ROD

[RH85] C. S. Raghavendra and S. Hariri. Reliability optimization in the design of distributed systems. IEEE Transactions on Software Engineering, SE-11 (10):1184–1193, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701934.

Rutenstreich:1986:TDP

[RH86] S. Rutenstreich and W. E. Howden. Two-dimensional program design. IEEE Transactions on Software Engineering, SE-12(3):377–384, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312880.

Rea:1987:AAD

[RJ87] K. Rea and R. De B. Johnston. Automated analysis of discrete communication behavior. IEEE Transactions on Software Engineering, SE-13 (10):1115–1126, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702151.

Ramamritham:1983:SSP

[RK83] K. Ramamritham and R. M.

Keller. Specification of synchronizing processes. IEEE Transactions on Software Engineering, SE-9(6):722–733, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703117.

Reuter:1984:ADI

[RK84] A. Reuter and H. Kinzinger.

Automatic design of the internal schema for a CODASYL database system. IEEE Transactions on Software Engineering, SE-10(4):358–375, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010249.

Ramesh:1987:MDD

[RM87] S. Ramesh and S. L. Mehndiratta. A methodology for developing distributed programs. IEEE Transactions on Software Engineering, SE-13(8):967–976, August 1987. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702314.

Ramamurthy:1988:SSS

[RM88] B. Ramamurthy and A. Melton. A synthesis of software science measures and the cyclomatic number. IEEE Transactions on Software Engineering, 14(8):1116–1121, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7622.

Ramamoorthy:1981:AMD

[RMB^{+}81] C. V. Ramamoorthy, Y. R. Mok, F. B. Bastani, G. H. Chin, and K. Suzuki. Application of a methodology for the development and validation of reliable process control software. IEEE Transactions on Software Engineering, SE-7(6): 537–555, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702887.

Reed:1988:PDE

[RMM88] D. A. Reed, A. D. Malony, and B. D. McCredie. Parallel discrete event simulation using shared memory. IEEE Transactions on Software Engineering, 14(4):541–553, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4677.

Robers:1981:GC

[Rob81]  T. Robers.  Global com-

ments. IEEE Transactions on Software Engineering, SE-7(4):370, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702855.

Rombach:1987:CEI

[Rom87] H. D. Rombach. A controlled experiment on the impact of software structure on maintainability. IEEE Transactions on Software Engineering, SE-13(3):344–354, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702220.

Ross:1985:SRS

[Ros85a] S. M. Ross. Software reliability: The stopping rule problem. IEEE Transactions on Software Engineering, SE-11(12):1472–1476, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701970.

Ross:1985:SES

[Ros85b] S. M. Ross. Statistical estimation of software reliability.

IEEE Transactions on Software Engineering, SE-11(5):479–483, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702038.

Roussopoulos:1982:LAP

[Rou82] N. Roussopoulos. The logical access path schema of a database. IEEE Transactions on Software Engineering, SE-8(6):563–573, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702990.

Rowe:1985:AEO

[Row85]  N. C. Rowe.  Antisam-

pling for estimation: An overview. IEEE Transactions on Software Engineering, SE-11(10):1081–1091, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701923.

Rowe:1988:ABS

[Row88] N. C. Rowe. Absolute bounds on set intersection and union sizes from distribution information. IEEE Transactions on Software Engineering, 14(7):1033–1048, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=42743.

Ryder:1988:EOF

[RP88] B. G. Ryder and J. S. Pendergrast. Experiments in optimizing FP. IEEE Transactions on Software Engineering, 14(4):444–454, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4668.

Roe:1987:STC

[RR87] R. P. Roe and J. H. Rowland.

Some theory concerning certification of mathematical subroutines by black box testing. IEEE Transactions on Software Engineering, SE-13(6):677–682, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702273.

Ries:1982:NTD

[RS82] D. R. Ries and G. C. Smith.

Nested transactions in distributed systems. IEEE Transactions on Software Engineering, SE-8(3):167–172, May/ June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702933.

Rowe:1983:PLC

[RS83] L. A. Rowe and K. A. Shoens.

Programming language constructs for screen definition. IEEE Transactions on Software Engineering, SE-9(1):31– 39, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703010.

Ramanthan:1988:PCA

[RS88a] J. Ramanthan and S. Sarkar. Providing customized assistance for software lifecycle approaches. IEEE Transactions on Software Engineering, 14(6):749–757, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6155.

Rotem:1988:AMP

[RS88b] D. Rotem and A. Segev. Algorithms for multidimensional partitioning of static files. IEEE Transactions on Software Engineering, 14(11): 1700–1710, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9056.

Rullo:1988:APD

[RS88c] P. Rullo and D. Sacca. An automatic physical designer for network model databases. IEEE Transactions on Software Engineering, 14(9):1293– 1306, September 1988. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
6173.

Ramamritham:1986:PTR

[RSBV86] K. Ramamritham, D. Stemple, D. A. Briggs, and S. Vinter. Privilege transfer and revocation in a port-based system. IEEE Transactions on Software Engineering, SE-12(5):635–648, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312959.

Reingold:1981:TDT

[RT81] E. M. Reingold and J. S.

Tilford. Tidier drawings of trees. IEEE Transactions on Software Engineering, SE-7(2): 223–228, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702828.

Rubin:1983:SDP

[Rub83] L. F. Rubin. Syntax-directed pretty printing — a first step towards a syntax-directed editor. IEEE Transactions on Software Engineering, SE-9(2): 119–127, March/April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703028.

Russell:1980:SRS

[Rus80] D. L. Russell. State restoration in systems of communicating processes. IEEE Transactions on Software Engineering, SE-6(2):183–194, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702715.

Rus:1988:PLP

[Rus88] T. Rus. Parsing languages by pattern matching. IEEE Transactions on Software Engineering, 14(4):498–511, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4672.

Ruschitzka:1989:HDT

[Rus89] M. Ruschitzka. Heterogeneous data translations based on environment grammars. IEEE Transactions on Software Engineering, 15(10):1236–1251, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=559774.

Rapps:1985:SST

[RW85] S. Rapps and E. J. Weyuker. Selecting software test data using data flow information. IEEE Transactions on Software Engineering, SE-11(4):367–375, April 1985. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702019.

Riddle:1987:TSE

[RW87a] W. E. Riddle and L. G.

Williams. Technology selection: An educational approach. IEEE Transactions on Software Engineering, SE-13(11):1199–1206, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702168.

Robb:1987:PCC

[RW87b] H. W. Robb and H. R. Weistroffer. Performance criteria for constrained nonlinear programming codes. IEEE Transactions on Software Engineering, SE-13(4):479–489, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702239.

Roussopoulos:1985:SST

[RY85] N. Roussopoulos and R. T. Yeh.

SEES — a Software testing Environment Support System. IEEE Transactions on Software Engineering, SE-11(4):355–366, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702018.

Sidhu:1988:CSS

[SA88] D. P. Sidhu and J. Aristizabal. Constructing submodule specifications and network protocols. IEEE Transactions on Software Engineering, 14(11):1565–1577, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=9045.

Son:1989:DCG

[SA89] Sang Hyuk Son and A. K. Agrawala. Distributed checkpointing for globally consistent states of databases. IEEE Transactions on Software Engineering, 15(10):1157–1167, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=559763.

Samet:1982:COC

[Sam82] H. Samet. Code optimization considerations in list processing systems. IEEE Transactions on Software Engineering, SE-8(2):107–112, March/ April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702918.

Samadi:1989:TKB

[Sam89]  B. Samadi.  TUNEX: a

knowledge-based system for performance tuning of the UNIX operating system. IEEE Transactions on Software Engineering, 15(7):861–874, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=29486.

Sanden:1989:CED

[San89a] B. Sanden. The case for electric design of real-time software. IEEE Transactions on Software Engineering, 15(3):360–362, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21764.

See correction [San89b].

Sanden:1989:CCE

[San89b]  B. Sanden.  Correction to

‘The case for electric design of real-time software’. IEEE Transactions on Software Engineering, 15(7):926–??, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=29491.

See [San89a].

Sandhu:1989:RIE

[San89c] R. S. Sandhu. Recognizing immediacy in an N-tree hierarchy and its application to protection groups. IEEE Transactions on Software Engineering, 15(12):1518–1525, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=58764.

Sarraga:1984:SDF

[Sar84] R. F. Sarraga. Static data flow analysis of PL/I programs with the PROBE system. IEEE Transactions on Software Engineering, SE-10 (4):451–459, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010259.

Selby:1987:CSD

[SBB87] R. W. Selby, V. R. Basili, and F. T. Baker. Cleanroom software development: An empirical evaluation. IEEE Transactions on Software Engineering, SE-13(9):1027–1037, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702325.

Spector:1985:SDT

[SBD^{+}85] A. Z. Spector, J. Butcher, D. S. Daniels, D. J. Duchamp, J. L. Eppinger, C. E. Fineman, A. Heddaya, and P. M. Schwarz. Support for distributed transactions in the TABS Prototype. IEEE Transactions on Software Engineering, SE-11(6):520–530, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702049.

Serazzi:1984:AOS

[SC84] G. Serazzi and M. Calzarossa.

Adaptive optimization of a system’s load. IEEE Transactions on Software Engineering, SE-10(6):837–845, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010312.

Schultz:1987:AMA

[SC87] R. D. Schultz and A. F. Cardenas. An approach and mechanism for auditable and testable advanced transaction processing systems. IEEE Transactions on Software Engineering, SE-13(6):666–676, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702272.

Sidhu:1988:ELS

[SC88] D. P. Sidhu and C. S. Crall.

Executable logic specifications for protocol service interfaces. IEEE Transactions on Software Engineering, 14(1):98–112, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4626.

Scacchi:1984:MSE

[Sca84] W. Scacchi. Managing software engineering projects: A social analysis. IEEE Transactions on Software Engineering, SE-10(1):49–59, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010198.

Shen:1983:SSR

[SCD83] V. Y. Shen, S. D. Conte, and H. E. Dunsmore. Software science revisited: A critical analysis of the theory and its empirical support. IEEE Transactions on Software Engineering, SE-9(2):155–165, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703032.

Schwarz:1982:UAM

[Sch82] J. Schwarz. Using annotations to make recursion equations behave. IEEE Transactions on Software Engineering, SE-8(1):21–33, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702902.

Schlichting:1985:TEP

[Sch85] R. D. Schlichting. A technique for estimating performance of fault-tolerant programs. IEEE Transactions on Software Engineering, SE-11(6):555–563, June 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702053.

Scholz:1986:SRM

[Sch86] F. W. Scholz. Software reliability modeling and analysis. IEEE Transactions on Software Engineering, SE-12(1):25–31, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312916.

Schneidewind:1987:ISS

[Sch87a] N. F. Schneidewind. Introduction to the special section on software maintenance. IEEE Transactions on Software Engineering, SE-13(3):301, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702214.

Schneidewind:1987:SSM

[Sch87b]  N. F. Schneidewind.  The

state of software maintenance. IEEE Transactions on Software Engineering, SE-13(3):303–310, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702216.

Schneidewind:1989:DSS

[Sch89]  N. F. Schneidewind.  Dis-

tributed system software design paradigm with application to computer networks. IEEE Transactions on Software Engineering, 15(4):402–412, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=16601.

Scott:1987:LSL

[Sco87] M. L. Scott. Language support for loosely coupled distributed programs. IEEE Transactions on Software Engineering, SE-13(1):88–103, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702136.

Sarkar:1989:SIRb

[SD89a] D. Sarkar and S. C. De Sarkar. A set of inference rules for quantified formula handling and array handling in verification of programs over integers. IEEE Transactions on Software Engineering, 15(11):1368–1381, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=41330.

Sarkar:1989:TPV

[SD89b] D. Sarkar and S. C. De Sarkar. A theorem prover for verifying iterative programs over integers. IEEE Transactions on Software Engineering, 15(12):1550–1566, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58767.

Sarkar:1989:SIRa

[SD89c] D. C. Sarkar and S. C. De Sarkar. Some inference rules for integer arithmetic for verification of flowchart programs on integers. IEEE Transactions on Software Engineering, 15(1):1–9, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21720.

Soloway:1984:ESP

[SE84] E. Soloway and K. Ehrlich.

Empirical studies of programming knowledge. IEEE Transactions on Software Engineering, SE-10(5):595–609, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010283.

Scott:1988:SMT

[SF88] Michael L. Scott and Raphael A.

Finkel. A simple mechanism for type security across compilation units. IEEE Transactions on Software Engineering, 14(8):1238–1239, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7631.

Sanford:1981:SAN

[SG81] R. Sanford and J. Gibson. Simplifld Alfu-N2merik FOnetiks (Sanf). IEEE Transactions on Software Engineering, SE-7 (2):241–248, March/April 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702830.

Spooner:1984:UAD

[SG84] D. L. Spooner and E. Gudes. A unifying approach to the design of a secure database operating system. IEEE Transactions on Software Engineering, SE-10(3):310–319, May/ June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010240.

Skillicorn:1989:RTS

[SG89] D. B. Skillicorn and J. I. Glasgow. Real-time specification using lucid. IEEE Transactions on Software Engineering, 15(2):221–229, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=21748.

Scott:1987:FTS

[SGM87] R. K. Scott, J. W. Gault, and D. F. McAllister. Fault-tolerant software reliability modeling. IEEE Transactions on Software Engineering, SE-13(5):582–592, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702258.

Shankar:1982:FAM

[Sha82]  K. S. Shankar.  A func-

tional approach to module verification. IEEE Transactions on Software Engineering, SE-8(2):147–160, March/ April 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702922.

Shatz:1985:PFR

[Sha85]  S. M. Shatz.  Post-failure

reconfiguration of CSP programs. IEEE Transactions on Software Engineering, SE-11 (10):1193–1202, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701935.

Shatz:1988:TCM

[Sha88] S. M. Shatz. Towards complexity metrics for Ada tasking. IEEE Transactions on Software Engineering, 14(8):1122–1127, August 1988. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7623.

Shaw:1989:RAT

[Sha89]  A. C. Shaw.  Reasoning

about time in higher-level language software. IEEE Transactions on Software Engineering, 15(7):875–889, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=29487.

Shaw:1989:SSM

[SHMM89] W. H. Shaw, J. W. Howatt, R. S. Maness, and D. M. Miller. A software science model of compile time. IEEE Transactions on Software Engineering, 15(5):543–549, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24703.

Shneiderman:1981:CLW

[Shn81] B. Shneiderman. Comments from a letter written on July 19, 1978. IEEE Transactions on Software Engineering, SE-7(4):370–371, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702856.

Stonebraker:1988:PRM

[SHP88] M. Stonebraker, E. N. Hanson, and S. Potamianos. The POST-GRES rule manager. IEEE Transactions on Software Engineering, 14(7):897–907, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=42733.

Shrivastava:1981:SDS

[Shr81]  S. K. Shrivastava.  Struc-

turing distributed systems for recoverability and crash resistance. IEEE Transactions on Software Engineering, SE-7(4):436–447, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702864.

Shriver:1985:IIS

[Shr85] B. D. Shriver. An introduction to IEEE Software. IEEE Transactions on Software Engineering, SE-11(9):829, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702100.

Siegrist:1988:RSMa

[Sie88a]  K. Siegrist.  Reliability of

systems with Markov transfer of control. IEEE Transactions on Software Engineering, 14(7):1049–1053, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=42744.

Siegrist:1988:RSMb

[Sie88b] K. Siegrist. Reliability of systems with Markov transfer of control, II. IEEE Transactions on Software Engineering, 14(10):1478–1480, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6192.

Silberschatz:1983:CNT

[Sil83a] A. Silberschatz. A case for non-two-phase locking protocols that ensure atomicity. IEEE Transactions on Software Engineering, SE-9(4): 535–538, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703088.

Silberschatz:1983:ECA

[Sil83b] A. Silberschatz. Extending CSP to allow dynamic resource management. IEEE Transactions on Software Engineering, SE-9(4):527–531, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703086.

Silberschatz:1984:CDC

[Sil84] A. Silberschatz. Cell: A distributed computing modularization concept. IEEE Transactions on Software Engineering, SE-10(2):178–185, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010220.

Simon:1986:WSE

[Sim86] H. A. Simon. Whether software engineering needs to be artificially intelligent. IEEE Transactions on Software Engineering, SE-12(7):726–732, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312974.

Sinha:1983:NDA

[Sin83]  M. K. Sinha.  Nonsensi-

tive data and approximate transactions. IEEE Transactions on Software Engineering, SE-9(3):314–322, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703059.

Sinha:1985:AAR

[Sin85] M. K. Sinha. Atomic actions and resource coordination problems having nonunique solutions. IEEE Transactions on Software Engineering, SE-11(5):461–471, May 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702036.

Sinha:1986:CTT

[Sin86] M. K. Sinha. Commutable transactions and the time pad synchronization mechanism for distributed systems. IEEE Transactions on Software Engineering, SE-12(3):462–476, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312887.

Sipala:1981:OCS

[Sip81]  P. Sipala.  Optimum cell

size for the storage of messages. IEEE Transactions on Software Engineering, SE-7 (1):132–134, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702810.

Schulthess:1983:ASP

[SJ83] P. Schulthess and C. Jacobi. Anatomy of a small Pascal compiler. IEEE Transactions on Software Engineering, SE-9(2):185–191, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703036.

Schwan:1986:FSD

[SJ86] K. Schwan and A. K. Jones.

Flexible software development for multiple computer systems. IEEE Transactions on Software Engineering, SE-12(3):385–401, March 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312881.

Sevick:1987:CTP

[SJ87] K. C. Sevick and M. J. Johnson. Cycle time properties of the FDDI token ring protocol. IEEE Transactions on Software Engineering, SE-13(3):376–385, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702224.

Shen:1985:IEP

[SjYTP85] V. Y. Shen, Tze jie Yu, S. M.

Thebaut, and L. R. Paulsen. Identifying error-prone software — an empirical study. IEEE Transactions on Software Engineering, SE-11(4):317–324, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702015.

Silberschatz:1982:FLP

[SK82] A. Silberschatz and Z. M.

Kedam. A family of locking protocols for database systems that are modeled by directed graphs. IEEE Transactions on Software Engineering, SE-8(6): 558–562, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702989.

Steier:1985:REA

[SK85] D. M. Steier and E. Kant. The roles of execution and analysis in algorithm design. IEEE Transactions on Software Engineering, SE-11(11): 1375–1386, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701953.

Shin:1989:ODC

[SKhL89] Kang G. Shin, C. M. Krishna, and Yann hang Lee. Optimal dynamic control of resources in a distributed system. IEEE Transactions on Software Engineering, 15(10):1188–1198, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559767.

Smith:1985:RKB

[SKW85] D. R. Smith, G. B. Kotik, and S. J. Westfold. Research on knowledge-based software environments at Kestrel Institute. IEEE Transactions on Software Engineering, SE-11(11):1278–1295, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701947.

Sugihara:1989:MSO

[SKY89] K. Sugihara, T. Kikuno, and N. Yoshida. A meeting scheduler for office automation. IEEE Transactions on Software Engineering, 15(10):1141–1146, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559760.

Shin:1984:EER

[SL84] K. G. Shin and Y. H. Lee. Evaluation of error recovery blocks used for cooperating processes. IEEE Transactions on Software Engineering, SE-10(6): 692–700, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010298.

Sarin:1987:DOI

[SL87] S. K. Sarin and N. A. Lynch.

Discarding obsolete information in a replicated database system. IEEE Transactions on Software Engineering, SE-13(1):39–47, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702131.

Sidhu:1989:FMP

[SL89] D. P. Sidhu and T. K. Leung. Formal methods for protocol testing: a detailed study. IEEE Transactions on Software Engineering, 15(4):413–426, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=16602.

Singer:1981:AAS

[SLH81a] A. Singer, H. Ledgard, and J. F. Hueras. The annotated assistant: A step towards human engineering. IEEE Transactions on Software Engineering, SE-7(4):353–373, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702854.

Singer:1981:AR

[SLH81b] A. Singer, H. Ledgard, and J. F. Hueras. Author’s reply. IEEE Transactions on Software Engineering, SE-7(4): 373–374, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702857.

Shu:1982:SFP

[SLTC82] N. C. Shu, V. Y. Lum, F. C. Tung, and C. L. Chang. Specification of forms processing and business procedures for office automation.

IEEE Transactions on Software Engineering, SE-8(5): 499–512, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702978.

Sneed:1985:ASQ

[SM85] H. M. Sneed and A. Merey.

Automated software quality assurance. IEEE Transactions on Software Engineering, SE-11(9):909–916, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702108.

Sumita:1986:ASA

[SM86] U. Sumita and Y. Masuda.

Analysis of software availability/reliability under the influence of hardware failures. IEEE Transactions on Software Engineering, SE-12(1):32–41, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312917.

Smith:1981:ALT

[Smi81a] A. J. Smith. Analysis of long term file reference patterns for application to file migration algorithms. IEEE Transactions on Software Engineering, SE-7(4):403–417, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702861.

Smith:1981:ISM

[Smi81b] A. J. Smith. Internal scheduling and memory contention. IEEE Transactions on Software Engineering, SE-7(1):135– 146, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702811.

Smith:1988:ASP

[Smi88] C. U. Smith. Applying synthesis principles to create responsive software systems. IEEE Transactions on Software Engineering, 14(10):1394–1408, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6185.

Smoliar:1981:ORA

[Smo81] S. W. Smoliar. Operational requirements accommodation in distributed system design. IEEE Transactions on Software Engineering, SE-7(6): 531–537, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702886.

Sinha:1985:PBD

[SN85] M. K. Sinha and N. Natarajan. A priority based distributed deadlock detection algorithm. IEEE Transactions on Software Engineering, SE-11(1):67–80, January 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701899.

Snodgrass:1983:OOC

[Sno83]  R. Snodgrass.  An object-

oriented command language. IEEE Transactions on Software Engineering, SE-9(1): 1–8, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703005.

Snow:1984:USE

[Sno84] T. C. Snow. Use of software engineering practices at a small MIS shop. IEEE Transactions on Software Engineering, SE-10(4):408–413, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010253.

Sheltzer:1986:ILE

[SP86] A. B. Sheltzer and G. J.

Popek. Internet locus: Extending transparency to an Internet environment. IEEE Transactions on Software Engineering, SE-12(11):1067–1075, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312996.

Selby:1988:LEG

[SP88] R. W. Selby and A. A. Porter. Learning from examples: generation and evaluation of decision trees for software resource analysis. IEEE Transactions on Software Engineering, 14(12):1743–1757, December 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9061.

Shi:1987:VHL

[SPSP87] Y. Shi, N. Prywes, B. Szymanski, and A. Pnueli. Very high level concurrent programming. IEEE Transactions on Software Engineering, SE-13(9):1038–1046, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702326.

Schwan:1988:LSC

[SRVO88] K. Schwan, R. Ramnath, S. Vasudevan, and D. Ogle. A language and system for the construction and tuning of parallel programs. IEEE Transactions on Software Engineering, 14(4):455–471, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4669.

Srinivasan:1981:ADD

[SS81] B. Srinivasan and R. Sankar.

Algorithms to distribute a database for parallel searching. IEEE Transactions on Software Engineering, SE-7(1): 112, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702807.

Skeen:1983:FMC

[SS83] D. Skeen and M. Stonebraker.

A formal model of crash recovery in a distributed system. IEEE Transactions on Software Engineering, SE-9 (3):219–228, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703048.

Stewart:1984:SER

[SS84] W. J. Stewart and W. P.

Stohs. Some equivalence results for load-independent exponential queueing networks. IEEE Transactions on Software Engineering, SE-10(4): 414–422, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010254.

Singpurwalla:1985:ASR

[SS85] N. D. Singpurwalla and R. Soyer. Assessing (software) reliability growth using a random coefficient autoregressive process and its ramifications. IEEE Transactions on Software Engineering, SE-11(12):1456–1464, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701968.

Stelovsky:1988:SSR

[SS88] J. Stelovsky and H. Sugaya. A system for specification and rapid prototyping of application command languages. IEEE Transactions on Software Engineering, 14(7):1023–1032, July 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=42742.

Shannon:1989:MID

[SS89] K. Shannon and R. Snodgrass. Mapping the Interface Description Language type model into C. IEEE Transactions on Software Engineering, 15(11):1333–1346, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=41327.

Schoen:1988:DKB

[SSB88] E. Schoen, R. G. Smith, and B. G. Buchanan. Design of knowledge-based systems with a knowledge-based assistant. IEEE Transactions on Software Engineering, 14(12):1771–1791, December 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9063.

Sheth:1985:AEN

[SSL85] A. P. Sheth, A. Singhal, and M. T. Liu. An analysis of the effect of network parameters on the performance of distributed database systems. IEEE Transactions on Software Engineering, SE-11(10):1174–1184, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701933.

Szymanski:1985:SDT

[SSP85] B. Szymanski, Yuan Shi, and N. S. Prywes. Synchronized distributed termination. IEEE Transactions on Software Engineering, SE-11 (10):1136–1140, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701929.

Siegel:1982:PME

[SSS82] L. J. Siegel, H. J. Siegel, and P. H. Swain. Performance measures for evaluating algorithms for SIMD machines. IEEE Transactions on Software Engineering, SE-8(4):319–331, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702955.

Sahner:1987:PRA

[ST87] R. A. Sahner and K. S. Trivedi. Performance and reliability analysis using directed acyclic graphs. IEEE Transactions on Software Engineering, SE-13(10):1105–1114, October 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702150.

Shyamasundar:1989:LCS

[ST89] R. K. Shyamasundar and J. W. Thatcher. Language constructs for specifying concurrency in CDL. IEEE Transactions on Software Engineering, 15(8):977–993, August 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=31354.

Stankovic:1982:GSS

[Sta82] J. A. Stankovic. Good system structure features: Their complexity and execution time cost. IEEE Transactions on Software Engineering, SE-8(4): 306–318, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702954.

Standish:1984:ESR

[Sta84] T. A. Standish. An essay on software reuse. IEEE Transactions on Software Engineering, SE-10(5):494–497, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010272.

Stankovic:1985:SDS

[Sta85] J. A. Stankovic. Stability and distributed scheduling algorithms. IEEE Transactions on Software Engineering, SE-11(10):1141–1152, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701930.

Sunshine:1982:SVC

[STE^{+}82] C. A. Sunshine, D. H. Thompson, R. W. Erickson, S. L. Gerhart, and D. Schwabe. Specification and verification of communication protocols in AF-FIRM using state transition models. IEEE Transactions on Software Engineering, SE-8(5):

460–489, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702976.

See correction [STE^{+}83].

Sunshine:1983:C

[STE^{+}83] C. A. Sunshine, D. H. Thompson, R. W. Erickson, S. L. Gerhart, and D. Schwabe. Correction to “Specification and Verification of Communication Protocols in AFFIRM Using State Transition Models”. IEEE Transactions on Software Engineering, SE-9(1): 113, January/February 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703019.

See [STE^{+}82].

Stetter:1986:CNF

[Ste86]  F. Stetter.  Comments on

“Number of faults per line of code”. IEEE Transactions on Software Engineering, SE-12(12):1145, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313010.

See [Lip82].

Stuebing:1984:SEE

[Stu84]  H. G. Stuebing.  A Soft-

ware Engineering Environment (SEE) for weapon system software. IEEE Transactions on Software Engineering, SE-10(4):384–397, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010251.

Subrahmanyam:1985:SEE

[Sub85] P. A. Subrahmanyam. The “software engineering” of expert systems: Is Prolog appropriate? IEEE Transactions on Software Engineering, SE-11(11):1391–1400, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701955.

Summers:1987:RSS

[Sum87] R. C. Summers. A resource sharing system for personal computers in a LAN: Concepts, design, and experience. IEEE Transactions on Software Engineering, SE-13(8):895–904, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702308.

Sung:1987:PAD

[Sun87]  Y. Y. Sung.  Performance

analysis of disk modulo allocation method for cartesian product files. IEEE Transactions on Software Engineering, SE-13(9):1018–1026, September 1987. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702324.

Sarikaya:1987:TDM

[SvBC87] B. Sarikaya, G. v. Bochmann, and E. Cerny. A test design methodology for protocol testing. IEEE Transactions on Software Engineering, SE-13(5):518–531, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702252.

Svobodova:1984:RDC

[Svo84] L. Svobodova. Resilient distributed computing. IEEE Transactions on Software Engineering, SE-10(3):257–268, May/June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010234.

Stemple:1986:FAG

[SVR86] D. W. Stemple, S. T. Vinter, and K. Ramamritham. Functional addressing in gutenberg: Interprocess communication without process identifiers. IEEE Transactions on Software Engineering, SE-12(11):1056–1066, November 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312995.

Schick:1978:ACS

[SW78] G. J. Schick and R. W.

Wolverton. An analysis of competing software reliability models. IEEE Transactions on Software Engineering, SE-4(2):104–120, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702504.

See summary [Goe80].

Shoshani:1985:SSD

[SW85] A. Shoshani and H. K. T.

Wong. Statistical and scientific database issues. IEEE Transactions on Software Engineering, SE-11(10):1040–1047, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701919.

Steinholtz:1987:AIS

[SW87] B. Steinholtz and K. Walden.

Automatic identification of software system differences. IEEE Transactions on Software Engineering, SE-13(4):493–497, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702241.

Strom:1986:TPL

[SY86] R. E. Strom and S. Yemini. Typestate: A programming language concept for enhancing software reliability. IEEE Transactions on Software Engineering, SE-12(1):157–171, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312929.

Symons:1988:FPA

[Sym88] C. R. Symons. Function point analysis: difficulties and improvements. IEEE Transactions on Software Engineering, 14(1):2–11, January 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4618.

Sevinc:1988:ESB

[SZ88] S. Sevinc and B. P. Zeigler. Entity structure based design methodology: a LAN protocol example. IEEE Transactions on Software Engineering, 14(3):375–383, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4657.

Tai:1980:PTC

[Tai80] Kuo-Chung Tai. Program testing complexity and test criteria. IEEE Transactions on Software Engineering, SE-6(6): 531–538, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702779.

Takahashi:1980:ACD

[Tak80] H. Takahashi. An automaticcontroller description language. IEEE Transactions on Software Engineering, SE-6(1):53– 64, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702695.

Tamir:1980:AAD

[Tam80]  M. Tamir.  ADI: Auto-

matic Derivation of Invariants. IEEE Transactions on Software Engineering, SE-6 (1):40–48, January/February 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702693.

Tansel:1989:TEQ

[TAO89] A. U. Tansel, M. E. Arkun, and G. Ozsoyoglu. Time-by-example query language for historical databases. IEEE Transactions on Software Engineering, 15(4):464–478, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=16597.

Taylor:1986:CFR

[Tay86] D. J. Taylor. Concurrency and forward recovery in atomic actions. IEEE Transactions on Software Engineering, SE-12(1):69–78, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312921.

Thomasian:1985:IPM

[TB85] A. Thomasian and P. F. Bay. Integrated performance models for distributed processing in computer communication networks. IEEE Transactions on Software Engineering, SE-11(10):1203–1216, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701936.

Teitelman:1985:TTC

[Tei85] W. Teitelman. A tour through cedar. IEEE Transactions on Software Engineering, SE-11(3):285–302, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702005.

Tenny:1988:PRP

[Ten88]  T. Tenny.  Program read-

ability: procedures versus comments. IEEE Transactions on Software Engineering, 14(9):1271–1279, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6171.

Terry:1987:CHD

[Ter87] D. B. Terry. Caching hints in distributed systems. IEEE Transactions on Software Engineering, SE-13(1):48–54, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702132.

Tyrrell:1986:DRS

[TH86] A. M. Tyrrell and D. J.

Holding. Design of reliable software in distributed systems using the conversation scheme. IEEE Transactions on Software Engineering, SE-12(9):921–928, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313047.

Tripathi:1987:LAN

[THJ87] S. K. Tripathi, Yennun Huang, and S. Jajodia. Local area networks: Software and related issues. IEEE Transactions on Software Engineering, SE-13(8):872–879, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702306.

Thomasian:1985:PEC

[Tho85] A. Thomasian. Performance evaluation of centralized databases with static locking. IEEE Transactions on Software Engineering, SE-11(4):346–355, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702017.

Tanaka:1988:VUI

[TI88] M. Tanaka and T. Ichikawa. A visual user interface for map information retrieval based on semantic significance. IEEE Transactions on Software Engineering, 14(5):666–670, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6144.

Takahashi:1989:ESM

[TK89] M. Takahashi and Y. Kamayachi. An empirical study of a model for program error prediction. IEEE Transactions on Software Engineering, 15(1):82–86, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=21729.

Tsay:1983:MNO

[TL83] Duen-Ping Tsay and M. T.

Liu. MIKE: A network operating system for the distributed double-loop computer network. IEEE Transactions on Software Engineering, SE-9(2):143–154, March/ April 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703031.

Theimer:1989:FIM

[TL89] M. M. Theimer and K. A.

Lantz. Finding idle machines in a workstation-based distributed system. IEEE Transactions on Software Engineering, 15(11):1444–1458, November 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=41336.

Troy:1985:ASR

[TM85] R. Troy and R. Moawad. Assessment of software reliability models. IEEE Transactions on Software Engineering, SE-11(9):839–849, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702103.

Taylor:1980:RDSa

[TMB80a] D. J. Taylor, D. E. Morgan, and J. P. Black. Redundancy in data structures: Improving software fault tolerance. IEEE Transactions on Software Engineering, SE-6(6): 585–594, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702785.

Taylor:1980:RDSb

[TMB80b] D. J. Taylor, D. E. Morgan, and J. P. Black. Redundancy in data structures: Some theoretical results. IEEE Transactions on Software Engineering, SE-6(6):595–602, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702786.

Tsuchiya:1986:DDM

[TMB86] M. Tsuchiya, M. P. Mariani, and J. D. Brom. Distributed database management model and validation. IEEE Transactions on Software Engineering, SE-12(4):511–520, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312898.

Tsubotani:1986:HLL

[TMTI86] H. Tsubotani, N. Monden, M. Tanaka, and T. Ichikawa. A high level language-based computing environment to support production and execution of reliable programs. IEEE Transactions on Software Engineering, SE-12(1):134–146, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312927.

Tavernier:1980:MBC

[TN80] K. R. Tavernier and P. H.

Notredame. Macro-based cross assemblers. IEEE Transactions on Software Engineering, SE-6(4):334–340, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702741.

Taylor:1980:ADC

[TO80] R. N. Taylor and L. J. Osterweil. Anomaly detection in concurrent software by static data flow analysis. IEEE Transactions on Software Engineering, SE-6(3):265–278, May/ June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702728.

Topkis:1985:CBI

[Top85] D. M. Topkis. Concurrent broadcast for information dissemination. IEEE Transactions on Software Engineering, SE-11(10):1107–1112, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1701926.

Towsley:1986:APC

[Tow86] D. Towsley. Allocating programs containing branches and loops within a multiple processor system. IEEE Transactions on Software Engineering, SE-12(10):1018–1024, October 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313018.

See comments [Kau90].

Thayer:1981:MIS

[TPW81] R. H. Thayer, A. B. Pyster, and R. C. Wood. Major issues in software engineering project management. IEEE Transactions on Software Engineering, SE-7(4):333–342, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702851.

Troy:1986:SMS

[TR86] R. Troy and Y. Romain. A statistical methodology for the study of the software failure process and its application to the ARGOS center. IEEE Transactions on Software Engineering, SE-12 (9):968–978, September 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313051.

Thomasian:1989:RSM

[TR89] A. Thomasian and In Kyung Ryu. A recursive solution method to analyze the performance of static locking systems. IEEE Transactions on Software Engineering, 15 (10):1147–1156, October 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=559761.

Trachtenberg:1983:ODD

[Tra83]  M. Trachtenberg.  Order

and difficulty of debugging. IEEE Transactions on Software Engineering, SE-9(6): 746–747, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703119.

Trachtenberg:1986:VHT

[Tra86] M. Trachtenberg. Validating Halstead’s theory with System 3 data. IEEE Transactions on Software Engineering, SE-12(4):584, April 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312906.

Tripathi:1989:OND

[Tri89] A. R. Tripathi. An overview of the Nexus distributed operating system design. IEEE Transactions on Software Engineering, 15(6):686–695, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24722.

Taylor:1985:SAA

[TS85a] R. N. Taylor and T. A. Standish. Steps to an Advanced Ada^{1} Programming Environment. IEEE Transactions on Software Engineering, SE-11(3):302–310, March 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702006.

Tuomenoksa:1985:TSP

[TS85b] D. L. Tuomenoksa and H. J. Siegel. Task scheduling on the PASM parallel processing system. IEEE Transactions on Software Engineering, SE-11(2):145–157, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701982.

Trueblood:1986:DAE

[TS86] R. P. Trueblood and A. Sengupta. Dynamic analysis of the effects access rule modifications have upon security. IEEE Transactions on Software Engineering, SE-12(8):866–870, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312989.

Tseng:1986:RTS

[TSSP86] J. S. Tseng, B. Szymanski, Y. Shi, and N. S. Prywes. Realtime software life cycle with the model system. IEEE Transactions on Software Engineering, SE-12(2):358–373, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312949.

Tai:1986:CCC

[TT86] K. C. Tai and A. L. Tharp. A comparison of computed chaining to predictors. IEEE Transactions on Software Engineering, SE-12(8):870–874, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312990.

Tohma:1989:SAE

[TTNM89] Y. Tohma, K. Tokunaga, S. Nagase, and Y. Murata. Structural approach to the estimation of the number of residual software faults based on the hypergeometric distribution. IEEE Transactions on Software Engineering, 15(3):345–355, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=21762.

Tenma:1988:SGL

[TTTI88] T. Tenma, H. Tsubotani, M. Tanaka, and T. Ichikawa. A system for generating languageoriented editors. IEEE Transactions on Software Engineering, 14(8):1098–1109, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=7620.

Turski:1984:PI

[Tur84] W. M. Turski. On programming by iterations. IEEE Transactions on Software Engineering, SE-10(2):175–178, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010219.

Troy:1989:LNR

[TYZ89] D. J. Troy, C. T. Yu, and W. Zhang. Linearization of nonlinear recursive rules. IEEE Transactions on Software Engineering, 15(9): 1109–1119, September 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=31368.

Upadhyaya:1986:WPB

[US86] J. S. Upadhyaya and K. K.

Saluja. A watchdog processor based general rollback technique with multiple retries. IEEE Transactions on Software Engineering, SE-12(1):87–95, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312923.

Unnikrishnan:1988:TLH

[USV88] A. Unnikrishnan, P. Shankar, and Y. V. Venkatesh. Threaded linear hierarchical quadtrees for computation of geometric properties of binary images. IEEE Transactions on Software Engineering, 14(5):659–665, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6143.

Urban:1985:UES

[UUD85] S. D. Urban, J. E. Urban, and W. D. Dominick. Utilizing an executable specification language for an information system. IEEE Transactions on Software Engineering, SE-11(7):598–605, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702064.

vBochmann:1987:SIC

[vBGS87] G. v. Bochmann, G. W. Gerber, and J. M. Serre. Semiautomatic implementation of communication protocols. IEEE Transactions on Software Engineering, SE-13(9):989–1000, September 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702321.

vBochmann:1987:SC

[vBV87] G. v. Bochmann and J. P. Verjus. Some comments on “transition-oriented” versus “structured” specification of distributed algorithms and protocols. IEEE Transactions on Software Engineering, SE-13(4):501–505, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702243.

VanLamsweerde:1988:GLS

[VDD^{+}88] A. Van Lamsweerde, B. Delcourt, E. Delor, M. C. Schayes, and R. Champagne. Generic lifecycle support in the ALMA environment. IEEE Transactions on Software Engineering, 14(6):720–741, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6153.

Vermeulen:1982:CPC

[Ver82] F. L. Vermeulen. On the combined problem of compaction and sorting. IEEE Transactions on Software Engineering, SE-8(4):432–435, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702965.

Voges:1980:SAT

[VGV80] U. Voges, L. Gmeiner, and A. A. Von Mayrhauser. SA-DAT — an automated testing tool. IEEE Transactions on Software Engineering, SE-6(3):286–290, May/ June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702730.

Vuong:1988:SIP

[VLC88] S. T. Vuong, A. C. Lau, and R. I. Chan. Semiautomatic implementation of protocols using an Estelle-C compiler. IEEE Transactions on Software Engineering, 14(3):384–393, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4658.

Volz:1989:TED

[VMBK89] R. A. Volz, T. N. Mudge, G. D. Buzzard, and P. Krishnan. Translation and execution of distributed Ada programs: Is it still Ada? IEEE Transactions on Software Engineering, 15(3):281–292, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21756.

VonBochmann:1988:DID

[Von88] G. Von Bochmann. Delayindependent design for distributed systems. IEEE Transactions on Software Engineering, 14(8):1229–1237, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7630.

Voss:1980:UPT

[Vos80] K. Voss. Using predicate/ transition-nets to model and analyze distributed database systems. IEEE Transactions on Software Engineering, SE-6(6): 539–544, November/December 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702780.

VanDenBos:1983:IOT

[VPH83] J. Van Den Bos, M. J. Plasmeijer, and P. H. Hartel. Input– output tools: A language facility for interactive and realtime systems. IEEE Transactions on Software Engineering, SE-9(3):247–259, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703052.

Varvel:1989:CCE

[VS89] D. A. Varvel and L. Shapiro.

The computational completeness of extended database query languages. IEEE Transactions on Software Engineering, 15(5):632–638, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24712.

Vessey:1984:RSP

[VW84] I. Vessey and R. Weber. Research on structured programming: An empiricist’s evaluation. IEEE Transactions on Software Engineering, SE-10(4):397–407, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010252.

Wang:1988:DSM

[Wan88] Y. Wang. A distributed specification model and its prototyping. IEEE Transactions on Software Engineering, 14(8):1090–1097, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7619.

Wang:1996:TII

[Wan96] Lin Lin Wang. Thorough investigation into “An improved algorithm based on subset closures for synthesizing a relational database scheme”. IEEE Transactions on Software Engineering, 22(4):271–274, April 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=491651.

See [YLN88].

Warburton:1983:MPC

[War83] R. D. H. Warburton. Managing and predicting the costs of real-time software. IEEE Transactions on Software Engineering, SE-9(5): 562–569, September/October 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703096.

Ward:1986:TSE

[War86]  P. T. Ward.  The trans-

formation schema: An extension of the data flow diagram to represent control and timing. IEEE Transactions on Software Engineering, SE-12(2):198–210, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312936.

Wasserman:1985:EST

[Was85] A. I. Wasserman. Extending state transition diagrams for the specification of human– computer interaction. IEEE Transactions on Software Engineering, SE-11(8):699–713, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702079.

Waters:1982:PAK

[Wat82]  R. C. Waters.  The pro-

grammer’s apprentice: Knowledge based program editing. IEEE Transactions on Software Engineering, SE-8(1):1– 12, January/February 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702900.

Waters:1985:PAS

[Wat85] R. C. Waters. The programmer’s apprentice: A session with KBEmacs. IEEE Transactions on Software Engineering, SE-11(11):1296–1320, November 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701948.

Waters:1988:PTA

[Wat88] R. C. Waters. Program translation via abstraction and reimplementation. IEEE Transactions on Software Engineering, 14(8):1207–1228, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7629.

Weiss:1985:ESD

[WB85a] D. M. Weiss and V. R. Basili. Evaluating software development by analysis of changes: Some data from the software engineering laboratory. IEEE Transactions on Software Engineering, SE-11 (2):157–168, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701983.

Wuu:1985:FDD

[WB85b] G. T. Wuu and A. J. Bernstein. False deadlock detection in distributed systems. IEEE Transactions on Software Engineering, SE-11(8):820–821, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702090.

White:1980:DSC

[WC80] L. J. White and E. I. Cohen. A domain strategy for computer program testing. IEEE Transactions on Software Engineering, SE-6(3):247–257, May/ June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702726.

Wolf:1988:MVC

[WCW88] A. L. Wolf, L. A. Clarke, and J. C. Wileden. A model of visibility control. IEEE Transactions on Software Engineering, 14(4):512–520, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4673.

Wolf:1989:ATS

[WCW89] A. L. Wolf, L. A. Clarke, and J. C. Wileden. The AdaPIC tool set: supporting interface control and analysis throughout the software development process. IEEE Transactions on Software Engineering, 15(3):250–263, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21753.

Wolf:1989:MCC

[WDIY89] J. L. Wolf, D. M. Dias, B. R.

Iyer, and P. S. Yu. Multisystem coupling by a combination of data sharing and data partitioning. IEEE Transactions on Software Engineering, 15(7):854–860, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=29485.

Weber:1986:SMS

[WE86] H. Weber and H. Ehrig. Specification of modular systems. IEEE Transactions on Software Engineering, SE-12(7):784–798, July 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312979.

Wiener-Ehrlich:1984:MSB

[WEHR84] W. K. Wiener-Ehrlich, J. R. Hamrick, and V. F. Rupolo. Modeling software behavior in terms of a formal life cycle curve: Implications for software maintenance. IEEE Transactions on Software Engineering, SE-10(4):376–383, July/ August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010250.

Weiser:1984:PS

[Wei84]  M. Weiser.  Program slic-

ing. IEEE Transactions on Software Engineering, SE-10 (4):352–357, July/August 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010248.

Weihl:1987:DVM

[Wei87a] W. E. Weihl. Distributed version management for readonly actions. IEEE Transactions on Software Engineering, SE-13(1):55–64, January 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702133.

Weiss:1987:TSD

[Wei87b] D. M. Weiss. Teaching a software design methodology. IEEE Transactions on Software Engineering, SE-13(11): 1156–1163, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702162.

Wesselkamper:1982:CPS

[Wes82] T. C. Wesselkamper. Computer program schemata and the processes they generate. IEEE Transactions on Software Engineering, SE-8(4): 412–419, July/August 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702963.

Weyuker:1986:AST

[Wey86] E. J. Weyuker. Axiomatizing software test data adequacy. IEEE Transactions on Software Engineering, SE-12(12):1128–1138, December 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6313008.

See comments [ZG89].

Weyuker:1988:ESC

[Wey88]  E. J. Weyuker.  Evaluat-

ing software complexity measures. IEEE Transactions on Software Engineering, 14(9): 1357–1365, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6178.

Weinberg:1984:RWI

[WF84] G. M. Weinberg and D. P. Freedman. Reviews, walkthroughs, and inspections. IEEE Transactions on Software Engineering, SE-10(1): 68–72, January/February 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010200.

Wu:1989:SCD

[WF89] J. Wu and E. B. Fernandez. A simplification of a conversation design scheme using Petri nets. IEEE Transactions on Software Engineering, 15(5):658–660, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24717.

Wu:1984:ACS

[WH84] M. L. Wu and T. Y. Hwang.

Access control with singlekey-lock. IEEE Transactions on Software Engineering, SE-10(2):185–191, March/ April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5010221.

Woodward:1980:EPA

[WHH80] M. R. Woodward, D. Hedley, and M. A. Hennell. Experience with path analysis and testing of programs. IEEE Transactions on Software Engineering, SE-6(3):278–286, May/ June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702729.

White:1983:MIA

[Whi83] J. R. White. On the multiple implementation of abstract data types within a computation. IEEE Transactions on Software Engineering, SE-9(4):395–411, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703074.

Wile:1981:TT

[Wil81] D. S. Wile. Type transformations. IEEE Transactions on Software Engineering, SE-7(1):32–39, January/February 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702800.

Winner:1988:NBV

[Win88] R. I. Winner. Naming and binding in a vertical migration environment. IEEE Transactions on Software Engineering, 14(5):599–607, May 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6138.

Wah:1985:DDD

[WL85] B. W. Wah and Yao-Nan Lien. Design of distributed databases on local computer systems with a multiaccess network. IEEE Transactions on Software Engineering, SE-11(7):606–619, July 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702065.

Wing:1989:EIJ

[WN89] J. M. Wing and M. R. Nixon. Extending ina jo with temporal logic. IEEE Transactions on Software Engineering, 15(2):181–197, February 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21744.

Weyuker:1980:TPT

[WO80] E. J. Weyuker and T. J.

Ostrand. Theories of program testing and the application of revealing subdomains. IEEE Transactions on Software Engineering, SE-6(3):236–246, May/June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702725.

Wilson:1985:ODF

[WO85] C. Wilson and L. J. Osterweil.

Omega — a data flow analysis tool for the C programming language. IEEE Transactions on Software Engineering, SE-11(9):832–838, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702102.

Wong:1983:DRP

[Won83] E. Wong. Dynamic rematerialization: Processing distributed queries using redundant data. IEEE Transactions on Software Engineering, SE-9(3):228–232, May/ June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703049.

Wong:1984:SSD

[Won84] C. Wong. A successful software development. IEEE Transactions on Software Engineering, SE-10(6):714–727, November/December 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
5010300.

Wong:1985:F

[Won85]  H. K. T. Wong.  Fore-

word. IEEE Transactions on Software Engineering, SE-11 (10):1038–1039, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701918.

Woodside:1986:CCP

[Woo86] C. M. Woodside. Controllability of computer performance tradeoffs obtained using controlled-share queue schedulers. IEEE Transactions on Software Engineering, SE-12 (10):1041–1048, October 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6313020.

Wortman:1987:SPA

[Wor87] D. B. Wortman. Software projects in an academic environment. IEEE Transactions on Software Engineering, SE-13(11):1176–1181, November 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702165.

Wasserman:1986:BRI

[WPS86] A. I. Wasserman, P. A. Pircher, and D. T. Shewmake. Building reliable interactive information systems. IEEE Transactions on Software Engineering, SE-12(1):147–156, January 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312928.

Wasserman:1986:DII

[WPSK86] A. I. Wasserman, P. A. Pircher, D. T. Shewmake, and M. L. Kersten. Developing interactive information systems with the user software engineering methodology. IEEE Transactions on Software Engineering, SE-12(2):326–345, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312947.

Walke:1980:DOS

[WR80] B. Walke and W. Rosenbohm. Deadline-oriented servicing: Waiting-time distributions. IEEE Transactions on Software Engineering, SE-6(3):304–312, May/June 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702733.

Wetherell:1981:LAP

[WS81] C. Wetherell and A. Shannon. LR — automatic parser generator and LR(1) parser. IEEE Transactions on Software Engineering, SE-7(3):274–278, May/June 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702840.

Wegner:1983:PTM

[WS83] P. Wegner and S. A. Smolka. Processes, tasks, and monitors: A comparative study of concurrent programming primitives. IEEE Transactions on Software Engineering, SE-9(4):446–462, July/August 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703079.

Wang:1985:ER

[WS85] Ching-Chy Wang and M. L. Soffa. Eager reclamation. IEEE Transactions on Software Engineering, SE-11(4):437–439, April 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702025.

Woodside:1986:OAF

[WT86] C. M. Woodside and S. K. Tripathi. Optimal allocation of file servers in a local network environment. IEEE Transactions on Software Engineering, SE-12(8):844–848, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312986.

Wegner:1989:EH

[WT89] L. M. Wegner and J. I. Teuhola.

The external heapsort. IEEE Transactions on Software Engineering, 15(7):917–925, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=29490.

Weiss:1988:EDB

[WW88] S. N. Weiss and E. J.

Weyuker. An extended domainbased model of software reliability. IEEE Transactions on Software Engineering, 14(10):1512–1524, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6196.

Wojcik:1989:SCC

[WW89] B. E. Wojcik and Z. M.

Wojcik. Sufficient condition for a communication deadlock and distributed deadlock detection. IEEE Transactions on Software Engineering, 15(12):1587–1595, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=58770.

Weller:1984:RRA

[WY84] D. L. Weller and B. W. York. A relational representation of an abstract type system. IEEE Transactions on Software Engineering, SE-10 (3):303–309, May/June 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=5010239.

Wah:1985:SMB

[WY85] B. W. Wah and Chee Fen Yu. Stochastic modeling of branchand-bound algorithms with best-first search. IEEE Transactions on Software Engineering, SE-11(9):922–934, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702110.

Wang:1987:MRA

[WYD87] Je-Hao Wang, Tak-Sun Yuen, and D. H. C. Du. On multiple random accesses and physical data placement in dynamic files. IEEE Transactions on Software Engineering, SE-13(8):977–987, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702315.

Xu:1989:MLC

[XH89] Z. Xu and K. Hwang. Molecule:

a language construct for layered development of parallel programs. IEEE Transactions on Software Engineering, 15(5):587–599, May 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24708.

Yang:1989:DGA

[Yan89] Chao-Chih Yang. Deduction graphs: an algorithm and applications. IEEE Transactions on Software Engineering, 15(1):60–67, January 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=21726.

Yu:1988:DTR

[YBL88] P. S. Yu, S. Balsamo, and Y. H. Lee. Dynamic transaction routing in distributed database systems. IEEE Transactions on Software Engineering, 14(9):1307–1318, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6174.

Yau:1980:ACC

[YC80a] S. S. Yau and Fu-Chung Chen.

An approach to concurrent control flow checking. IEEE Transactions on Software Engineering, SE-6(2):126–137, March/ April 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702710.

Yau:1980:SSM

[YC80b] S. S. Yau and J. S. Collofello. Some stability measures for software maintenance. IEEE Transactions on Software Engineering, SE-6(6):545–552, November/December 1980. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702781.

Yau:1983:DSS

[YC83] S. S. Yau and M. U. Caglayan. Distributed software system design representation using modified Petri nets. IEEE Transactions on Software Engineering, SE-9(6):733–745, November/December 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1703118.

Yau:1985:DSM

[YC85] S. S. Yau and J. S. Collofello. Design stability measures for software maintenance. IEEE Transactions on Software Engineering, SE-11(9):849–856, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702104.

Yang:1989:ACM

[YCC89] C. C. Yang, J. J. Y. Chen, and H. L. Chau. Algorithms for constructing minimal deduction graphs. IEEE Transactions on Software Engineering, 15(6):760–770, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24729.

Yu:1985:QPF

[YCT^{+}85] C. T. Yu, C. C. Chang, M. Templeton, D. Brill, and E. Lund. Query processing in a fragmented relational distributed system: Mermaid. IEEE Transactions on Software Engineering, SE-11(8):795–810, August 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702088.

Yuen:1986:DFS

[YD86] Tak-Sun Yuen and David Huang-Chang Du. Dynamic file structure for partial match retrieval based on overflow bucket sharing. IEEE Transactions on Software Engineering, SE-12(8):801–810, August 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312983.

Yau:1981:MRP

[YG81] S. S. Yau and P. C. Grabow. A model for representing programs using hierarchical graphs. IEEE Transactions on Software Engineering, SE-7(6): 556–574, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702888.

Yu:1989:PSD

[YGBC89] C. T. Yu, K. C. Guh, D. Brill, and A. L. P. Chen. Partition strategy for distributed query processing in fast local networks. IEEE Transactions on Software Engineering, 15(6):780–793, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=24731.

Yau:1988:VCC

[YH88] S. S. Yau and W. Hong. Verification of concurrent control flow in distributed computer systems. IEEE Transactions on Software Engineering, 14(4):405–417, April 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4662.

Yao:1987:ADS

[YHYM87] S. B. Yao, A. R. Hevner, and H. Young-Myers. Analysis of database system architectures using benchmarks. IEEE Transactions on Software Engineering, SE-13(6):709–725, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702276.

Young:1980:OSC

[YL80] T. Y. Young and P. S.

Liu. Overhead storage considerations and a multilinear method for data file compression. IEEE Transactions on Software Engineering, SE-6(4):340–347, July/August 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702742.

Yang:1988:IAB

[YLN88] C. C. Yang, G. Li, and P. A. B. Ng. An improved algorithm based on subset closures for synthesizing a relational database scheme. IEEE Transactions on Software Engineering, 14(11):1731–1738, November 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=9058. See

investigation [Wan96].

Yellin:1986:AIA

[YM86] D. M. Yellin and E. M. M.

Mueckstein. The automatic inversion of attribute grammars.

IEEE Transactions on Software Engineering, SE-12(5):590–599, May 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312955.

Yang:1989:PMP

[YM89] C. Q. Yang and B. P. Miller. Performance measurement for parallel and distributed programs: a structured and automatic approach. IEEE Transactions on Software Engineering, 15(12):1615–1629, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=58772.

Yoshida:1985:TCE

[YMW^{+}85] M. Yoshida, K. Mizumachi, A. Wakino, I. Oyake, and Y. Matsushita. Time and cost evaluation schemes of multiple copies of data in distributed database systems. IEEE Transactions on Software Engineering, SE-11(9):954–959, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702114.

Yuasa:1985:IMP

[YN85] T. Yuasa and R. Nakajima. IOTA: A modular programming system. IEEE Transactions on Software Engineering, SE-11(2):179–187, February 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701985.

Yau:1988:ILC

[YNTL88] S. S. Yau, R. A. Nicholl, J. J. P.

Tsai, and S. S. Liu. An integrated life-cycle model for software maintenance. IEEE Transactions on Software Engineering, 14(8):1128–1144, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7624.

Yamada:1985:SRG

[YO85] S. Yamada and S. Osaki. Software reliability growth modeling: Models and applications. IEEE Transactions on Software Engineering, SE-11(12):1431–1437, December 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1701965.

Yu:1988:ASS

[YSD88] T. J. Yu, V. Y. Shen, and H. E. Dunsmore. An analysis of several software defect models. IEEE Transactions on Software Engineering, 14(9):1261–1270, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6170.

Yu:1985:AFA

[YSLC85] C. T. Yu, Man-Keung Siu, K. Lam, and C. H. Chen. Adaptive file allocation in star computer network. IEEE Transactions on Software Engineering, SE-11(9):959–965, September 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702115.

Yau:1986:SSD

[YT86] S. S. Yau and J. J. P. Tsai. A survey of software design techniques. IEEE Transactions on Software Engineering, SE-12(6):713–721, June 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312969.

Yau:1987:KRS

[YT87] S. S. Yau and J. J. Tsai. Knowledge representation of software component interconnection information for largescale software modifications. IEEE Transactions on Software Engineering, SE-13(3):355–361, March 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702221.

Young:1988:CSC

[YT88] M. Young and R. N. Taylor. Combining static concurrency analysis with symbolic execution. IEEE Transactions on Software Engineering, 14(10):1499–1511, October 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6195.

Young:1988:SEA

[YTT88] M. Young, R. N. Taylor, and D. B. Troup. Software environment architectures and user interface facilities. IEEE Transactions on Software Engineering, 14(6):697–708, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6151.

Yu:1988:EBB

[YW88a] C. F. Yu and B. W. Wah. Efficient branch-and-bound algorithms on a two-level memory system. IEEE Transactions on Software Engineering, 14(9):1342–1356, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6177.

Yu:1988:LDR

[YW88b] C. F. Yu and B. W. Wah.

Learning dominance relations in combined search problems. IEEE Transactions on Software Engineering, 14(8):1155–1175, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=7626.

Yao:1982:VMI

[YWH82] S. B. Yao, V. E. Waddle, and B. C. Housel. View modeling and integration using the functional data model. IEEE Transactions on Software Engineering, SE-8(6): 544–553, November/December 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702987.

Yau:1981:ADC

[YYS81] S. S. Yau, Chen-Chau Yang, and S. M. Shatz. An approach to distributed computing system software design. IEEE Transactions on Software Engineering, SE-7(4): 427–436, July/August 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702863.

Zave:1982:OAR

[Zav82] P. Zave. An operational approach to requirements specification for embedded systems. IEEE Transactions on Software Engineering, SE-8 (3):250–269, May/June 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702941.

Zeil:1983:TPP

[Zei83] S. J. Zeil. Testing for perturbations of program statements. IEEE Transactions on Software Engineering, SE-9(3):335–346, May/June 1983. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1703062.

Zeil:1989:PTD

[Zei89]  S. J. Zeil.  Perturbation

techniques for detecting domain errors. IEEE Transactions on Software Engineering, 15(6):737–746, June 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=24727.

Zweben:1989:CR

[ZG89] S. H. Zweben and J. S. Gourlay.

Comments, with reply, on “Axiomatizing software test data adequacy” by E. J. Weyuker. IEEE Transactions on Software Engineering, 15(4):496–501, April 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/

                                        stamp.jsp?arnumber=16609.

See [Wey86].

Zhou:1988:TDS

[Zho88] S. Zhou. A trace-driven simulation study of dynamic load balancing. IEEE Transactions on Software Engineering, 14(9):1327–1341, September 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6176.

Zhao:1987:VTC

[ZR87] Wei Zhao and K. Ramamritham. Virtual time CSMA protocols for hard real-time communication. IEEE Transactions on Software Engineering, SE-13(8):938–952, August 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702312.

Zhao:1987:STR

[ZRS87] Wei Zhao, K. Ramamritham, and J. A. Stankovic. Scheduling tasks with resource requirements in hard real-time systems. IEEE Transactions on Software Engineering, SE-13(5):564–577, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702256.

Zave:1986:SFE

[ZS86] P. Zave and W. Schell. Salient features of an executable specification language and its environment. IEEE Transactions on Software Engineering, SE-12(2):312–325, February 1986. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312946.

Zhang:1988:IPS

[ZTSN88] Y. X. Zhang, K. Takahashi, N. Shiratori, and S. Noguchi. An interactive protocol synthesis algorithm using a global state transition graph. IEEE Transactions on Software Engineering, 14(3):394–404, March 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4659.

Zhongxiu:1987:CDP

[ZX87] Sun Zhongxiu and Li Xining.

CSM: A distributed programming language. IEEE Transactions on Software Engineering, SE-13(4):497–500, April 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702242.

Zhu:1987:SED

[ZY87] Chuan-Qi Zhu and Pen-Chung Yew. A scheme to enforce data dependence on large multiprocessor systems. IEEE Transactions on Software Engineering, SE-13(6):726–739, June 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702277.